A Complete Bibliography of Publications in Cold War History

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

25 April 2024 Version 1.00

Title word cross-reference

circ [Ons16]. ^{o} [P´ 24].

//doi.org/10.1080/14682745.2014.959500 [Ano15c]. //doi.org/10.1080/14682745.2014.971017 [Ano15b].

10.1080 [Ano14a]. 10.1080/14682745.2011.625160 [Ano14a]. 11 [Mas02]. 12 [Ano14a]. 14682745.2011.625160 [Ano14a]. 14682745.2014.959500 [Ano15c]. 14682745.2014.971017 [Ano15b]. 1928 [Zub16]. 1930s [Wil24]. 1940 [Arn03]. 1946 [Pra02]. 1950s [Bor01, Byr19, Cas19, Che21, Far13, Grü03, LAGE21, Lib11, Nav11, Ris11, Rob16b, Tob02, Tun15]. 1954 [Ass11]. 1956 [Dei06, Gon19, Kap14]. 1957 [Pea12]. 1960 [Dra13]. 1960s [Blu01, Byr19, Cul18, Fri10, Kat24, Las24, Ped22, Ris11, Rob16b]. 1961 [Her07]. 1962 [Far15b]. 1964 [QJ09]. 1965 [Lüt09]. 1966 [Eas05]. 1967 [Mag22]. 1968 [Roy02, Kra11]. 1969 [BK03, Fis09, Pec17, Lud09]. 1970s [Cha12, Gat22, Ger19b, Kir20, Kom20, Mit17, Pal14, Pet14, Rea12, Rom14a, Sar13, Wil24, Zub08]. 1972 [Bur22]. 1974 [Lak21]. 1975 [Zac20]. 1979 [Gal19]. 1980 [Cor13]. 1980s [Pal14, Pet12a, Rub14, Sny10, Tre20, ZM20, Zub17]. 1986} [Har23]. 1989 [Wes10]. 1990 [Mar19, Sny13b]. 1990s [hHC15]. 1992 [Ons16]. 1994 [Ons24].

2 [Dra13]. 2010 [Han13].

4 [Lüt09].

5 [Fis09]. 53/55 [Mue06].

986pp [Han17].

A. [Gra20b]. Abdel [McA11]. Able [Man09]. absorb [Com20]. Academic [Dee10]. accidents [Arn21]. Accommodating [Pop10]. According [Lun06]. accords [Kal08, Dai19, Sch20a]. Acero [P´ 24]. Acheson [Kit12]. across [Gat22, Str23]. Act [Dav09, Sch20a]. Action [Hus05, Jon20]. Active [Eas15]. Activism [Ped13, Col18, Ped16]. activists [Ren18]. Actors [Woo13]. adapt [Glo18]. Adenauer [Gra11]. administration [Bra21, Bro20a, Gar13, Mac11, Sny13a, Tys23, Gua02, Moo02, Var03]. Admiral [Dav03]. adoption [Sha17]. adult [Jon20]. Adversary [Ren00, Rob16a]. Advertising [Tob11a]. advice [Zai09]. advisors [Sny12]. advisory [Che22]. Aegean [Riz09]. aerospace [Sam14]. Affair [QJ05, Oña11, Sba09, Her03]. affairs [Pal14, Mar22b]. affected [Tre20]. Afghan [Kli23, Sta24]. Afghanistan [Her03, Kal08, Kal14, Lah13, Sta24]. Africa [Lop13, Ons16, vdH22, Ahl13, Ano07a, Ano07b, Gas18, Ger19a, Gue10, Gue24, Ian12, Mil12, O’M12, Ons05a, Ons05b, Ons13, Ons22, Sau17, Say16, Sch23, Shu07, Shu12, Shu15, Shu24, Web22, vW15, vW24]. African [Ell16, Mac22, Bai18, Cam15, Dai19, Eis18, Kat24, Kir20, Kon21, Mar22b, Mil22, RB24, vW07, vW10]. Africanism [Kat24]. Afro [Ewi19, Ger16]. After [Cas11a, Ale10, Arn21, Do22, Heu14, Lan22, Rei16, Rot23, ZM20]. aftermath [Mat15, Sta13]. Against [Set13b, Bar15, Bau16, Cli19, Fai15, Nie21, Sal18, Wet09, Whi09, Zub14]. age [Cas19, Lev14, Mor14a, Pal14, Syl22, Cig17, Cra15]. agency [Kir20, vW15, Coo22, Cul03, Hol16, Mös23, Roe16, Sha10]. aggravation [Nad04]. Aggression [Goo07]. Agreement [Bid24, Bur22, Gal19, Ban21]. Agricultural [Dra13, Kol18]. Aid [Lon24, Maz07, Cal19, Gat22, Grü03, GU22b, Sch23, XZ19]. Aiping [Liu18]. Air [Heu09]. airlift [Lev08]. al [Fie12]. Albania [Lon24, Mar17, Var12]. Albanian [Lon24]. Alert [BK03, Nut17]. Alessandro [Hof13]. alien [Cox21]. Aligned [Her07, Mav21]. alignment [DeR07, Hay18, Kir20]. Allemands [Lud09]. Allen [Mag19, Zub16]. Allende [Gus12, Per11].

Alliance [Blu01, Fri17, Cas15b, CyJ17, Fre14, Gol09, Gra11, Loc09a, Loc09b, Lud11, Ris19, Lou17b]. allies [Riz09, Mag13]. Ally [Har00, Ren00, vW07]. almost [Sfi14]. already [Mad05]. Alvandi [Cha15]. am [Gra20b]. ambassador [LS19, Bro12]. Ambiguity [Cos00]. Ambiguous [Tub05, Tul12]. ambitions [Ada21]. America [Han13, Han17, Arn13, Ass12, Cas19, Cha15, Che21, Cox13, Cra15, Cra16, Dip18, Dri11, Dül12, Fog20, Hei17, Hei11, How21, Loc09a, Mis12, Nat18, Nav11, New11, Noe12, Pag21, Pet11, Rey23, Roy02, Set13b, Spo15, Tob11a, Tun15, Wee21]. américains [Mag22]. American [Ano15c, Cas20, Ces13, Cox13, Dro13, Jos19, MZ22, kM10, Par12, Pri22, Rob16a, See02, AK15, Bar17, Bou08, Bra16, Bur01, Cam08, Con05, Cox07, Cre02, Dro15, Ell06, Fai15, Foo11, Gol09, Gri08, Gus12, GU22b, Han11, Han14, Her19, Hul12, Hyu22, Joh23, Kin24, Kit12, Kra14, Kui00, Lat11, Lib11, Lon24, Lüt09, Mag09, Min16, Mit17, Nic02, Oso07, Peo08, Per17, Per16, Pra11, Pre09a, Pre13, Rut13, Sar09b, Sar15, Sba09, SS09, Sha10, Sny12, Sny13c, Spe13, Ste18, Sur08, Tob13, Tob02, Tót20, Woo10, Zhu14, vW11]. Americanization [Ber10]. Americanized [Tob11b]. Americans [Cam15, Eli10]. amid [Nem10]. Amidst [Oso07]. amministrazione [Chi14]. among [Ber10]. Amoureuse [Die02]. Analysing [ZCP^{+}21]. Analysis [Sto03, Pat11]. Anatomy [Lev08, Hof13]. ANC [Sim24]. Andreas [KC19, Naf13]. Andrei [Kos11]. Andropov [Che14, Say09]. Angela [Rom18a]. Angeles [Sar09a]. Angeli [Chi14]. Anglo [Cre02, Cul18, Ell06, Fur09, Hul12, Lon24, Sar15, Spe13]. Angolan [Wid18]. angry [Dum14a]. Ann [Rie14]. année [Wes10]. Anniversary [Kad14]. anonymity [Con23]. Anti [Low07, QJ05, Col18, Mar21, Nut14, Ste16]. anti-colonial [Ste16]. Anti-communism [Low07, Nut14]. anti-nuclear [Col18]. Anti-Party [QJ05]. anti-revisionist [Mar21]. Anticommunism [Wil24, Bec23, Joh23]. anticommunist [Bau23, Ger19c]. antinuclear [Rub14]. anxieties [Sar15]. anyway [Sun22]. apart [Mil12]. Apartheid [vW10, Bai18, Ons24, vW11, vW15]. Apocalypse [McM10]. appeal [Bra16]. appearing [Bra16]. Application [Par06]. appointment [Zan12]. apprivoisée [Lud09]. Approach [SL15, Sny13a]. approaches [Boe19]. appuntamento [Zan12]. April [Kap14, Ons24]. Arab [All24, Cit12a, Cit12b, Gfe11, Gin02, Hir20, Pop10, Sto12]. Arabia [Par12]. Arabian [Par12]. Arabism [Aba02]. Arabists [Cha15]. Arbor [Rie14]. Arc [All24]. arcane [Jac16]. Archer} [Man09]. Archie [Hil21, Pat07]. Archival [GR06, Sty01, Pat11, Pre22]. Archives [Gro05, Kap03, Lud01, Maz07, Mel07, Muk06, Ons05b, Rad03, Yeg06, Cav11, Mar22b, Web22]. Arctic [Laj13]. area [Cho12a, Hat13]. Arena [Pet14]. Argentina [Har15, Kel22]. Argentine [Sch12]. Armageddon [Bjo23]. Armed [Ped01]. armée [Mag22]. Armenian [Aha19]. armies [Jag22]. Arms [Blu12, Gas24, Eis18, Lev14, McD12, Sun22, Wys12]. Army [Nge23, Rei10, Win09]. Arras [QJ09]. Art [Per16, Wen22, Bro20b]. Article [Ano15b, McM06, Ald01, Gua02, MV23, PP23, Ano05]. artistic [Vel19].

Artois [QJ09]. Asia [Ben05, Bes12, Cas11a, Che21, Ewi19, Kra14, MM05, Oso07, Shi20, Tan14]. Asian [Ger16, BR23, Cav11, Jag22, SX14]. Aspect [GR06]. assassination [McP21, Rob17]. assassinations [Nie21]. Assessing [Rob16a]. assessment [Gle16]. assimilation [BR23]. Assistance [Sto03, Coo22, LAGE21, Lev08]. Association [LS24]. Astrain [P´ 24]. astray [New11]. Atlantic [Loc09b, Sea21]. Atomic [Coo22, Ghe13, Hol16, Mös23, Roe16, Tun15, Van01a, vW10, Arn03, Ell22b, Gil09, Gri23, Cig17, Cra15, Twi16]. Atoms [vW15, Twi16]. atrocity [Lea23]. attack [Lüt09]. attempts [App08, Asi24, CS20]. Attitudes [Con05]. Attrition [Ada06, Gin02]. Audience [Myt11, Cla13]. August [Pec17, Per19]. Australia [Ben05, BR23]. Australian [BR23, Gra20b]. Austria [Mue06]. authoritarian [Spe10]. Averell [Soa09]. aviation [Ell22a, TL23]. avoid [Gra10]. Awakenings [Sur06]. away [Cla08, Sar15]. Azerbaijan [Rai01, Sha17].

B [Ano14a, Ell14, Ler11, Ped13, Sau13]. Back [Ald01, Heu14, Zub14]. backyard [Mou15]. balance [Cas20, Fis10]. Balancing [Ger19b]. Baling [HCL19]. Balkan [Cav11, Cho12a, Sta19]. Ballet [Gon19, Gon19]. Baltic [Ale10, Pii09, Rea06]. bamboo [Dum10]. ban [Zai09]. Bandung [Zha07, Ewi19, Ger16]. Bangladesh [Lea23]. Bank [Spi06, Ked18]. Bargaining [Gat22, Sch08a]. Bark [Goo07]. barricades [Bjo21]. Barton [Dan22]. basculé [Wes10]. base [Alv14]. based [Ris07]. Basic [Cha15, Fra18]. basin [Mou15]. Ba’thist [Sit18]. Battle [Ons16, Roy02, Set13b, Sca23]. Battleground [Ahl13]. Bay [Pet12b]. BBC [Maj13, Myt11]. Be [Fil06, Gar23]. Bear [Del07, Sch08a]. Bears [Ada06]. beating [Liu18]. Beautifully [Sim05]. bedfellows [McP21]. been [Cas20]. begining [Nad04]. Behind [Ful11, Maj13]. beholder [Hyu22]. Beijing [Alb14, Kha11, XZ19]. Beiträge [Kra11]. Belgian [Maz07]. Belgium [Duj07, Gij16]. Belgrade [Her07, Rak14]. Benedikt [Lud09]. Berghahn [Sny13b]. Berlin [Glo18, Ano15c, Ash09, Buf09a, Car11, Dei13, Dro15, Ful11, Giv22, Gro10, Grü03, Har00, Lun03, Tra06, Tra21]. Bernard [Her12]. Bernd [Sny13b]. Best [Hae15, Mav21, Tre20]. betrayal [Jer12]. better [Ger16, Gar23]. Between [Duj07, Kir20, Pec17, Rea06, Bur22, CyJ17, Ger19b, hHC15, LK20, Mis12, Nad04, Pon02, Ris07, Ris09, Riz09, Sit18, Smi12, TL23, WN08, Fis21]. Bevin [Sha17]. Beyond [All24, Lev13a, Sny13a, ZVU18, Bur19, Gno18, Kie20, Tag20]. bid [Sch08a]. Big [Mag13, Pec14]. Bilateral [Tag17, Far17]. bind [Gol09]. Bingchang [Foo09]. bipolar [Glo18]. bipolarity [Mar23]. bipolarization [vA09]. birth [Lea23]. birthdays [Jac16]. Bitterness [Ste16]. Black [Ano07a, Pri22, Gue10, Gas18, Rom18a]. bloc [Boe19, Sta13, Tag20, UT18, Far22, Sza17]. Bloody [Hug06]. Blue [Gin02]. bluff [BB14]. Board [Ano10b, Ano11b, Ano12a, Ano13, Ano14b, Ano17b, Ano00a, Ano01a, Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano03b].

Bogotá [Kar16]. Böhlau [Kra11]. Boiling [Eas05]. Bolshoi [Gon19]. Bomb [Ell15, Bar15, Gil09, Tob16, ZM20]. bombs [Arn03]. bonds [Sea21]. Bonn [Ban21]. Book [Ano03a, Ano04, Bor20, Boy14, Ces13, Cha15, Chi14, Cox13, Cox24, Cro08, Dan22, Doc08, Dro13, Fie12, Fis21, Han13, Han17, Hil13a, Hil21, Hof13, Kar21, Ken08, Ken16, Kra11, Lak21, Lat09, Lin13, Lop13, Lor24, Lud09, Mag19, Mag22, Man13, Mar19, Mue24, Ons16, Ped13, P´ 24, Pri22, QJ09, Rie14, Rob16a, Ruo05, Sau13, Sch22, Sch20b, Sch01, Sch02a, Sch02b, Sch03, Set13a, Sny13b, Som13, Tag20, Tho05, Wes10, Wes12, Zub16]. Books [Ano09a, Ano10a, Ano11a, Ano15a, Ano16a, Ano17a, Ano18, Cha15, Han17, Sny13b, Kad14]. Boothe [Bro12]. Border [Bur01, Jos19, Kui00, SL15, Jos20, Kra12]. borders [Cas20, Dum14b]. Botswana [Kir20]. boundaries [Cav11]. Bourgeois [Fil06]. boy [Old20]. boycott [Cor13]. Bozo [Sny13b]. Brandt [Ell11, Lot09]. Brazil [Ban21, Con23, Har12a, Her07, Lou17b, Lou17a, Sea21]. Brazilian [Lou17b, Pat15]. Breaking [Mar23]. Brent [Sny13c]. Bretton [Roy02]. Brezhnev [Lot01]. bridge [Dum10]. bridge-building [Dum10]. Brief [BBF24, Kim23]. Brigades [Hof13]. Brill [Tag20]. Brinkmanship [Lun03]. Britain [Aba02, Ban05, Bes12, Boz13, Bus01, Cra16, Dei13, Die02, Dip18, Ell10, Gre09, Kap14, Loc09a, Maw13, Par06, Ped01, Rak14, Roy02, Sai21, Smi12, Var03, You02]. British [Ben05, Cor13, Cox10, Day22, Dei06, Dum14a, Ell15, Foo13, Ger16, Hop01, Hug02, Hug04, kM10, Par11, Rab09, Rie19, Say16, Sco10, Sha17, Sme23, Sto10, Twi16, Ull11, UK14]. Broadcasting [Ris13, Maj13, Tob13]. brotherhood [Koc08a]. Brown [Hil21, Pat07]. Bucharest [Pec17]. buckskin [Hum20]. Budget [Boy14]. build [Lod10]. build-up [Lod10]. Building [Har00, Mou15, Dum10, Lor24, Pea12, Rei10]. Built [Som13]. Bukovsky [Uli14]. Bulganin [Kap14]. Bulgaria [Wol10, Zof18]. Bulgarian [Par11]. Bullard [Sha17]. Burma [FZ23, Nge23, ZF19]. Burn [Rot23]. Bush [Fre02, Hor20, Sny13a]. bushy [Fis09]. Business [Alb14, GR06]. Busting [Kad14]. Butcher [MO10]. buy [Sch08a, Tob13].

C [Dro13]. c.1945 [Hae15]. ca.1920s [Low07]. cables [Mat15]. Caetano [MO10]. Cairo [Ste16]. calculation [ZX15]. Calculations [Rea06]. Call [Ano21, Ger16, Gut21]. Callaghan [Ell15]. Calling [BB14]. Cambodia [QJ11]. Cambodian [Ban05]. Cambridge [Cox13, Han13, Kar21, Lak21, Lop13, Lor24, Ped13, Pri22, Cro11]. came [Cas11b, Vel19]. Camp [Dai19]. campaign [LK20, Nie21, Sch12, Wet09, Zof18]. Campaigning [LK20]. campaigns [Fai15]. Campbell [Cox13]. Canada [Hum20, McK12, Mus21, Rhé11, Web21]. Canadian [Col18, Laj13]. cannot [Cra16]. capability [Ger10]. capitalisation [Her19]. Capitalism [AC17, Cox18, kM17]. capitalist [Asi24]. Captive [Cla13]. care [Sit18]. Caribbean [Maw13, Mou15]. Carole [Lak21]. Carolina [Dro13]. Carré [Old20]. Carter [Chi14, All24, Fra18, Lud11, Mit07, Sau17, Sny12, Tys23, Woo10]. Casals [Jos20]. Case [Mis06, Och06, Sty05, Tsv24, Fen20, FM15, Gno18, Jac16, LS19, LS24, Mar09, Sco10, WN08]. cash [Sun22]. Castro [Pet11, Pra24, Whi09]. Catarate [Fie12]. Catholic [Liê05, Man21]. Catholics [Bjo21]. Caucasus [Küs02]. Caught [Kal14]. caused [Lar14]. Causes [Kho05]. Cautious [Rhé11]. CCP [Cli19]. CEE [Boy14]. celebrity [Bel23]. Celluloid [Buf09a]. Central [Byr17, Kra14, Mus21]. Centre [Pop15]. Centro [P´ 24]. Century [Han13, Har15, Bro10, Dri11, Fra20, Heu14, Sca23, Sco10]. certain [Mag19]. Césari [QJ09]. challenge [Fri10, Gij16, Glo18, Gre09, Jos20]. challenged [Wid18]. Challenges [Her19]. chance [Gud04, Sta19]. Chancellor [Sch08a, Sch14]. Change [Yam07, Doo21, Pha12, Sur08]. changed [Soa09]. changing [Min16]. chaos [Nem10]. Character [Han15]. Charles [Mag19]. Charlie [Man21, Tra06]. charms [Zhu14]. Che [Nav11]. Checkpoint [Tra06]. Chernenko [Che14]. Chester [Col12]. Chevaline [Sto10]. Chiang [Foo09]. chicken [Gar23]. Chiefs [BK03]. child [Die11]. Childhood [Bar17]. Children [Cos23, Kal11]. Chile [Gus12, Ked18, Per11, Pum22, Sal18, TL23]. Chilean [Uli14, Vel19]. China [Tag20, Alb14, Aut21, Bör18, Bur22, Che19, Cio14, Dum10, FZ23, Far17, For21, FS23, Gno18, Ham21, HCL19, Jun11, Kra12, Li11, Liu18, Lüt07, kM17, Mar17, Min16, QJ11, SX10, SL15, Sun20, Tag17, UT18, Vám18, Wes17, Wil24, Zha07, Zhu12, ZVU18, Zof18, ZF19, Zub17]. Chinas [Kir20]. Chinese [Che14, Che22, Con05, Eis18, Fri10, Gu22a, Kha11, Kim10, Kim12, Knü20, Kra14, Lüt09, kM10, Mar21, Sun22, Wil24]. Cho [LS19]. choix [Boy14]. Christian [Fis21, Sca23]. Christians [Pre12]. Christmases [Jar18]. Chung [RS17]. CIA [Cha15, Bec23, Gle16, Mar12, Tok22]. cinema [Dav03, Dum14a, Mag09, Zhu14]. Circle [Tre20]. citizen [Fog20]. City [Mag09]. Civil [Ell22a, Kim10, Oña11, Tob02, Bar15, Bjo23, Jar18, Syl22, Twi16]. civilian [Gut21]. civilisation [Bel15]. Clandestine [Ald01]. Clare [Bro12]. Clarifying [BW07]. Clash [Kui00, Die11]. Claude [Och06]. Clay [Tra06]. client [Cas15b]. climate [Doo21]. close [Gut21]. Closed [Lud01]. clothing [Fai15]. Co [SS07, Kha11]. Co-existence [SS07]. co-operation [Kha11]. Coast [Ger16]. Coercion [Bor01]. Coexistence [Boq05, QJ11, Smi12, Zha07]. Cold [Ano14a, Ano15b, Ano15c, Bor20, Bra07, Cas20, Ces13, Cha15, Cos23, Cox13, Dro13, Far22, Fis21, Gus12, Her07, Hil21, How21, Hus05, Jos19, Kin24, Lak21, Lop13, LS24, Mag09, Mar16, Mue13, Mue24, Pec01, Pec14, Ped13, Pop10, Pop15, Rie19, Rob16a, Sau13, Sch20b, Set13a, Set13b, SL15, Sny13b, Som13, Tag20, Tob11b, Ulu03, Wee21, Gle16, Mat10, Sha03b, Wid18, vW24, Far15a, Aba02, Aha19, Ahl13, Alc18, Ald01, Ale10, Ano07a, Ano21, Arn21, Arn13, Bar17, Bau16, Bau23, Bec23, Ben05, BBF24, Bes12, Boe19, Boo16, Bou08, Boz07, Bra16, BR23, Bro11, BW07, Bro07, Bro20b, Cam08, Car11, Çel22, Cho13, Cho12b, Cit12a, Cit12b, Cit19, Com20, Con10, Cor13, Cos00, Cox07, Cox14, Cox21, Cra11, Cro09, Cro11, Cul13, Cul18]. Cold [Dal14, Dav03, DeV11, Dee06, Dei13, Die11, Dim08, Doo21, Dri11, Dül08, Dum14a, Ede17, Eli10, Ell22a, Ell11, Eme10, Far15a, Far17, Fis10, Fog20, Foo13, For19, FH14, Gas18, Ger19c, Gil09, Gil10, Gle18, Glo18, Gno17, Gon19, Gri23, Gro15, Gue10, Gue24, Hae15, Hal20, Ham21, Han14, Har12a, Har12b, Har08, Her19, Her12, Heu14, Hil13b, Hir20, Hop01, Hop08, Hor20, Hug02, Hum20, Hyu22, Jag22, Jon20, Kan06, Kec24, Ked18, Kel10, Kha11, KK13, Kin13, Knü20, Kra14, LPT16, Law11, Lev13a, Lev14, Lev13b, Lib11, Lop22, Lot03, LK20, Lou17b, Lou17a, Lun00, Lun06, Lüt17, Mac22, Mad05, Mag17, MM03, Maj03, Maj13, Man13, MZ22, Mar23, Mar17, Mas02, Mas14, Mat15, Mav21, Maw13, McD12, McP21, Mey21, Mil22]. Cold [Mis12, Mit07, Mit18a, Mit18b, Mor14b, Mou15, Mue06, Muk06, Mum17, Myt11, Nad04, Nat18, Neh04, Noe12, Nut14, O’M12, Ole08, Ons22, Ons24, Oso07, Pag21, Pai06, Pat21, Pat07, Pea12, Ped16, Per17, Pet11, Pii09, Ple13, Pre12, Pre13, Pro23, Puc20, Pum22, Rab09, RB24, Rea06, Rei10, Rey13, Rey23, Ris13, Rob16b, Roe16, Rog12, Rom14a, Rom11, Rom14b, Rom18b, Rut13, Sal10, Sar09a, Sar13, Sar15, Sau17, Say16, Sba09, Sca12, Sch23, Sch20a, Sch12, Sch11, Sea21, Sha10, Sha17, Shi09, Sho18, Shu12, Shu24, Sim24, Smi01, Smi12, Sny18, Soa09, Spi06, Spo15, Sta23, Sun20, Sur06, Tam10, Tan14, Tau09, Tob13, Tok22, Tót20, Tra21, Tsv24, Tub05, Var08, Vit17, Voß16, Web13, Web21, Web22, Wen22, WN08, Wes06, Wes12]. Cold [Wes17, Wet09, Wil11, Wil24, Wol10, Wys12, Yeg06, Zhu14, ZVU18, Zub02, ZCP^{+}21, dK09, vA09, Lin13]. collaboration [Her19]. collapse [Cox14, Mil12, Riz12, Woo13, ZCP^{+}21, Cox24]. Collection [Ons22, vdH22]. Cologne [Kra11]. Colombo [Ewi19]. Colonels [Mar09]. colonial [Ste16]. colonialism [Byr19]. Colonies [Maz03]. colour [Cam15]. Columbia [Cox13, Dro13]. Come [Cos00]. COMECON [Ger19b]. comics [Lop22]. Cominform [SX14]. Coming [Muk06, Fog20]. Comité [Boy14]. commander [Tra21]. Commemoration [Far15a]. comment [Cam08]. commentary [Jos20]. Commercial [TL23]. Commitment [Boq05, Cre02, Hop01]. Committee [Kec24, Wid18, Goo07]. Common [Rey04, Yam07]. communism [Bag11, Gno18, Low07, Nut14, Var12, Wol10, Zub10, Kat24, Mor14a, Sal18, Set13b, Sun20]. Communist [Eis18, Ell16, Pon02, Sba09, Har21, Lus23, Mar21, Shi09, Efi05, Far13, Foo11, Maj03, Pon02, Pra02, Sta13, Whi10, Zan12]. communists [Bor04, Kli23, Mis12, Lod10, Sfi14, UK14]. Community [Par06, Sim05, Loc09b, Mac11, Dwa01, Kom20]. companion [Ell14]. Company [Par12]. Comparative [Gue24, MM03, Sho18]. Competition [Lüt17, Bra16, Gri08, Lot19, Pal14]. complementary [LPT16]. Complex [Say09, Set13b, Mey21]. complicated [Gar13]. Complications [Zar05]. compromise [Cal19]. compromiso [Fie12]. Comptes [Boy14]. computers [Bou08]. Comrades [Do22, Eis18, Zar05]. Comrades-in-arms [Eis18]. comradeship [CyJ17, Cra11]. comunisti [Zan12]. concept [BR23, Rey04]. Cone [Har12a]. Confederation [Gri23]. conference [Rak14, Dra13, Ass11, Ger16, Her07]. Conferences [Zha07]. confidence [Tro22]. Conflict [Par11, FH14, Noe12, Riz09, Rog12, Sto12, Wes12, Wid18, Wil11, Nic02, Lin13]. conflicto [Fie12]. conflicts [Lar13, Mou15]. conflitto [Wes12]. confront [Sam14]. ’Confrontation [Gua02, Cra11, Nad04, Soa09]. confrontations [Gri08]. Confronting [Mis12, SS07]. confronts [D’H18, Rab09]. Confusion [Her07]. Congo [Ahl13, Gij16, Mar22a, Maz07, Noe12, Shu12, Sun22]. Congress [Man21, Shi20, Ull11]. Consensu [Pin24]. consensus [Pea12]. consequences [Cro09, Sim24]. conservatism [Joh23, Mor14a]. Conservative [Han08, Jur17, Sam12, Koc08a, Lib11]. consideration [Kim16]. consolidation [Koc08a]. Constitucionales [P´ 24]. constitutional [Çel22]. Constructing [Zha07, Sau13]. construction [Arn03, Lat11]. contacts [Gut21]. containment [Fis21, Pre09a, Sny13a]. contested [Koc08b]. Contesting [Per17]. context [Alb14, Neh04]. Continent [Die02]. Continental [Cre02]. continuing [Jos20]. continuity [Pha12]. contours [Che21]. Contra [Her03]. contribution [Liu18]. Contributions [Kin24]. control [Blu12, Lev14, Par12]. Controlling [Par12]. Controversial [Han17]. Controversies [MN01a, MN01b]. Controversy [Ell15, Bar15]. Conversations [Ano21]. conversion [Woo10]. Cooper [Col12]. Cooperation [Die02, Com20, Gut21, Hul12, Kol18, Kon21, Sar21, Sta19, Bur22]. cooperatives [Kol18]. core [Gro15]. Cornell [Ces13, Hof13, Mue24, Som13]. Corporate [Par12]. correspondence [Pec14]. Corriere [Lib11]. Corrigendum [Ano14a, Ano15b]. Corvalán [Uli14]. Costa [Bro20a]. costs [Smi12]. Council [MD10]. counter [Sto12]. counter-revolution [Sto12]. Countering [Sto03]. counterinsurgency [Mar12]. counterinsurgent [Dri11]. counterrevolution [Bel15]. countries [Bau16, Cav11, TL23]. country [Foo13, Sai21]. Course [Han15]. Courting [Mav21]. Couscous [McA11]. Coventry [Bar15]. Covert [Eas05, Hus05, Lon24, Lop22, Voß16, Whi09]. Cox [Cam08]. crafting [Ewi19]. Craig [Cox13]. Cravens [Man13]. create [Mat10]. Creation [Rai01, Hol16, Kit12, MD10, Roe16]. creole [Bec23]. crises [Gno17, Maj21]. Crisis [All24, Fre02, Roy02, Sta13, Sza05, Cho14, Cho12b, Eas15, Fil23, Ger19a, Gre09, Jun11, Kra11, LS19, Nat18, Nay18, Per12, Pra12, Rea10, Sha17, Tre20, Ano15b, Ash09, Ban05, Cam15, Ell06, Far15b, Glo18, Kad14, KC19, Kel22, Lun03, Mar22a, Sto14, Sty01, Sun22, Tra06]. Critic [vW07]. Criticism [CP00, Sco09]. critique [Ass11, NZ12]. critique. [Ano14a]. Cross [Wid18]. crossed [Cas20]. crossing [Dum14b, Jos20]. crossings [Jos19]. Crossroads [WS01, Dül12, Rom14b]. Crucible [Sta24]. Cruising [Col18]. crusade [Sca23]. crusader [Col12, Man21]. crush [Cit12a, Cit12b]. cry [Sun20]. CSCE [Bad18, Boz13, Dav09, Fis09, Sny10, Sny13a]. Cuba [Boo16, Gle07, Gri08, Hul12, Hus05, McK12, Pra24, Whi09, Fie12]. Cuban [Ano15b, Far15b, Kad14, Cam15, Col23, Kar16, Kel22, Oña11, Ped22]. Cultural [Cos00, Shi20, Som13, Tsv24, Ber10, Dum14a, Ger19a, MM03, Mik20, Sba09, Vel19, Wen22, Mar17]. Culture [Low07, Car11, Ped22, Syl22, Bro20b]. curiously [Tro22]. Curricula [KW17]. curtain [Dum10, Hum20, Maj13, Web21, Gat22, KK13, Sch11, Str23]. Cyberneticizing [Bou08]. Cyprus [Nic02]. Czechoslovak [CS20, Kol18]. Czechoslovakia [Bra21, Puc20].

danger [Col18]. Danish [Ped16, Syl22]. Dar [Mil22, Rob17]. David [Dai19, Dan22]. Davis [Rom18a]. dawn [Sba09]. dawning [Per17]. Day [Bat09, Dai09, GR06, Lar14]. Days [Cox15, MO10]. DC [Rob16a]. dead [McD12]. deal [Cra16]. Dealing [Byr17]. Deals [Gas24]. Dean [Kit12]. débat [Boy14]. Debate [BW07, QJ05, Ber10, Çel22, See03, Zai09]. debated [Pre12]. debates [Cor13, Ger19b]. debt [Tre20]. December [Gal19]. Déception [Die02]. Decision [Ada06, Bus01, Ano14a, Gal19, NZ12, vD12, CS20]. Declaration [Bék06, Ons05a, Buf09a]. Declassified [Ris07, Mar12, Mat15]. decline [Fre14, Lar14]. decolonisation [Jag22]. Decolonization [RB24, Sur06, Noe12]. Defeating [Ell06]. defection [LS19]. defector [Fen20, Rie19]. Defence [Die03, Dwa01, Tob02, Van01a, Bar15, Bjo23, Cho12a, Syl22, Mar22b]. defense [Woo10]. deferential [Kap19]. deferment [Rut13]. Defying [Che14]. del [Fie12, P´ 24]. della [Lib11, Wes12]. demanding [Sar21]. demands [Kos11]. demise [Tul12, Zub14]. Democracy [Man13, Mue24, Riz12, Wol10]. Democrat [Naf13]. Democratic [FM15, Fil06, Kay09, Rai01, Sco09, Cox18, Sfi14, Ass11, Gro05]. Deng [AC17, kM17]. deniable [Voß16]. Denmark [Bjo23, Far15a, Pet12a]. Dennis [Old20]. denuclearisation [Mus21]. Department [Sai21, Sto12]. dependable [Sar21]. dependence [LAGE21]. depiction [Buf09a]. deployment [Foo11, LAGE21, MV23, PP21, PP23]. Deputy [CP00]. Derek [Dro13]. Desires [Wes06]. Despite [Sea21]. destalinization [Cze13]. destroy [Win09]. destruction [Mag09]. Détente [Sur08, Alb14, Alv14, Cas15a, Dra12, Fur09, Han08, Koc08b, Lot09, Rea10, Sch08b, Sch11, Sta19, Zub08, Duj07, Pec17, WS01, Yam07, See02]. deterioration [Lou17b]. determinism [Peo08]. Deterrence [Blu01, Fre14, Kie20, Rei10]. developing [Cal19, Fri10, Roe16]. Development [Hil13b, Kap03, Nic02, Sto14, vW07, Asi24, Bor20, Boz13, Ell22b, Ger16, Ger10, MV23, PP21, PP23, Rob16b, Spe10, Sto10, Wet09, Ian12]. Devices [Wes06]. Devils [Kin24]. dialogue [Fra18]. Diasporic [Wil24]. dictator [Bel23]. dictators [Byr17, Sny21]. Dictatorship [Har15, Die11, Uli14]. Did [Tob13, Win09]. Didn’t [Goo07]. Diez [P´ 24]. Diez-Astrain [P´ 24]. Different [Dei06, Pet14, Gra20b]. Differentiation [Pec17, Che19]. difficult [Foo13]. dilemma [Dai09, Dai19, Li11, Nat18]. dimension [Kie20]. Dimona [Dai09, Ger10]. Diplomacy [BK03, Byr19, Kie20, Mag17, Mas02, Nut17, Per16, Rea06, Sto12, Sto14, Tag17, Woo13, Zha07, Aha19, AK15, Bel23, Bra16, BR23, Bro20b, Ewi19, Fog20, Gar23, Gut21, Kal08, Kel10, Lüt09, Mav21, Mik20, Pal10, Sco10, Sha10, SX10, Twi16]. diplomat [Dei13, Riz12]. diplomates [Mag22]. Diplomatic [GR06, WS01, Mat15, Pra24, Rey04]. Diplomats [Lev13a, Pac03]. director [Cul03]. diritti [Chi14]. Dirty [Har15, Sch12]. Disagreement [Lud09]. disappearance [Sal18]. disarmament [BB14]. disbandment [Che22]. disconnected [Sca12]. discourse [McM10]. discovered [Mad05]. discussion [Sch08b]. Disease [Kal14]. disoriented [Tro22]. dispatch [Gu22a]. dispute [FZ23, Riz09]. Disputes [SL15]. dissenso [Chi14, Zan12]. Dissent [Lev13a, WS01, Pit03, Rom18a, Zan12]. Distant [Ken16, Shu12]. distensione [Chi14]. Disturbing [Bro20a]. Divide [Hei17]. Divided [Dee06, Sca12, Car11, Doo21]. divorce [Zub17]. Do [Ano14a, NZ12]. Dockrill [Kom20, Ano09b]. doctrine [Sam14, Lot01]. Document [Küs02, MN01b, MN01a, Van01a]. documentary [Pra11]. Documents [Cox10, Ons07, Pon02, Ris07, Sty01, Lot10]. Dog [Goo07, Liu18]. Dog-beating [Liu18]. DOI [Ano14a]. doi.org [Ano15b, Ano15c]. Dokumente [Kra11]. domestic [Hum20, Lak21, Sca11]. domesticity [Rut13]. Don [Dei13]. Don’t [Old20, Ris09]. Doppler [Mag22]. Doppler-Speranza [Mag22]. double [Maj21]. down [Hum20, Mar23, Old20, Rhé11, Wes10, Gas18]. Drafting [Rut13]. Dragon [Aut21, Nge23]. dramas [Old20]. Drawing [D’H18]. DRC [Sch23]. dream [Ell11]. Driving [Har00]. Drogan [Ano15c]. DRV [QJ05]. dual [Ano14a, Gal19, NZ12, vD12, CS20]. dual-track [Ano14a, Gal19, NZ12, vD12, CS20]. Duke [Hil13a, Sch22]. Durham [Hil13a, Sch22]. During [Rob16a, Tub05, Ale10, Boe19, Bro20a, Che14, Cul03, Dal14, Doo21, Dra12, Dri11, Dum10, Eas15, Eli10, Ell11, FS23, Fur09, Gar13, Ger19b, Giv22, Gle16, Gno18, Her07, Hyu22, Kat24, Ked18, Knü20, KW17, Lev08, Mad05, Maj21, Mar23, Moo02, Myt11, Nat18, Oña11, Pet14, Pii09, Rub14, Rut13, Sar15, SL15, Uli14, Var03, Voß16, Zhu14]. Dutch [Van01b, Wen22]. Dynamics [Bra07, Eme10, Gno17, Loc09a].

each [Gno17]. Eagles [Cha12]. Early [Joh06, Alc18, App08, Bau16, Cav11, Cra11, Dav03, Dei13, Foo13, Gri23, hHC15, Lev13b, Mas09, Mas14, Mat15, Sar15, Syl22, Wys12, Kin24, LS24, Pec01, Rie19, Set13b, SL15]. East [Bel23, Bes12, Bur19, Byr17, Cha15, Che14, Che19, Che21, Cit19, Ful11, Gas24, Gfe11, Gil10, Gro05, Ing01, Jur17, Lot19, LK20, MM03, Maj13, Nad04, Pop10, Pop15, Pro23, Sim05, Sit18, Tan14, Tre20, Wil11]. Eastern [Dim10, Com20, Far22, FM15, Har21, Hug04, Mas09, MN01a, MN01b, Rom14a, Sta12, UT18, ZVU18]. Eating [Tun15]. EC [Ell10, Mas09, Yam07]. EC/EU [Mas09]. eccentric [Han11]. echo [Maj13]. Economic [Bur22, Ked18, Kom20, Sza05, Ber10, Ell22b, Hay18, Kie20, Pat11, Vám18, XZ19, Zof18]. economics [Ell22b]. Économie [Boy14]. économique [Boy14]. Economists [Spe10]. economy [Sar13, ZF19]. ecosystem [Dum14a]. Ecuador [Bec23]. ed [Lor24, Set13a, Sny21]. Edgar [Cul13].

Edited [Tag20]. edition [Pec14]. Editor [Buf09b]. Editorial [Ano00a, Ano01a, Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano03b, Ano10b, Ano11b, Ano12a, Ano13, Ano14b, Ano17b]. eds [Man13, Sny13b]. Eduardo [Pum22, Rob17]. educate [kM17]. Education [KW17, Jon20, Pit03]. Edward [Cox13, Cul03]. EEC [Ger19b, MD10, Par06, Rom14a]. Effects [Kho05, Rub14, SX10]. efforts [Say16]. Egypt [Ste16]. Egyptian [Ada06, Lar14, Win09]. Einaudi [Wes12]. Eisenhower [Cas19, Cox13, Gra11, McM10, Moo02, Zai09]. El’be} [dK09, dK09, Str23]. Election [Mis06]. elections [Mik20]. elite [Ger19b, Sta12]. Ellenzéki [Rot23]. Ely [Som13]. emancipation [Spo15]. embassy [Riz12, Lüt09]. Emergence [Pra02, War13, Law11, Lot09, Tul12, Wet09]. Emergency [HCL19]. emerging [Çel22]. Emerson [Ons16]. émigrés [Nie21]. Emotion [Cos00]. Empire [Bor01, Hil13a, Kan06, Puc20, Som13, Byr19, Har21, Cox15, Mil12]. enacting [Syl22]. encounter [Kim23]. encounters [Vel19]. encrypted [Mat15]. End [Ano00b, Ano01d, Ano01e, Ano01f, Ano02d, Ano02e, Ano02f, Ano03c, Boz07, BW07, Bro07, Cam08, Cox07, Gas18, Gue24, Lun00, Pat07, Ped13, Rey13, Sny13b, Zub02, Byr19, Fog20, Hil21, Hor20, Lop22, Mar22a, McP21, Min16, Ons24, Pal14, Rog12, Sal10, Shu24, Tau09, Var12, Wes17, vW24]. Endgame [Rea06, ZCP^{+}21]. ending [FH14, Sny21]. Endings [Sau24]. Enduring [Ris19]. Enemies [Tub05]. enemy [Fai15]. Energy [Coo22, Hol16, Mös23, Roe16, Ell22b, Gri23]. engagement [Pat21]. Enlai [SX10]. Enlightening [Ris07]. enmity [Do22, Fog20]. Entering [Pet14]. entre [P´ 24]. environment [How21]. Environmental [Kin13, Bra16, Fil23, LPT16, Mac11, Rob16b]. envoy [Gar23]. Epilogue [MM05]. Epoch [Aba02]. Epstein [Gra20b]. equal [Sha03b]. Equilibrium [MM05]. equivocal [Gij16]. era [Eis18, Lev13b, Vám18, Web21, ZVU18, AC17]. Erhard [Sch08a, Sch14]. Eric [Sch22]. Erratum [Ano11c, Ano15c]. España [P´ 24]. Espionage [Hug06]. Essay [Lot03]. essential [Gal19]. establishment [Whi10]. Estado [MO10]. Estados [Fie12]. Estudios [P´ 24]. État [Boy14]. Ethel [Dip18]. Ethiopia [Dai19, Tsv24]. Ethnicity [Mue13]. EU [Mas09]. Euratom [Kri15]. Eurocurrency [Kim23]. Euromissile [KC19]. euromissiles [Wet09]. Europe [Boy14, Han13, Kar21, Sny13b, Apo20, Bau16, Byr17, Com20, Dim10, Dül12, Ell06, Ell09, FM15, Har21, Hug04, Kom20, Lot09, Mar21, Mas09, Mat10, MN01a, MN01b, Mus21, Pet14, Rea10, Rey04, Rom14a, Sar10, Sca23, SS09, Sfi14, Spo15, Sta12, Tag20, Tob11b, War13, Wes12, Wes17, Woo13, ZVU18, Zub15]. European [Kec24, AC17, App08, Bör18, Boy19, Boz13, Cam08, Cox07, Day22, Die02, Die03, Dwa01, Ell11, Fur09, Ger19b, Gfe11, Gud04, Hae15, Kom20, Lod10, Lud11, Mav21, MD10, Mue13, Neh04, Pra12, Sch11, Sur08, Tre20, War13, Woo13, Zac20]. Events [Dei06]. Evidence [Fil06, GR06, Gro05, Mue06, Cav11, Kha11]. evolution [Bro12, Che22, Rie19, Sza17]. evolving [Byr19]. examining [Bid24, Min16]. example [Sch14]. exchange [Uli14]. exchanges [Smi12, Tag20, Web21]. Exclusion [Smi01]. excuse [Win09]. execution [Dip18]. Exercise [Pec01]. exhibitions [Ris11]. exhibits [Jer18]. exile [Col23, Rob17]. exiles [KK13, Tok22]. existence [SS07]. experiences [Rei16, Sca12]. Experiencing [Mey21]. expertise [How21]. Explaining [Ass12]. exploring [Mac22]. Exposes [GR06]. Extended [Blu01]. External [Myt11]. extremes [Mor14a]. exuded [Kra12]. eye [Nay18]. Eyes [Gin02, Hyu22].

F [Aut21, Cox13, Gol09, See02, Tro22, Whi11]. face [Glo18, Lea23]. Facing [Cig17]. Factor [Con05, HCL19, Hil21, Pat11]. Failure [RS17]. Faith [Pre12]. Fall [Ple13, Cox24, Do22, Fra20, Gro10, Ian12, Mil12, Zub10, vW24]. fallout [Coo22, Fil23]. false [New11]. Fardella [Ano15b]. Fascism [Har15, Mor14a]. Fascist [Ped01]. fate [Apo20]. fateful [Gup19]. Favouritism [Mar09]. FBI [Cul13]. Fear [ZM20]. February [Dra13, Arn03]. Fedayeen [Gas24, Sto12]. Federal [Cla08]. Federico [Wes12, Ano16b]. Ferguson [Han17]. Festival [Pea12]. fiction [Maj03, Man09]. fictional [Rot23]. Fidel [Pet11]. fields [LPT16]. fiery [Her12]. Fiftieth [Kad14]. fight [Hor16, Pat21]. Fighting [Gij16]. File [Fil06, Gra20b]. Filling [MD10]. Film [Cho13, Buf09a]. films [LK20, Ris09]. Final [Cox15, Sch20a, Dav09]. Finances [Boy14]. Financial [Roy02]. financière [Boy14]. Finding [Pre22, Gro15]. Fink [Lak21]. Finland [LK20, Ren00]. Finnish [Fis09, Mik20]. Finnished [GR06]. fire [Lar10, Mar16]. first [Che22, Cho14, Ped22, Rak14, Sny13a, Kra12]. Five [Kad14, Ede17, QJ11]. flag [Bag11]. flank [Mar09]. flock [Del20]. Flood [Kin24]. flying [Vel19]. FMLN [Oña11]. Focus [Rot23]. Foe [Sha03a]. Food [Tun15]. Foot} [Hug06]. Footnoting [Pet12a]. Forbidden [Rei16]. force [Liu18, Heu09]. forced [Sal18]. forces [Foo11, Ped01]. Ford [Cze13, Jon20, Tro22]. Foreign [Ces13, Hal20, Her07, Hil13b, Ing01, Lon24, Mit17, Nut02, Ons13, Pec01, Sny13c, Bor04, Cal19, Cox18, Ell11, Han11, Han14, Lak21, Lea23, Ler11, Mar23, Ped16, Per12, Pre09a, Pro23, Ren18, Sar09b, Sny12, Tau09, Töm19, Vaï12, Wol10, Jag22, Mar22b, Lop13]. forest [Rei16]. format [Ano15d]. Formation [Cos00, Mag17, Che22, Lot09]. Former [Maz03, Maz07, Muk06]. Fouchet [Van01b]. foundation [Sny10, Cze13, Jon20, Tro22]. foundational [BR23]. foundations [Gol09]. Foxbats [Dai09]. Fragmentary [RB24]. France [Boy14, Lud09, Mag19, Mag22, Mar19, Bad18, Bjo21, Boz07, Cre02, Ell09, Foo11, Foo13, Gfe11, Mis12, Per17]. Franco [Chi14, Sar15]. Francois [Mag22]. Frank [Sha17]. Franks [Hop01]. ’Fraternal [Ano05, Ano20]. Fred [Joh23]. fredda [Wes12]. Frédéric [Sny13b]. Frederik [Cox13]. Free [Gri23, Pit03, Tok22, Tro22]. Freedom [Shi20, Shu15, WD10, Cho13, Mue24]. freer [Boz13]. Freeze [Joh06]. Frei [Pum22]. Freire [Har12b]. Frelimo [Ons16, Rob17]. French [Ada21, Byr19, Ger16, Ger19a, Kon21, Lud01, MV23, Par06, PP21, PP23, Per17, Pin24, Sar15, Sar21, Vaï12]. Friedman [Lor24]. Friend [Cos00, Sha03a]. friendly [Mik20]. Friends [Cre02, Tub05, Gno17, Puc20]. friendship [Jer12, Kra12]. Frisch [Arn03].

Froide [Mar19]. front [Con10, Mar21, Sca11, Shu12]. Frontier [Cho13, DeR07, Car11]. Frontiers [Ald01, Ful11, Pri22]. frontlines [Tra21]. Frühling [Kra11]. Frustration [Pec01]. Fu [Foo09]. full [Sai21]. functioning [Jag22]. Fund [BR23, Tro22]. Future [Maj03, kM17, Ped16].

Gaddis [Cox13, Lun06]. Gaiduk [Ano11d]. Galbreath [Dan22]. Gamal [McA11]. gamble [Lar14, Tau09]. Game [Cha15, Maj21]. Garthoff [Rob16a]. gate [Eli10]. gathering [Pum22]. Gaulle [Boy14, Lud09, Glo18, Mag19, Sch14, Vaï12, Van01a]. Gaullist [Gre09]. GDR [FM15, Buf09a, Cla13, Die11]. gendarmes [Lar13]. Gendering [Tam10]. General [Ell06, Lot03, Kap19, Liu18]. generation [Ped22]. Geneva [Ass11, Kal08, Zha07]. George [Cox13, Hor20, Sha03b, Tro22]. Georges [Mar19]. Georges-Henri [Mar19]. Georgetown [Rob16a]. Georgiy [Fil06]. German [Ano15c, Che14, Ban21, Bel23, Boz07, Che19, Cox10, Day22, Del20, Dro15, FM15, Fur09, Giv22, Gro05, Ing01, Kin24, Lot10, Sal10, Sca12, Sch08a, Sme23, Spo15, Töm19, Zub14]. Germans [Lud09]. Germany [Lak21, Alb14, Bur19, Cla08, Dül08, Ell09, Fil06, Fis21, Ful11, Gup19, Hir20, Hug02, Maj13, Sit18, Spo15, Töm19]. Germanys [Tre08, Vit17]. gestures [Mik20]. ghosts [FH14]. Gilbert [Cas20]. Giving [Hug06]. Gizenga [Maz07]. glass [Shi20]. Gleamed [Gin02]. Global [Lin13, Say16, Set13a, Sur06, Ada21, Alb14, Bur19, Cha12, Cro09, Doo21, Fil23, Mac11, Mav21, Mil22, Pha12, Pre13, Pri22, Rea12, Ris19, Rub14, Sch22, Sch20b, Bor20, Mar16, Tre20, Wee21]. Globalising [Lor09]. globalization [Ell22a]. Globalizing [Vaï12]. Gluckman [Gra20b]. Go [Duj07]. Go-Between [Duj07]. goals [Han08]. Goedde [Sch20b]. Goes [Cul13, Man21, Maw13]. Gold [Ger16]. golden [Lev14, Nge23]. good [Han14, Sch22]. Goodbye [Zub15]. Gorbachev [Hil21, Bro10, Küs02, Lun00, McD12, Rea10, Rey04, Sam12, Tau09, Zub02, Zub14]. Goulart [Lou17b]. Government [Ell15, Sto10, Gil10, Lah13, Pra24, Maz07, Ull11]. governments [Fur09]. grandes [QJ09]. Grandin [Han17]. Great [Cha15, QJ09, Jer12, Lat11, Say16, Aba02, Die02, Dip18, Lot19, Rak14, Sta23, Zhu12]. greater [Mou15]. Greece [Joh06, Kar11, Riz09, Riz12, Ulu03]. Greek [Mar09, Naf13, Sfi14]. green [Bra16, Bra12, Kha17]. Greg [Han17]. grew [Fis09]. Grocer [MO10]. Grosser [Wes10, Ano16b]. Group [Pop15]. groups [Che22]. guerillas [Cha12]. guerra [Wes12]. Guerre [Mar19]. Guianese [WD10]. Guinea [Eas15, Ger19a]. guns [Spe10].

H [Ano14a, Hor20, Bar15]. H-bomb [Bar15]. Half [Ren00]. Half-Adversary [Ren00]. Half-Ally [Ren00]. halt [Sch12]. Hamburg [Str23]. Hamilton [Man13]. Hammarskjöld [O’M12, Rie14]. hand [McD12, Rhé11]. Hanoi [Do22, XZ19]. hardening [Woo10]. Harold [Gil10]. Harriman [Soa09]. Harvard [Cox13, Lor24, Pri22]. Haslam [Ken16]. Hating [Dip18]. hats [Ede17]. hatters [Jur17]. Havana [Shu15]. Havel [Dan22]. Haven [Cox24]. Hawk [Gas18]. head [Gar23]. health [Pag21].

Hear [Smi01]. Hearts [Mag17]. Heat [Oso07]. Heath [Ell10, Ell11]. heavens [Dit11]. Hee [RS17]. hegemonic [FH14]. hegemonism [Dra12]. Hegemony [Sau13, SS09]. Helmet [Gin02]. Helmut [Day22, Hae15, Kie20]. Help [Cre02, Gno17]. helping [Rhé11]. Helsinki [Ped13, Bad18, Dav09, Rea10, Sar10, Sch20a]. Henri [Mar19]. Henry [Han11, Per11]. Heonik [Cox13]. heritage [Far15a]. Hermann [Che19]. Heroes [Shu07]. hidden [Dal14]. hideously [Foo13]. High [Her07]. High-Spirited [Her07]. highland [Hyu22]. Hiroshima [Cas11a]. histoire [Boy14]. Historia [P´ 24]. historians [Ber10]. historic [Cal19]. Historical [Sur06]. Histories [Kin13]. historiographical [Cas20, Sau24]. Historiography [Dim08, Rom14b]. History [Ald01, Ano14a, Ano15b, Ano15c, Ano21, BBF24, Cra15, Fri17, Gle18, Han15, Ken16, Mit18a, Mit18b, Pai06, Ped13, RB24, Rom18b, Sny18, WS01, Wes06, Zub02, Arn03, Cav11, Cha14, Cox14, Cro11, Dül08, Far15a, Gra20a, Gu22a, Kar11, LPT16, Lan22, Lop22, MM03, Mös23, Neh04, Nie21, Pra11, Ris19, Rob16b, Ste18, Var08, Vel19, Sch20b, Wes12, Bor20, Kar21]. Hollywood [Cul13, Mag13, Zhu14]. Home [Cla08, Rey04]. Hong [kM10, kM17]. Hoover [Cul13]. Horn [Gas18]. hostile [vA09]. hot [DeV11, Rab09]. Hounds [Dee10]. Hour [Ada06]. House [Cox13, Gal19, Sar09a, Sny12]. https [Ano15b, Ano15c]. Hugh [Cha15]. Human [Cos23, Fra18, Man13, Rey13, Sch12, Tys23, Bau16, Cha12, Hil21, McP21, Mue24, Ren18, Sal18, Sny21, Sur08, Tul12, Ped13]. humanitarian [Gat22, Mar22a]. humanitarianism [Sch23]. Hungarian [LS24, Bék06, Bor04, Ger19b, Kec24, Rot23, Say09, Sty01, Vám18, Web13, Zhu12]. Hungary [Bor01, Bör18, Byr17, Gra10, Pit03]. hunger [Sal18]. hunts [Lib11]. Hurley [FS23, Pal10]. Hussein [All10, Bra12].

IAEA [Kri15, Roe16, vW15]. ice [Laj13]. Iceland [Ing01]. ICEM [Kec24]. idea [Mag19]. Idealist [Han17]. Ideological [Har15, Li11, QJ05, Sha03a, Bra16, Nad04, KW17]. Ideologies [Mor14a, Cit12a, Cit12b, New11]. ideology [Cox21, Dim08]. II [Hay18, Lan22, Mag13, Mat15, Pec14, Rei16, vW11]. III [Cox10]. Illegally [Sim05]. Illusions [Sit18, New11]. Ilya [Ano11d]. imaginary [Uli14]. imagination [Car11]. IMF [Con23, Con23]. immigration [Cla08]. Impact [Ben05, Low07, GU22b, Hug04, Zhu14]. imperatives [Kay09]. ’Imperial [Lun00, Mum17]. imperialism [Mar22a]. imperialists [Lus23]. implementing [Gal19]. implosion [Whi10]. imports [Wys12]. Impossible [Van01b]. impoverished [Lon24]. Imre [Sty05]. inch [Rey23]. Incident [MM05, Man09]. incipient [dK09]. Independence [Gle07, Ons05a, Rea06, Rab09, WN08]. Independent [Her07]. Index [Ano03d]. India [Boq05, Cha14, Jag22, Nay18, Pra24, Pro23, Sar21]. Indian [Bur22, Cra16, Gup19, Sar15]. Indications [Joh06]. Indies [Rab09]. indigenous [Jar18, Hum20]. Indochina [Kra12, Boq05, Che22, Oso07, Sza17]. Indonesia [Eas05, Eas15, Efi05, Ham21, Töm19]. Indonesian [Eas05, Jag22, Sim05, Spe10]. Industrial [Set13b, Mey21]. Industrialization [Sto14]. Industrie [Boy14]. inevitable [Kom20]. Inexpensive [Lon24]. influence [Sba09]. informants [Per17]. Information [Cul03, Ris07, Ris09, Sha10, Sai21]. initial [Kim16]. initiative [Fis09]. initiatives [Sto12]. Innocent [Bar17]. Inquiring [Ada06]. insecurity [McM10, Cox13]. insidious [Zhu14]. insights [Kim16]. inspection [Kri15]. Inspectors [Mös23]. Institute [vdH22]. instrument [Pet12a]. insurgencies [Sun22]. insurgency [D’H18]. insurgent [Dri11]. Integration [Yam07, Fur09, Ger19b, Gud04, Hae15, Zac20]. intellectual [Cze13]. intellectuals [Hum20, Sam12]. Intelligence [Bra21, Goo07, Gra20b, Ken16, Mor14b, Rob16a, CS20, Gra20a, Knü20, LS24, Per17]. intelligentsia [Kal11]. inter [Gus12]. inter-American [Gus12]. interactions [Cli19]. interdependence [WN08]. Interdoc [Nut14, SS07]. Interests [Ben05, Ban21]. Interference [Mik20]. Intergovernmental [Kec24]. Interior [Bla22, Pri22]. Interkit [Sza17]. International [Ano05, Ano20, Cas15a, Fri17, Pet14, Roy02, Sim05, Smi01, Alc18, Boy19, Kra11, Lan22, Mar21, Mou15, Sar13, Sfi14, Spo15, Sur08, Vaï12, Wen22, Coo22, Gri23, Hol16, Mös23, Roe16, WD10, Wid18]. internationale [Kra11, Nut14]. Internationalising [Gra20a]. Internationalism [Sny13c, Bau23, Boy19, Ger19c, Puc20, Vel19]. Interpreting [vA09, Neh04]. interregnum [Che14]. interrogating [Old20]. Intersections [Sur06]. Intervene [Ada06]. Intervening [Kal14]. Intervention [Eas05, Lop13, Kim16, Lar14, McA11, Nat18, Per19, Pro23, WD10]. interventionism [Mue24]. Interventions [Woh06, Tar12, Voß16]. Introduction [Ano07a, Lan08, NO15, Tob11b, Zau10]. Invasion [Sim05, Bra21, Bra12, Gra10, Lah13, Sta24]. invented [Cul03]. Investigating [Kho05, Smi01]. invitation [Aha19]. Involvement [Mis06, Var03]. Iran [Cha15, Alv14, Asi24, Bra12, Eme10, Her03, Nem10, Rai01]. Iranian [Cas15b, Maj21]. Iraq [For19, Tre08]. Iraqi [Sit18]. Irish [Cox21]. Iron [Gat22, KK13, Sch11, Str23]. Irreconcilable [Blu01]. Isaiah [Dei13]. Islam [Kha17]. Islamic [Kli23]. Island [Grü03, Con10, Kui00]. isle [Sfi14]. Israel [Lak21, Ger10, Hei17, Koc08a, Lev08, Vit17, Win09]. Israeli [Ada06, Aha19, Gol09, Lar10, McA11, Sto12]. issue [FZ23]. Issues [Cig17, Tob16]. Italian [Cli19, Ped01, Cal19, Cig17, Maz03, Mis06, Nut02, Pon02, Zan12]. italiana [Zan12]. Italy [Bro11, Bro12, Cal19, Far17, Lib11, Mis12, Tam10, Tob11a, Tob13, Var08, Var12, Zac20]. Ithaca [Ces13, Hof13, Mue24, Som13]. Itineraries [How21]. iv [Lak21]. ix [Boy14, Som13].

J. [Cul13]. Jackson [Mag19]. Jacobs [Ces13]. Jaguar [Cra16]. Jakarta [Töm19]. Janick [Tag20]. János [Bor04]. Japan [Hei11, Mum17, Pat21, Sai21]. Japanese [Gar23]. Jawaharlal [Kha11].

Jean [Cox13, Dwa01]. Jeremy [Lor24]. Jerusalem [Hei17]. Jimmy [All24, Fra18, Mit07, Sau17, Sny12, Tys23]. João [Lou17b]. Jodrell [Spi06]. John [Dee06, Aut21, Cox13, Fen20, Gol09, Lun06, Old20, See02, Whi11]. Johnson [Bra21, Cas15b, Dum10, Ell14, Gar13, Giv22, Gua02, Gud04, Ler11, Loc09a]. Joining [Koc08a]. joint [Per19, BK03, Goo07]. Jonathan [Ken16]. Jordan [All10]. Joseph [Cas20, Pal10]. journalism [Boe19]. journalists [Fai15]. journals [Jur17]. judgment [Lan22]. Julian [Mag19]. Julius [Dip18, Mad05]. June [Bat09, Dei06]. just [Ell22b, Win09]. Justifications [Sto03].

Kádár [Bor04]. Kai [Foo09]. Kai-shek [Foo09]. kak [Kal14]. Kanupp [Ell22b]. kapcsolatok [Sza23]. Karachi [Ell22b]. Karol [Zar05]. Katangese [Lar13]. Kavkaz [Gin02]. Kazimierz [Gno18]. Keep [Eas05]. Kekkonen [LK20]. Kennan [Cox13, Tro22]. Kennedy [Aut21, Cul03, Fre02, Gol09, Loc09a, Lot09, See02, Var03, Whi11]. Kenya [Sun20]. Kenyan [Cul18]. Kersten [Man21]. Khrushchev [Alv14, Ash09, Kap14, Lot10, Lun03]. kidnapping [Fen20]. Kielce [Pra02]. killed [O’M12]. Kim [CyJ17, RS17]. kínai [Sza23]. Kind [Hil13b]. King [All10]. Kingdom [DeV11, Hop08, Hug06]. Kippur [Spe13]. Kiran [Kar21]. Kissinger [Cha15, AK15, Han11, Han17, Per11, Per12]. Kissingerism [Tul12]. Klaus [Kar21]. knew [Mad05]. knots [Rom14a]. Knowledge [Man13]. known [Cav11]. Koch [Sny21]. Kohl [Küs02]. Komsomol [Hor16]. Konfrontasi [Gar23]. Kong [kM10, kM17]. Korea [Ano05, Ano20, AK15, Cas11b, Cho13, CyJ17, Kim23, LS19, Mat11, Per19, RS17]. Korean [Cho14, Far13, Hop01, hHC15, Kim16, Pre09b]. Koreans [Kim12]. Koreas [Kir20]. Kotkin [Zub16]. Kremlin [Bro14, Muk06, ZVU18]. Kremlinology [Jac16, UT18]. Krisenjahr [Kra11]. Kristin [Som13]. Kwon [Cox13].

L [Col12, Gra20b, Kap19, Rob16a]. labor [Rab09]. Labour [Fur09, Hug02, Sto10, UK14, Gil10]. Lakes [Lot19]. landscapes [Rei16, Rob16b]. Lane [Mag19, Zub16]. Lanzmann [Och06]. Laos [QJ09, Ces13]. last [DeV11, Kal11, MO10, Sch22, Wes12, Wet09, Cox15]. late [Aha19, Doo21, Gat22, Mac22, Mue24, Sea21, Tót20, Cos23, Far22]. later [ZCP^{+}21]. Latin [MZ22, Pag21, Arn13, Cas20, Dri11, How21, Jos19, Nat18, Nav11, Pet11, Wee21]. Latvian [Rei16]. launch [Com20]. launching [Ewi19]. Laura [Far22]. Laurent [QJ09, Boy14]. Law [Och06]. ’LBJ [You02]. lead [Ano14a, NZ12, Rot23]. lead-up [Rot23]. leader [Kap19, Whi11]. Leaders [Rob16a, Tar12, Kap14]. Leadership [SX14, MD10]. Leap [Zhu12, Zof18]. Learning [Cox14]. Lebanon [McA11, Nay18]. led [New11]. left [Çel22, Cox21, MZ22, Zan12]. Legacies [Sch08b]. Legacy [Boo16, Mas02, Gro10, Koc08b, Nie21]. Legitimacy [Och06, Kli23]. legitimising [Ped16]. Leiden [Tag20]. Lemnitzer [Kap19]. lesser [Cav11]. let [Old20]. Letelier [McP21]. letters [KK13]. l’Europa [Wes12]. leveraging [Aha19]. Lewis [Cox13]. Liberal [CP00, Man13, Jon20]. Liberalism [BBF24, Mor14a]. Liberation [Ald01, Gas24, Kan06, Nge23, Shu07, Gue10, Mil22]. Liberty [Pum22]. Libya [Dai19]. Life [ZM20, All10, Kos11, Mag19, Cox13, Rie14]. Light [Sty01, Bor04, Bra12, Mat15]. limit [Say16]. limited [LAGE21, Rei10]. Limits [Her07, Nic02, Mac11, Mar21, Nat18, Ste16, Sza17, Tob13]. Lindenberg [Bel23]. line [Con10, D’H18]. links [Her03, Mac22]. Lipsey [Rie14]. list [Say16]. listener [Cla13]. Listening [Maj13]. literacy [Har12b]. Literature [Cho13, Lin13, KK13, Mat11]. Little [Cre02, Bad18]. Living [Hil13a]. Local [Cro09, Lar13, Ass10]. lodo [Har23]. Logevall [Cox13]. logic [Tag20]. London [Ces13, Hof13, Lin13, Mag19, Sau13, Sch22, Som13, Gon19, MO10]. Long [Han17, Bur22, Dri11, For19, Fre14, Ger19b, Rom14a, How21]. Longer [Lud01, Cas15b]. Looking [Gro15, Heu14, Shi20]. looking-glass [Shi20]. Lorenzini [Bor20]. losing [Sun22]. Lost [Col12, GU22b, Som13, Ass12, Sta19]. Love [Buf09a, Ban21]. Luce [Bro12]. Lucius [Tra06]. Ludlow [Sny13b]. Lumumbist [Sun22]. Lundestad [Gad07]. Lutherans [Del20]. Lyman [Kap19]. Lyndon [Ell14, Giv22, Ler11].

M [Ano15b, Ano15c, Cox24, P´ 24]. MA [Cox13, Lor24]. machine [Bou08]. machinery [Per12]. Macmillan [Man13, Ash09, Glo18]. mad [Jur17]. Made [War13, Par12]. Mader [Mad05]. Madman [Nut17]. Madria [Fie12]. Madrid [P´ 24]. Magazines [Cig17]. Maghreb [Aba02]. Magyar [Bör18, Sza23]. Magyar-kínai [Sza23]. mail [Cla13]. mainline [Pre13]. maintenance [Gra11]. Making [Ass10, Hal20, Mil22, Naf13, Pin24, Bag11, Coo22, Dai19, Fil23, Hum20, Kar11, Lou17a, Mum17, Nav11, Pal14, Per12, Rey23, Sny12, Sta24]. malaria [Par12]. Malayan [HCL19]. Malenkov [Fil06]. Mali [Las24]. Maltz [Sba09]. man [Bro10, Cul03, Par12]. man-made [Par12]. Management [Fre02, McM10, Sta13]. Managing [Riz09]. mancato [Zan12]. Mandela [Ell16]. manifestation [Sco09]. mankind [Rom11]. mantle [Con23]. Many [You06b]. Mao [Ano15b, Cli19, Eis18, Far13, Far15b, Kim16, Li11, Min16, SX14]. Map [Ano07b]. maps [Lev13b]. Marcello [MO10]. Marcelo [Jos20]. March [Eas05, Lüt09, Roy02, Lot04]. Marco [Tag20]. Margaret [Sme23]. María [P´ 24]. Marie [Sny13b]. Marie-Pierre [Sny13b]. Marina [Tag20]. Mark [Man13]. market [Cig17, Kim23, Yam07]. markets [Her19]. Marshall [Twi16, Pai09]. Martial [Och06]. Marxists [Tót20]. Mass [Cig17, Lea23, Sta12, vD12]. Mass-market [Cig17]. Matern [Che19]. Matter [Ano00b, Ano01d, Ano01e, Ano01f, Ano02d, Ano02e, Ano02f, Ano03c, Bel15]. Mauritius [Bur22]. Maus [Dro13]. Maverick [Ghe13, Naf13]. Max [Gra20b]. Maxwell [Tra21]. May [Fis09]. Mayaguez} [MM05]. Mayibuye [Web22]. MBFR [Blu12]. McCarthyism [Dee10]. meaning [Lot10, vA09]. means [Aha19]. Mecca [Zei17]. Media [Som13]. mediation [Gar23, Uli14]. medicine [Pag21]. Mediterranean [Hir20]. Meeting [Kos11, dK09]. Meetings [Küs02, Pon02]. meets [MO10]. Megan [Pri22]. membership [Mas09]. memorandum [Arn03]. Memorial [Kom20]. Memoriam [Ano09b, Ano11d]. memories [Jer12]. memory [Col23, Far15a]. men [Dum14a]. Mende‘s [Boy14]. Mental [Lev13b]. mentality [Sha17]. Mention [Ris09]. mercenaries [Voß16]. mésentente [Lud09]. Metropolitan [Han17]. Mexican [Boo16]. Mexico [Sch22, Boo16]. Michael [Mue24, Cam08]. Michigan [Rie14]. micro [Sta19, Vel19]. micro-détente [Sta19]. micro-history [Vel19]. mid [Sny10]. mid-1980s [Sny10]. Middle [Cha15, Bra07, Cit19, Gas24, Gfe11, Pop10, Pop15, Wil11]. migrants [Com20]. Migration [Ham21, Kec24]. Mijal [Gno18]. Mikhail [Lun00]. Milano [Chi14]. militaires [Mag22]. militancy [Ped22]. militant [Kra12]. Military [Ben05, Hil13a, Hop01, Shu07, Spi06, Sto03, Die11, Eas15, Fis10, GU22b, Lan22, Law11, LAGE21, Lev08, Lod10, Mey21, Nge23, Pal14, Pat21, Tar12, XZ19]. military-political [Die11]. mills [LS24]. Mind [Hof13, Çel22, Soa09]. Mind-Set [Hof13]. Minds [Mag17]. mineral [Pri22]. Ministère [Boy14]. ministerial [Mar22b]. minority [Hyu22]. mirage [Gro15]. mirror [Cla13]. miscalculation [Che19]. missed [Gud04, Zan12, ZCP^{+}21]. Missile [Ano15b, Cam15, Far15b, Kad14, Kel22, Alv14, Peo08]. Missiles [Cam15]. Missing [Her03]. mission [Col18, FS23, Gar13, KK13, Lat11, Bus01]. missionaries [Sch23]. mistakes [Say16]. misunderstanding [Day22]. Mitterrand [Boz07, Sal10]. MLF [Bra07]. mobility [Web21]. Model [Ian12, Vám18, Zof18]. Models [Pac03]. moderate [Sto12]. Modern [Cha15, Hof13, Bag11, Bör18, Jac16, Kar11, Sta24, Cha14]. modernisation [Cit12a, Cit12b, LAGE21, Liu18, Par12]. Modernism [Vel19]. Modernization [Hil13b, Lat11]. Molotov [CP00]. Moment [Hal20, Hir20]. monde [Wes10]. Mondlane [Rob17]. Monetary [War13, Sch14]. Money [War13]. Mongin [MV23]. Mongolian [Rad03]. Monnet [Dwa01]. morality [Lou17a]. Moro [Har23, Har23]. Moroccan [Ste16]. Moscow [Ano14a, Cor13, Pea12, Alb14, Che14, Dra13, Gon19, Küs02, Lot01, NZ12, Pon02, Ren18, Sar09a, Som13]. Most [Han17, McK12]. movement [Ano14a, Boz13, Mar21, NZ12, Ste16, Tót20, Kat24, Rot23]. Movies [Cul13]. Mozambique [Ons16]. Mozgalom} [Rot23]. Mr [Bro14]. Mr. [Maw13]. much [Bad18]. Muehlenbeck [Set13a]. Mugabe [CyJ17]. multi [Pre22]. multi-archival [Pre22]. multilateral [Con23]. multinational [Pre22]. Murrow [Cul03]. musical [Her19]. musings [Sau24]. Muslim [Kli23]. Mussolini [McA11]. mustard [Fis09]. my [Gra20b, Kal14]. Myths [Kad14, Dav09, New11].

N [Sny13b, Ell15]. N-Bomb [Ell15]. nacionalista [Fie12]. Nagy [Sty05]. NAI [vdH22]. Nakhimov [Dav03]. Namibia [Gle07]. napalm [Mar16]. narrated [Rei16]. Narratives [Cam08, Cox07, Kra12]. Nashville [Set13a].

Nasser [Lar14, McA11]. Nation [Kat24, Ano15c, Dro15]. Nation-States [Kat24]. National [Gra20b, Kan06, MN01a, MN01b, Ass10, Cox18, McM10, Ole08, Rei16, Rot23]. National-Territorial [MN01a, MN01b]. Nationalism [Ano07a, Mue13, Pet11, Pop10]. nationalist [Ste16, Wil24]. Nations [Han15, Tót20, BB14, For21, Joh06, Jur17, Zau10]. Native [Tót20]. NATO [Ano14a, Bjo23, CS20, Cho12a, Ell06, Fur09, Hat13, Kap03, Kap19, Loc09a, Mar09, Mas09, NZ12, Pal14, Pet12a, Ris07, Ris09, Ris11, Ris19, Say16, Wet09, vD12]. Nature [Man13]. naval [Che21, kM10]. Navigating [Bur19]. NC [Hil13a]. NEA [Bro20b]. Near [Ken16]. necessity [Par11]. needs [Fra18]. Negotiating [Bjo23, Mar22b]. negotiation [Bur22]. negotiations [AK15, BB14, Blu12, Ghe13, Tre08]. Nehring [Ano14a]. Nehru [Cha14, Kha11]. nei [Zan12]. neighbor [Sch22]. neighbour [Rhé11]. Neighbours [Ken16]. Neither [Cli19, Bad18]. Nelson [Ell16]. Nemzeti [Rot23]. nerve [Pag21]. Netherlands [vD12]. network [Nut14, Ped13]. networks [Knü20]. ’Neutral [Moo02]. Neutrality [Bék06, Wid18, Wys12, Dal14, Ked18]. neutron [Tob16]. next [FH14, Syl22]. Nguyen [Pre22]. Niall [Han17]. Nicaraguan [Jar18, Nat18]. Nicholas [Kap14]. nightmare [Mag09]. Nikita [Kap14]. Nina [Ede17]. Nixon [Bro20a, BK03, Cha15, Ell10, Koc08a, Lud11, Mac11, Nut17, Pec17, Per12, Sar09a, Sar09b]. No [Bid24, Cas15b, Lud01, Ano14a, Cam15, Bar15]. nominal [Cho12a]. Non [Bra07, DeR07, Her07, Woo13, Asi24, Cra16, Hay18, Kir20, Mav21, Nat18]. Non-Aligned [Her07, Mav21]. Non-alignment [DeR07, Hay18, Kir20]. non-capitalist [Asi24]. non-intervention [Nat18]. Non-Proliferation [Bra07, Cra16]. Non-State [Woo13]. nonaligned [Rak14]. nonlinear [Law11]. nor [Bad18, Cli19]. Norden [Ole08]. Nordic [Ale10, vdH22]. Normalisation [Arn21, Cli19, hHC15]. Norms [Gri08]. north [Cas11b, Ano05, Ano20, CyJ17, Fra18, Kim23, LS19, Liê05, Loc09b, Per19, Sza05]. Northern [Nge23]. Nose [Hug06]. nostalgia [Han14]. Note [Ano12b, Ano15d, Web22, vdH22, Kel22, Lot04]. Notes [Doc08, Mar22b, Ons22]. November [Dei06, Gon19]. Novo [MO10]. NPT [Bra07]. nuances [UT18]. Nuclear [Ada21, Bid24, BK03, Cig17, Cra15, Die02, Ell22b, Kon21, Nut17, Pin24, WS01, ZM20, vW07, vW11, Ano15c, Arn21, Arn03, Ban21, Cas11a, Cho14, Col18, Cra16, Dai09, Dro15, Fil23, Ger10, Ghe13, Gri23, Gut21, Liu18, MV23, Pat15, PP21, PP23, Rub14, Sar15, Sar21, See03, Syl22, Twi16, Zai09, vW10, vW24, Ell22b]. Nuremberg [Lan22]. NY [Ces13].

O. [Dee06]. Obituary [Ano10c]. Objections [Van01b]. obsession [Whi09]. Occupation [Hei11]. Ocean [Bur22]. October [BK03, Eas05, Gon19, Ple13, Roy02]. Odyssey [Dee06, Kel10]. Offensive [KW17, Cze13, Mue24, Wet09]. officials [Sha17]. Ogaden [Gas18]. Oil [Par12, Pai09]. Old [Kal08, Old20]. Oliver [Hop01]. Olympic [Sar09a].

Olympics [Car11, Cor13]. Oman [DeV11]. One [Mag13, Rey23]. onset [Jag22]. OPEC [Fra20]. Operation [Hug06, Kha11, Gin02, Mar22a, Rot23]. operations [Gle16, Laj13, Lop22, Nge23, Sai21, Whi09]. Operative [Bro14]. Opinion [Cig17, Ris07, Jur17, Pre09b, Sme23]. opportunities [Koc08b]. opportunity [ZCP^{+}21]. Opposition [Rot23]. Order [Hal20, Han15, Sau13, Cra16, Glo18, Kit12, Lat11, Ris19]. Organisation [Gas24, Gra20b]. organisations [Eis18]. Organization [Mis06, Loc09b, Lüt07]. Origins [Cul13, Har15, Sau13, Sto14, Bro11, Com20, Ell16, Gil09, Jar18, Lot04, Lüt09, McA11, Pat21, Pat15, Zhu12, vA09]. Orsini [Hof13]. Orwell [Sha03b]. Osimo [Zac20]. Ostermann [Fis21]. Ostpolitik} [Sco09, Kie20, Lor09, Lot09]. Other [Cox13, FH14, Gno17]. Otto [Fen20]. our [Cre02, Moo02, Rey04, Sny21]. Out-of-area [Hat13]. Outcomes [SL15, Han08]. outset [Wen22]. outsider [Sny12]. Overcoming [Lot01]. Overlapping [Vit17]. overseas [Cox10]. Overstepping [Cav11]. Overstretch [Lun00]. overview [Ris09]. Own [Hug02, Mou15]. ox [Gar23]. Oxford [Cha15, Ken16, Sch20b, Sny13b].

P [Pal10]. Pact [Lüt07, Blu12, Del07, Pal14]. page [Gu22a]. Painter [Gad07]. Pakistan [Pro23]. Palestine [Gas24]. Palestinian [Cha12]. Palgrave [Man13]. Pamphlet [vdH22]. pan [App08, Aba02, Kat24]. Pan-Africanism [Kat24]. pan-European [App08]. Panthers [Rom18a]. Papandreou [KC19, Naf13]. Paper [Hae15, Mav21, Tre20, LS24]. Papers [Lev13a]. paradigm [Neh04, Ole08]. Paradox [Blu01, Han08]. Paradoxes [Zub16, Gij16]. Parallel [Hul12]. paramilitary [Bau23, Ger19c, Gle16]. Paris [Boy14, Lud09, Mar19, Wes10]. Park [RS17]. Part [MN01b, Blu12, Giv22]. Partial [Pai06]. participation [Dra13, Kra12]. partisan [Rei16]. partner [Cas15b, Sar21]. Partnership [Hop01, Col18]. Party [Asi24, Efi05, QJ05, Smu06, Jur17, Cas15a, Eis18, Ell16, Hug02, Pon02, Rai01, Sba09, Sca11, UK14, Zar05]. passage [Thu21]. Past [Ped16]. Patel [Kar21]. path [Asi24]. paths [Ano14a, NZ12]. Patrick [FS23]. Paulo [Har12b]. pbk [Lin13]. PCI [Cli19]. PCI-CCP [Cli19]. peace [Ano14a, Cha14, Col18, For21, Fre14, HCL19, Her12, KC19, Kul22, Mös23, NZ12, Sch20b, Ull11, Cox13, Twi16]. Peaceful [Boq05, SS07, Smi12, QJ11, Zha07]. peacefulness [For21]. peacemaker [KC19]. peaceniks [CS20]. peculiar [LS19]. Peierls [Arn03]. Pelopidas [PP23]. Penguin [Cox13, Han17]. Peninsula [Cho14]. People [Mac22, Tob16, Gra20b, Lon24, Tro22, Bur22, Lüt07, Nge23]. perceived [Dra12]. perception [Cho12a, Lib11]. Perceptions [Ulu03, Cox21, Fis10, Ger16, Hat13, Hyu22, McA11, Sfi14, Sha17]. Perestroika [Bro07, Rey13]. perfect [Maj03]. Perforating [Sch11]. Performing [Bör18]. peril [Gij16]. Perils [Gri23, Pea12]. period [Dim10, KW17]. Peripheral [Pag21, Pre13]. Periphery [Ing01, Far17]. Perrin [Wes10]. Personal [Fil06]. Personality [Zub02]. Perspective [Gue24, hHC15, Rea12, dK09, Set13a]. Perspectives [Zub02, Ass10, Byr19, Dim10, Gij16, Gri23, Pra11, Rog12, Sur08, Vaï12]. Petra [Sch20b]. petro [Mav21]. petro-diplomacy [Mav21]. Pettina [Fie12]. Philip [Set13a]. Philippe [PP23]. Philippine [Wil24]. pianist [Her19]. Piecing [RB24]. Piercing [Dum10]. Pierre [Sny13b, Wes10, Ano16b, Col18]. Piers [Sny13b]. Pigs [Pet12b]. Pinch [Las24]. Pinetown [Ons16]. Pinker [Fre14]. Pinochet [Ked18, Sal18]. Pirates [Lus23]. pitfalls [Pro23]. Pittsburgh [Dan22]. Pius [Pal10]. PLA [Kim12]. Place [Bla22]. Plan [Lou17b, Mus21, Pai09, Twi16, Van01b]. planning [GR06]. Plans [Nic02, Sch14]. Plant [Ell22b]. Plausibly [Voß16]. Playing [Lar10, Cul18]. pledge [Alv14]. Plenum [Per19]. plotting [Tok22]. Pogrom [Pra02]. Pol [Cio14]. Poland [Bjo21, Gno18, Kom20, KW17, Kul22, Las24, Maj21, Och06, Pac03, Pra02, Pra10, Tys23]. polarised [MZ22]. police [Har21]. Policies [Ris07, Las24, Pop10, Vaï12, Zhu12]. Policy [Ces13, Die03, Hal20, Her07, Hil13b, Ing01, Moo02, Mue06, Nut02, Ren00, Sny13c, Str23, Van01a, Yam07, Alb14, Ale10, Bor04, Boz13, Bra21, Cal19, Che19, Che21, Cox10, Cox18, Dal14, Ell11, Fis10, Fri10, Han11, Han14, Hug04, Hum20, KC19, Lea23, Ler11, Nem10, Ped16, Per12, Pet12a, Pre09a, Ren18, Rhé11, Rie19, Riz09, Rut13, Sco09, Sny12, Sta13, Sza17, Tau09, Töm19, Tys23, Wol10, Woo10, Sha17]. policy-making [Hum20]. policymakers [Sam14]. Polish [Cze13, Dei06, Gas24, Gno17, Liê05, Lus23, Mac22, Pra12, Wen22, Zar05]. Political [Fil23, Low07, MM05, Mis06, Naf13, Rea06, Sza05, All10, Car11, Cla08, Die11, Eis18, Ell11, Koc08a, Lib11, Nie21, Pet12a, Sar13, Sho18, Spo15, ZX15, ZF19]. políticas [P´ 24]. Politicians [Pac03]. Políticos [P´ 24]. Politics [Bar17, Blu01, Cox13, ZM20, Alc18, Boy19, Col23, Cul18, Far15a, Gil10, Gon19, Har12b, Kal08, Lak21, Mac11, Mar16, Ole08, Pha12, Pre09b, Rob17, Rom18a, Sca11, Sch23, Sch20b, Wen22, WN08]. Pompidou [Ell11]. Ponomareva [Ede17]. Poor [Kin24]. Pope [Pal10, Zar05]. popular [Bar15, Car11, Tob11b]. population [Tob16]. Populism [Har15]. Portrayal [KW17]. Portugal [MO10, Sea21]. Portuguese [Byr19, Cas15a, Mil12, Per11]. Poseidon [Che21]. possible [Arn03]. Post [Hal20, Lud01, Mas02, MM05, Och06, Ped01, RB24, Sau13, Sny13c, Apo20, App08, Che21, Hor16, Mat10, Ole08, Rey23, Sai21, Spo15, WN08]. Post-11 [Mas02]. Post-1945 [Lud01]. Post-Cold [Hal20, RB24, Mat10, Ole08, Rey23, Spo15, WN08]. Post-Fascist [Ped01]. Post-Martial [Och06]. post-Stalin [Hor16]. Post-Vietnam [MM05, Sny13c]. Post-War [Sau13, Apo20, App08, Che21, Sai21]. postcolonial [Ger19a]. Postwar [Sto14, Dim10, Joh23, Ris19, SS09]. Pot [Eas05, Cio14]. potential [Say16]. Potter [Old20]. pour [Boy14]. Power [Ano07a, Nic02, Pra02, Zub16, Ell22b, Ful11, Gud04, Gue10, Gut21, Hay18, Kar11, Lar14, Puc20, Twi16, Sta23, Pri22]. Powers [QJ09, BB14, Ewi19]. pp [Cha15, Cox13]. pp. [Boy14, Ces13, Cha15, Chi14, Cox13, Cox24, Dan22, Dro13, Fie12, Fis21, Han13, Han17, Hil13a, Hof13, Kar21, Ken16, Kra11, Lak21, Lin13, Lop13, Lor24, Lud09, Mag19, Mag22, Mar19, Mue24, Ons16, Ped13, P´ 24, Pri22, QJ09, Rie14, Rob16a, Sau13, Sch22, Sch20b, Set13a, Sny13b, Som13, Tag20, Wes10, Wes12, Zub16]. practical [Cas15a, Laj13]. practice [Boy19, Rie19, Sco10]. Practicing [Kel10]. Prager [Kra11]. pragmatism [Che19]. Prague [Kra11, Gra20a, Hug04, Lot01, Sta13]. PRC [For21, hHC15]. Pre [GR06, Syl22]. Pre-enacting [Syl22]. Pre-planning [GR06]. Preface [Buf09b]. Prelude [Kim12]. Present [Hil13b, Ped16, vW11, Hil13a]. Presidencia [P´ 24]. presidency [Giv22]. President [Koc08a, Lou17b]. presidential [LK20]. presidents [Tar12]. Press [Bor20, Ces13, Cha15, Cox13, Dan22, Dro13, Fis21, Han13, Han17, Hil13a, Hof13, Kar21, Ken16, Lak21, Lop13, Lor24, Mue24, Ped13, Pri22, Rie14, Rob16a, Sch22, Sch20b, Set13a, Som13, Bar15, FM15, Lus23, Rom18a, UT18]. Presses [Lud09, Mag22, QJ09]. pressure [Gra11]. Pretoria [Shu15]. Preventing [Ull11]. Prevention [Bid24]. preventive [Lod10]. Price [Hop01]. Priest [Bla22]. Prime [Som13]. Princeton [Bor20]. principles [QJ11]. prisoners [Del20, Uli14]. privilege [Sha03b]. Prize [Hae15, Mav21, Tre20, Kom20]. PRL [Mac22]. problem [Doo21, Kri15, McK12]. process [Sar09a]. produced [Ris09]. Production [Man13]. Professors [Tsv24]. program [Par12]. programme [Ada21, Pat15]. programmes [Rob16b]. Progress [Lou17b]. Project [Ell22b, Ger16, Kol18, Kar21]. Projecting [Che21]. proletarian [Lüt09]. Proletariat [Smu06]. Proliferation [Bra07, Cra16]. promises [Gri23]. Promoting [Gno18]. promotion [Mue24]. Propaganda [Pec01, Ris11, Foo13, Gon19, Pre09b, Sai21, Sha10, Tob11b]. Proper [Bla22]. property [Tob16]. prospection [Ada21]. Prospects [Yeg06, Mas09]. protection [Bar15]. protest [Rub14]. Protestant [Sch23]. Protestants [Pre13]. Provincial [Kra12]. Psychological [SS07, Bro12]. psychosis [vD12]. Public [Cig17, Mag17, Ris07, Tag17, BR23, Bro20b, Kel10, Pre09b, Sha10, Sme23]. publics [Boy14]. Pudovkin [Dav03]. puissances [QJ09]. purges [Sta12]. purpose [MV23, PP21, PP23]. Putin [Bro14]. Pyongyang [Do22]. Pyrrhic [Zac20].

quarter [Heu14]. Quemoy [Con10]. Quest [Kan06, vW11]. Question [Hei17, Liê05, Ons05a, Ano15c, Dro15, For21, FH14, Lot10, Mus21, Pii09, Töm19].

R [Cul03]. Race [Sau17, Cas11a, McD12]. Racial [DeR07]. Radical [Kha17, Old20]. Radicals [All24]. Radio [Spi06, Cla13, KK13, Web13, Ris13]. ragion [Har23]. RAI [Tob13]. raise [Web21]. Ramos [P´ 24]. Random [Cox13]. Rapacki [Mus21]. Rapprochement [Bur01, Con05, Kho05, Kui00, Tag17, Aha19]. rationality [Dal14]. Raymond [Rob16a]. RDA [P´ 24]. Re [Bid24, hHC15, Min16]. Re-examining [Bid24, Min16]. Re-thinking [hHC15]. Reach [Han17, Ada21, ZVU18]. reaction [Pet11]. reactions [Bar15, Cam15]. reactor [Ano15c, Dro15]. Reader [Sha17, Pra11]. Readiness [BK03]. Reading [Kar16, Rot23]. Reagan [Hil21, Dob12, McD12, Sam12]. real [Uli14]. Realism [Kho05]. Realist [Han11]. realists [Bor04]. Reality [Dro13, New11, Sfi14]. really [Fog20, Han14]. Realpolitik} [Rea06]. Reappraisal [Boz07]. rearmament [Pat21]. reason [Har23, Kos11]. reassessing [MV23, PP21, PP23]. Reassessment [XZ19, Sny10]. rebellion [Dob12, Jar18]. Rebellious [Pra10]. rebirth [Ell10]. Received [Ano15a, Ano16a, Ano17a, Ano18, Ano09a, Ano10a, Ano11a]. Recently [Ris07]. reception [Pit03, Tob11b]. recognise [Pra24]. recognition [Gup19, Pra24]. reconciliation [Cra11, Spe13, Zub17]. Reconciling [Blu01]. Reconnecting [Str23]. reconsidered [Sco09]. reconsidering [Per19]. Reconstruction [Ano05, Ano20, Ped01, Del20, SS09]. record [Dav09]. recovery [App08]. Recrimination [Spe13]. recurrent [Mag09]. Red [Hof13, Wid18, Bag11, Gij16, Knü20, Mum17, Jar18, Oña11, RS17, Tam10, Tót20]. Rediscovering [App08]. Reduction [Bid24, Tob02]. Redux [Kho05]. Reflected [Maz07]. Reflection [MM05]. Reflections [Han15, Pat07]. Réforme [Boy14]. reforms [Vám18]. refuge [See03]. Refugee [Kec24, Mar22a]. refugees [Cha12, Eli10, Kec24, Tro22]. refuse [Cra16]. Regarding [Hei17]. regime [Bai18, Cio14]. regimes [Zan12]. regimi [Zan12]. region [Ale10, Kra12, Lot19]. regional [Rog12, Sta19, Wil11]. rejection [Sch14]. relaciones [P´ 24]. relation [Riz09]. Relations [AC17, Bur01, Ell06, Far13, Gas24, Kra14, Lou17b, Mar17, Mit17, Ons13, Sha03a, Smi01, Sty05, Aha19, Bor04, Bro20a, Cav11, Che14, Con23, Cul18, Day22, Eis18, Far17, Fur09, Giv22, Gno17, Jer18, Jun11, Lak21, Lud11, Mar23, kM10, McK12, Min16, Oña11, Par11, Sar09b, Sar15, Sch11, Spe10, Spe13, Sun20, TL23, Vám18, Vel19, vW10, vW15, Cli19]. relationship [Koc08a, McA11, Pop10, ZF19, Ell10, UK14]. relationships [Wen22]. Relief [Kin24, Gat22]. Religion [Pre09a, Set13a]. Religious [Hof13, Set13b]. relinquished [Sba09]. remake [Sca23]. Remaking [Mit17, Jos19]. Renamo [Ons16]. renegade [Rom18a]. renegotiation [Sar21]. repatriate [Mum17]. repatriation [Kim12]. Report [Mel07, Lud01, Nut02]. reporting [Mat15, UT18]. representation [For21, dK09]. representations [Foo11]. repression [Lou17a, Sta12]. Republic [Ass11, Bur22, Cla08, FM15, Fil06, Gro05, Lüt07, Mac22, Aut21, Ons05b]. Republican [Cox21, Sca11]. request [Coo22]. requirement [Laj13]. Research [Kel22, Lud01, Mar22b, Nut02, Ons22, Sai21, Web22, vdH22, Doo21, Hop08, Myt11, Pre22]. Researching [Koc08b]. Resolution [Nic02]. Response [Gad07, Hei17, MV23, Par06, PP23, Yam07, vW07, Ano16b, Bla22, Jos20, Lah13]. Responses [Dei06, Cho12a, Peo08]. result [Die11]. Resurgence [Çel22]. Retconning [Lop22]. Rethinking [FZ23, Ger19a]. RETRACTED [Ano05]. Retraction [Ano20]. retrospect [Mas14]. retrospective [Col23]. return [New11]. returned [Tan14]. Reunification [RS17, Sch08a, Sme23, Zub14]. Revelations [Muk06]. revenants [Mum17]. Review [Ald01, Bor20, Boy14, Ces13, Chi14, Cox24, Dan22, Dro13, Fie12, Fis21, Gua02, Han13, Hil13a, Hil21, Hof13, Kar21, Ken16, Kra11, Lak21, Lin13, Lop13, Lor24, Lot03, Lud09, Mag19, Mag22, Man13, Mar19, McM06, Mue24, Ons16, Ped13, P´ 24, Pop15, Pri22, QJ09, Rie14, Rob16a, Sau13, Sch22, Sch20b, Set13a, Sny13b, Som13, Tag20, Wes10, Wes12, Zub16]. Reviews [Ash07, Aun06, Blo05, Cha15, Cox13, Fal05, Han17, HB06, Lat09, Rey07, Ruo05, Tho05, You06a, Ano03a, Ano04, Cro08, Ken08, Sch01, Sch02a, Sch02b, Sch03]. Revisionism’ [Gro05]. revisionist [Ass11, Mar21]. revisited [Che19, Peo08, Sba09, UT18]. Revisiting [HCL19, Hei11, Tob02]. Revival [Efi05]. Revolution [Hil13b, Pum22, WS01, Zei17, Col23, Dri11, Gra10, Ham21, Kar16, Lor24, Min16, Nat18, Pal14, Per11, SX14, Sto12, Thu21, Zhu12, Bék06, Bjo21, Boo16, Cas15a, Cha12, Mar17, Ple13, Dan22]. Revolutionaries [Bau23, Ger19c]. Revolutionary [Mil22, Gro10, Han08, Nav11, Ped22, Ste16]. Reward [Bid24]. Rey [Sny13b]. Rhetoric [Rea06, For21, Kay09, Soa09]. rhetorical [Gri08]. Rhodesia [Mar22a, Mit07, Ons05a]. Rhodesian [Low07]. Rican [Bro20a]. Richter [Her19]. rift [Glo18]. Right [Hil13b, Bau23, Ger19c]. Rights [Cos23, Ped13, Rey13, Bau16, Cha12, Fra18, McP21, Mue24, Ren18, Sal18, Sch12, Sny21, Sur08, Tul12, Tys23]. Ripe [Lor24]. Rise [Hal20, Kha17, Sau13, Fra20, Ian12, Zub10]. rising [Gre09]. Risk [Bid24]. rivalries [Vit17]. rivalry [Sta23]. River [Jur17]. roads [Rak14]. Robert [CyJ17]. Roberts [Sha17]. rock [Bel23]. Roger [Rie14]. Rogge [Dee06]. Roham [Cha15]. ROK [hHC15]. Role [Ing01, SS07, Sty01, Woo13, Giv22, Ped16, Tob13, Zof18]. rollback [Fis21]. Rolling [Ald01]. Romania [Del07, Dra12, Dra13, Gat22, Ghe13, Gra10, Mav21, Sta13]. Romanian [Sta19]. Rome [Pon02]. Romero [Wes12, Ano16b]. Ronald [Dob12]. Rosenberg [Dip18]. Rostislav [Asi24]. Roth [Som13]. Roth-Ely [Som13]. Rother [Sny13b]. Round [Ons24]. Routledge [Lin13]. Rule [Hei17]. runaway [LS19]. Running [Dee10]. Russia [Tro22, Bes12, Ple13, Rey23, Tok22]. Russian [Dro13, Gin02, Jer12, Kal11, Maz07, Yeg06]. Russians [Fog20, Hug06]. Ruth [Ano09b, Kom20].

Sadat [Dai19]. Saddam [Bra12]. Sadler [Gon19]. Safeguarding [Cox18]. Saharan [Coo22]. Saigon [Do22]. Saints [Tam10]. Sake [Hug02]. Sakharov [Kos11]. Saki [Ano09b, Kom20]. Salaam [Mil22, Rob17]. Salt [Las24]. Salvador [Sto03, D’H18]. Salvadoran [Bel15, Oña11]. same [Giv22]. sanctions [Eme10]. Sandinista [Nat18]. Sandinistas [Jar18]. Sands [Gas18]. Sang [LS19]. Sang-Cho [LS19]. Santiago [TL23]. Sara [Bor20]. Sarah [Ped13]. satellite [Pra10]. Satire [Dro13]. Saudi [Par12]. Saving [Par06]. sceptical [Sny13a]. Schaufelbuehl [Tag20]. Schmidli [Mue24]. Schmidt [Day22, Hae15, Kie20]. Schoenborn [Lud09]. scholarship [Har08].

Schönau [Cha12]. School [Hae15, Mav21, Tre20]. Schriever [Her12]. Schwarz [Joh23]. Science [Man13, Per16, Spi06, ZM20, Bou08, Dim08, How21, Jac16, Jon20, Law11, Maj03, Rub14, SS09, Zai09]. sciences [Cze13]. Scowcroft [Sny13c]. Screen [Mag13]. sea [Sfi14]. Search [Och06, Bai18, Kli23, Nem10]. Seattle [Soa09]. Second [Par06, Dra12, Dum14b, Sca11, Glo18, Ull11]. Secondary [KW17, Mue06]. Secret [BK03, Cha15, Gas24, Nut17, Oso07, Har21, Mad05, Mey21, Pal10, ZX15]. secrets [Bro20a]. Secured [Cre02]. Securing [Lou17a]. Security [Die03, Gra20b, Har21, ZM20, Boz13, Cox18, FZ23, Web21]. seed [Fis09]. Seeking [Hug02]. seemed [Hor20]. Seen [Lun03]. selection [Sar09a]. self [Kri15]. self-inspection [Kri15]. Selling [BR23, Pre09b]. Selma [Ren18]. semi [Sau24]. semi-historiographical [Sau24]. send [Bus01]. sense [Ass10]. September [Mas02]. Sera [Lib11]. Serbian [Cav11]. series [Cox10]. serious [McK12]. Service [Ris09]. services [Mad05, Myt11]. Set [Hof13, Asi24]. Seth [Ces13]. setting [Dav09]. Settlement [Boq05]. Seven [Gra10]. sex [Lou17a]. Shaba [Lar13]. Shadow [Han17, Kat24, Lüt17, For19]. shadows [Pet11]. Shah [Cha15, Alv14]. Shaka [Mac22]. Shanghai [Mey21]. shape [Tar12]. Shaping [Cha15, Ell09, Han11, Ped22, Spo15]. Shattering [Shi20]. shed [Mat15]. sheep [Fai15]. shek [Foo09]. Shifting [ZX15, Che21]. Shoah [Och06]. shock [Rea12]. Shop [Lud01]. shoplifting [Ede17]. short [Ris09, Zau10]. should [Gno17]. shuttle [SX10]. Side [Moo02, Old20]. Sideshow [Pop15]. Significance [QJ05]. significant [Far17]. Silence [Smi01]. silent [Pat21, Sch14]. Since [Zub15, Liu18]. sinistra [Zan12]. Sino [Tag20, AC17, AK15, Bur01, Cli19, Con05, Dit11, Fri17, Gno17, Gno18, Jer18, Kho05, Kol18, Kra14, Kui00, Li11, Lüt17, Min16, Pat11, Per19, Shi09, Vám18]. Sir [Hop01, Sha17]. sites [Far15a]. Six [Bat09, Dai09, GR06, Lar14]. Six-Day [Lar14]. sixties [Dum14b, Sch22]. Sizwe [Ell16]. skill [Peo08]. Skoda [Smu06]. Small [Tag17]. Smith [Cox13, Maw13]. Snyder [Ped13]. Social [Man13, Sur06, WS01, Cze13, Jon20, Sco09]. Socialism [AC17, Bur19, Çel22, Lor24, Ano05, Ano20]. Socialist [Las24, Bör18, Boy19, Dum14b, FM15, Puc20, Sit18, TL23, Tre20, Cas15a, Gno17, Jer18]. socialists [Boy19]. society [Die11, Whi10]. socio [MM03]. socio-cultural [MM03]. soil [Jag22]. Soldiers [RB24]. Solidarity [Las24, Pra12, Sza17, Ste16, Pra12]. Solovey [Man13]. Some [Pat07, Sha03b]. Soul [Dee06, Pre09a, Rom11]. Sources [Liê05, Nut02, Pat11]. South [Dro13, Kon21, Ons05b, Ons16, Shi20, vW10, Cho13, Ell16, Fra18, Gue24, hHC15, Mac22, Mar22b, Mil12, Ons05a, Sea21, Shu24, Tre20, vW07, vW15]. Southeast [Ben05, MM05]. Southeastern [Mar09]. Southern [Har12a, Ano07a, Ano07b, Gue10, Mar22a, Ons13, Ons22, Shu07, Shu15, Web22, vW24]. Soutou [Mar19]. sovereignty [Apo20, McP21, Tót20]. Soviet [Bid24, Cox24, Gno18, Ken16, Kra14, Sha17, Som13, Tag20, Ada06, Alv14, Asi24, Bar17, Bat09, Boe19, Bor01, Bra16, Bra21, Bur01, Che19, Cox14, Cox15, Cra11, Dav03, Dit11, Dum14a, Eas15, Fai15, Fri10, Fri17, Gra10, Gri08, Her19, Heu09, Hol16, Hor16, Hug06, Ian12, Jer18, Kal08, Kan06, Kap14, Kat24, Kui00, Lah13, Lan22, Lar10, Li11, Lun03, Lüt09, Lüt17, Mat15, Maz07, Mik20, Mue06, Muk06, Mum17, Pec01, Pec14, Ped22, Peo08, Per19, Pon02, Puc20, Ren00, Rob16a, Rom18a, Rom11, Rub14, See02, Sha03a, Shi09, Shu07, Smi12, Sta24, Sto03, Sty01, Sty05, Sza17, Tag17, Tau09, Tro22, Tub05, Uli14, Ulu03, Wet09, Wol10, Yam07, ZX15, Zhu14, Zub08, Zub14, Zub17, ZCP^{+}21, dK09]. Sovietization [Dim10]. Soviets [Dai09, Har00, Ris09, Win09]. Spain [Far22, Joh06, LAGE21]. Spanish [Gar13]. Special [Ons22, Koc08a, McA11, Sea21, WN08, Ell10, UK14]. Spectre [Nut17]. spectres [FH14]. speech [Soa09]. Speranza [Mag22]. spheres [Hul12]. Spies [CS20]. Spirited [Her07]. Spiritual [Set13b]. Split [Cam08, Cox07, Do22, Li11, Pat11, Shi09, Gno18, Tag20]. spoon [Bur22]. sport [Ede17]. Spring [Ban05, Kra11, Gra20a, Hug04, Sta13]. Sputnik} [Peo08]. spy [Lop22]. stability [Nem10]. Staff [BK03]. Stage [Pop15]. stalemate [Rey23]. Stalin [Apo20, Bro11, Cho12b, CP00, Efi05, Far13, Hor16, Kim10, Lot04, Lot10, Mue06, MN01a, MN01b, Rai01, Thu21, Wol10, Zub16]. Stalinist [KW17, Sta12]. Stanford [Fis21]. star [Bel23]. Stasi [Sit18]. State [Mil22, Woo13, Çel22, Die11, Har23, McP21, Sha03b, Sto12, Sit18, Sto12, ZF19]. State-Making [Mil22]. Statecraft [Cra15, Min16]. Statement [Ano20, Bau16]. States [Cha15, Cul03, Fis21, Jun11, Kat24, Sha10, vW10, FM15, Boo16, Bra12, Byr17, Cas11a, Cit12a, Cit12b, Eli10, Ell09, For19, Gar13, Ger10, Gre09, Har08, Hum20, Jon20, Kar11, Kay09, Ked18, Kha17, KW17, Ler11, Lev08, Lus23, Mis12, Ons13, Pre09b, Rab09, Rak14, Riz09, Rom11, Rub14, Sim05, vW07]. Statesman [Han17]. stato [Har23]. Status [Cos23]. steering [FM15]. Stephen [Fre14, Ons16, Zub16]. Steppe [Kra14]. Steps [Bur01, Tag17]. stick [Liu18]. Storia [Wes12]. Stories [Mar21]. storm [Nay18, Pum22]. Story [Far22, Zub15, MZ22, McD12, Pec14, Spo15]. straight [Dav09]. strained [Lud11]. Strait [Jun11]. Strange [McP21, Fen20]. stranger [Man09]. Strasbourg [Mag22]. Strategic [Ben05, Kay09, Cha14, CyJ17, Ell09, Kim16, Kon21, LAGE21, Pal14]. Strategy [Cra15, Hus05, Mis06, Pin24, Blu12, Fur09, Ger19a, Liu18, Lod10, Maj21]. Strength [Hug02]. strike [Sal18, WD10]. Striking [WD10]. Struggle [Ons16, Rea06, Shu15, Sto14, Apo20, Dit11, Giv22, Hay18, Mar21, Mat10, Sta23]. struggles [Rab09]. Student [Kat24]. Students [Tsv24, Bur19, Gar13, Ped22, Sit18, Pea12]. Studies [Yeg06, Ole08, WN08, Jos19]. Study [Nut02, Ris07, FM15, Rey04, Sco10]. sublime [Rei16]. submarine [Pat21]. Submissions [Ano21]. Suburbanization [Tob02]. Subversion [Lon24]. Subversive [Dro13]. Suez [Sto14]. sugar [WD10]. Summer [Hae15, Mav21, Tre20]. summits [Sam12]. Sung [CyJ17, RS17]. suns [Dit11]. Sunset [vW10]. Super [Har00]. Super-Ally [Har00]. Superpower [Aha19, Har00, Kan06, Mit17, New11]. Superpowers [Bra07, Woh06, Gfe11, Wil11]. Supping [Bur22]. supplement} [Cas20]. Supplied [Sun22]. Support [Kan06, Bel15, Cas15a, Eas15, Sny21, Sun22]. Supporting [Bus01]. supremacy [Dit11, O’M12, vW11]. Surveillance [Far22, Web21]. survey [Mat11]. Survival [Tun15, Bai18, Mus21]. Sustaining [Grü03]. Sviatoslav [Her19]. Sweden [Dal14, Mat15, Moo02]. Swedish [Sco09]. Swiss [Wys12]. Switzerland [Knü20, WN08, Wid18, ZM20]. symbolic [CyJ17]. Syria [Tre08]. Syrian [Lar10]. System [War13]. systems [Die11, Law11].

Table [Ons24]. Tactics [Hei17]. tail [Gar23]. Taiwan [Jun11, Lus23, Sha03a, Tub05, ZX15]. takes [Pra24]. Taking [Las24]. talk [Jun11]. talks [HCL19, ZX15]. Tallandier [Mar19]. tamed [Lud09]. Taming [Aut21]. Tanks [Tra06]. Tanzania [Rob17]. Tanzanian [Bur19]. Targeting [Ing01]. Taunting [Del07]. Tauris [Sau13]. Taylor [Tra21]. tea [Jur17]. teaches [Sta23]. Teaching [Har08, Hop08, Joh23]. Tearing [Hum20]. Technical [Bur22]. techno [Gut21]. techno-diplomacy [Gut21]. technological [Peo08]. Technology [Spi06, Ghe13, How21]. Telescope [Spi06]. television [Old20, Zhu14]. tell [Mas14]. Telón [P´ 24]. Tending [Del20]. tentative [Dum10]. tenure [Cul03]. Territorial [MN01a, MN01b, FZ23]. terror [Sta12, Bro20b, Lop13]. terrorism [Aha19, Har23]. terrorismo [Har23]. Terrorists [Hof13]. Test [BK03, Cas15a, Zai09]. test-ban [Zai09]. testimony [Riz12]. testing [Fil23]. Textbooks [KW17]. Thai [Hyu22]. Thailand [Oso07]. Thatcher [Hil21, Lah13, kM17, Sme23]. Theatre [Mue06]. their [Che22, Mou15]. them [Sar15, Kha17]. themes [Boe19]. theory [Heu09]. There [Hil13a, Bra12]. these [Tro22]. theses [Boe19]. Thieu [Pre22]. Things [Mil12]. think [Cra16]. thinkers [Sai21]. thinking [hHC15, Peo08]. Third [Lor24, Win09, Ewi19, Hat13, Hor16, Las24, Lüt17, Mag17, Rog12, Sto14, Woh06]. thirty [ZCP^{+}21]. Thompson [Bus01]. thought [Kos11, Sho18]. Threat [Sto03, Cho12a, Dra12, Peo08, Kha17]. Threatened [For21]. Three [You06b, Sch08b, Lou17b, Pec14]. three-way [Sch08b]. Three-Year [Lou17b]. thwart [CS20]. tie [Sea21]. Ties [Gol09]. Time [Som13, Ham21]. Timing [Ons05a]. Timor [Sim05]. Tito [Hay18, Nie21]. today [Sta23]. Together [RB24]. Togliatti [Bro11]. tolerance [Çel22]. too [Bad18, Mas14]. totalitarian [Bau16]. Tough [Tre08]. tourists [kM10]. Town [Tun15]. Tra [Chi14]. track [Ano14a, Gal19, NZ12, vD12, CS20]. Trade [Gri23, Pra12]. traitor [Old20]. trans [KK13]. transaction [Ger19a]. Transatlantic [Cam08, Cox07, Han13, Col18, Eme10, Gre09, Gud04, Sch11]. transfer [SX14]. transformation [GU22b, Sar10, Sch20a]. Transformed [Mit17, Ren18]. translation [GU22b]. Transnational [Ped13, Ass10, Bel15, Lar13, Mou15]. transport [Kec24]. transsystemic [Gut21]. travelling [Ris11]. travels [Nav11]. través [P´ 24]. tread [Pro23]. Treaty [Loc09b]. tree [Fis09]. triangle [Ell09, Lar10, Nge23]. Trías [MZ22]. tribunal [Lan22]. Trident [Che21]. Trilateral [Far13]. trip [Che19]. Tropes [Mit07]. Troublemaker [KC19]. Truce [See02]. Trudeau [Col18]. truth [Cul03, Man09]. Tudeh [Asi24]. Tulli [Chi14]. Tunisian [Coo22]. Turin [Wes12]. Turkey [Çel22, Kay09, Riz09, Ulu03]. Turkish [Aha19, Cho12b, GU22b]. turned [Wes10]. tutelage [Sar15]. twentieth [Bro10, Fra20, Sca23, Har15, Sco10]. twentieth-century [Sca23, Sco10]. twilight [Sta23]. twinning [Kol18]. twins [Kon21]. twisted [Tro22]. Two [Dit11, Hil13a, Rak14, Sau24, You06b, Gra20b, Kir20, Tre08, Vit17]. Tyler [Bla22].

U.S [Hil13a, Rei10]. U.S. [Hal20, Hil13b, Lou17b, Per12, Ren18, Whi09]. Udo [Bel23]. Ukraine [ECH^{+}23]. Ulianovskii [Asi24]. ultimate [Her12]. ultimo [Wes12]. umani [Chi14]. Umberto [Chi14]. Umkhonto [Ell16]. Un-Finnished [GR06]. Uncertain [Sha03a]. uncertainty [App08]. Uncivil [Whi10]. under-ice [Laj13]. understanding [Pra24]. Undesired [SL15]. Uneasy [See02]. unfinished [Dan22]. Unfit [PP21, MV23, PP23]. unfortunate [Fen20]. Unidos [Fie12]. Unification [Boz07, Cox10, Sal10, Spo15]. Unilateral [Ons05a]. Union [Som13, Van01b, Sch14, Alv14, Bat09, Bor01, Cox15, Cra11, Hol16, Kat24, Lüt09, Mum17, Ped22, Pon02, Rom11, Rub14, Smi12, Tro22, Tub05, Uli14, Zhu14, Zub08, Zub17, Cox24]. unions [Pra12, Gri23]. Unipolar [Hal20]. unique [Sfi14]. United [Cha15, Fis21, Zar05, BB14, Boo16, Bra12, Byr17, Cas11a, Cit12a, Cit12b, Cul03, DeV11, Eli10, Ell09, For19, For21, Gar13, Ger10, Gre09, Har08, Hop08, Hug06, Hum20, Joh06, Jon20, Jun11, Jur17, Kar11, Kay09, Ked18, Kha17, KW17, Ler11, Lev08, Lus23, Mis12, Ons13, Pre09b, Rab09, Rak14, Riz09, Rom11, Rub14, Sha10, Sim05, Zau10, vW07, vW10]. Unity [Kom20, Ell11]. Universe [Ces13]. Universitaires [Lud09, Mag22]. Université [QJ09]. University [Bor20, Ces13, Cha15, Cox13, Cox24, Dan22, Dro13, Fis21, Han13, Hil13a, Hof13, Kar21, Ken16, Lak21, Lop13, Lor24, Mue24, Ped13, Pri22, Rie14, Rob16a, Sch22, Sch20b, Set13a, Som13, Dee10, Gra20b]. Unknown [Gu22a]. Unmasking [Fai15]. Unraveling [Ces13]. Unsung [Shu07]. untold [McD12]. Untying [Rom14a]. Unvarnishing [Dro13]. Uprising [Rot23, Web13, Muk06]. upside [Wes10]. uranium [Ada21]. Urban [Tob02]. Urho [LK20]. Urss [Chi14]. Uruguay [Sny21]. USA [DeR07]. Uses [Spi06, Wes06, Gut21]. USSR [Bad18, GR06, Gu22a, Kli23, Mat15, Maz03, Shu12, Sta13].

V [Ano11d, Kal14]. vacuum [MD10]. Vaïsse [MV23]. Valeria [Tag20]. valuable [Lon24]. value [Pre22]. values [Lib11]. Vanderbilt [Set13a]. vanishing [Jac16]. Vanni [Fie12]. Vatican [Pal10]. Vatican’s [Sca23]. Verlag [Kra11]. Versus [Kha17, Sar09a, Wen22]. very [Laj13, Lon24, Nay18, Zau10]. veterans [Alc18, LS24]. Victim [Fen20]. Victims [Cos23, Mum17]. victory [Zac20]. Vienna [Maw13, Sny10]. Vietnam [Pre12, Ass10, Ass11, BK03, Bus01, Col12, Con05, Gro05, Hun11, Liê05, kM10, Mar12, MM05, Nut17, Pra11, Rhé11, Sca11, Sco09, Sny13c, Ste18, Sza05, Var03]. Vietnamese [Pat11, Pra11]. Vietnams [You06b]. view [Law11, Myt11, Sta19]. viewed [Pec14]. Views [Lot03]. vii [Lak21, Sny13b, Cox10]. violence [Cla08, Dri11]. virtually [Bes12]. Virus [Kal14]. vis [Çel22]. vis-à-vis [Çel22]. Visions [Shu15, Pre13, Sny13b]. Visit [Kap14, Kha11, MO10, kM17, Pec17]. Visits [BR23, kM10, Pac03]. visual [Syl22, dK09]. Vivian [MZ22]. Vladislav [Cox24]. VOA [Tob11a]. voice [Gfe11, Pit03]. Voices [KK13, Lev13a, Mag17]. Vol [Ons13, Zub16, Han17, Kra11]. Volk [Bai18]. vols [Kra11]. Volume [Ano03d, Cox10]. Vstrecha [dK09]. vtorzheniem [Kal14]. Vulnerability [Tob02].

W [Hor20]. wages [Sar09a]. Waging [Com20]. Waiting [Ger10]. wall [vW24, Buf09a, Ful11, Gro10, Har00, Ple13, Zub14]. War [Ada06, Alc18, Ano14a, Ano15b, Ano15c, Bec23, Bor20, Bur01, Bus01, Ces13, Cha14, Cox13, Dro13, Fis21, Hae15, Har15, Hay18, Hil13a, Jos19, Lak21, Lan22, Lib11, Lin13, Lop13, Lou17b, Mag13, Man13, Mat15, Nut14, Ole08, Oso07, Pag21, Pec14, Ped13, Puc20, Rei10, Rei16, Sau13, Sch23, Sch20b, Set13a, Sny13b, Tag20, WN08, You02, vW11, Apo20, App08, Ass12, Bes12, Bou08, Cas11b, Che21, DeV11, Del20, ECH^{+}23, Far15a, FS23, Gle16, Han14, Her03, Hun11, Jar18, Kel22, Kim12, Kul22, Lar13, Law11, Lev08, Man21, Mat10, Mat11, Pre12, Sai21, Sca11, Sch12, Sha03b, Ste18, Syl22, Tar12, Uli14, Wid18, vW24, Aba02, Aha19, Ahl13, Alc18, Ald01, Ale10, Ano07a, Ano21, Arn21, Arn13, Bar17, Bat09, Bau16, Bau23]. War [Ben05, BBF24, Bes12, Bid24, Boe19, Boo16, Boq05, Bou08, Boz07, Bra16, Bra07, BR23, Bro11, BW07, Bro07, Bro20b, BK03, Cam08, Car11, Cas20, Çel22, Cha15, Che22, Cho13, Cho12b, Cit12a, Cit12b, Cit19, Col12, Com20, Con05, Con10, Cor13, Cos23, Cos00, Cox07, Cox13, Cox14, Cox21, Cra11, Cro09, Cro11, Cul13, Cul18, Dai09, Dal14, Dav03, DeV11, Dee06, Dei13, Die11, Dim08, Doo21, Dri11, Dül08, Dum14a, Ede17, Eli10, Ell22a, Ell11, Eme10, Far22, Far15a, Far13, Far17, Fis10, Fog20, Foo13, For19, FH14, Gas18, Ger19c, Gil09, Gil10, Gin02, GR06, Gle18, Glo18, Gno17, Gon19, Gri23, Gro15, Gue10, Gue24, Gus12, Hal20, Ham21, Han14, Har12a, Har12b, Har08, Her19, Her07, Her12, Heu14, Hil13b, Hir20, Hop01, Hop08]. War [Hor20, How21, Hug02, Hum20, Hus05, Hyu22, Jag22, Jon20, Kan06, Kec24, Ked18, Kel10, Kha11, Kim10, Kim16, KK13, Kin13, Kin24, Knü20, Kra12, Kra14, LPT16, Lar14, Law11, Lev13a, Lev14, Lev13b, Lop22, Lot03, LK20, Lou17a, Lun00, Lun06, Lüt17, LS24, Mac22, Mad05, Mag09, Mag17, MM03, Maj03, Maj13, MZ22, Mar23, kM10, Mar17, Mar16, Mas02, Mas14, Mat15, Mav21, Maw13, McD12, McP21, MM05, Mey21, Mil22, Mis12, Mit07, Mit18a, Mit18b, Mor14b, Mou15, Mue13, Mue06, Muk06, Mum17, Myt11, Nad04, Nat18, Neh04, Noe12, Nut17, O’M12, Ole08, Oña11, Ons22, Ons24, Oso07, Pai06, Pat21, Pat07, Pea12, Pec01, Ped16, Per17, Pet11, Pii09, Ple13, Pop10, Pop15, Pra11, Pre09b, Pre13, Pro23, Pum22, Rab09, RB24]. War [Rea06, Rey13, Rey23, Rhé11, Rie19, Ris13, Rob16a, Rob16b, Roe16, Rog12, Rom14a, Rom11, Rom14b, Rom18b, Rut13, Sal10, Sar09a, Sar13, Sar15, Sau17, Say16, Sba09, Sca12, Sch20a, Sch12, Sch11, Sea21, Set13b, Sha10, Sha17, SL15, Shi09, Sho18, Shu12, Shu24, Sim24, Smi01, Smi12, Sny13c, Sny18, Soa09, Som13, Spe13, Spi06, Spo15, Sta23, Sun20, Sur06, Tam10, Tan14, Tau09, Tob11b, Tob13, Tok22, Tót20, Tra21, Tsv24, Tub05, Ulu03, Var03, Var08, Vit17, Voß16, Web13, Web21, Web22, Wee21, Wen22, WN08, Wes06, Wes12, Wes17, Wet09, Wid18, Wil11, Wol10, Wys12, XZ19, Yeg06, Zhu14, ZVU18, Zub02, ZCP^{+}21, dK09, vA09, Hil21, Mue24, Sau13]. War-era [Web21]. Warfare [Mis06, SS07, Bau23, Bro12, Ger19c]. warily [Pro23]. Warlouzet [Boy14]. Warning [Goo07, Bra21]. Warriors [Pec14, Wil24]. Wars [Gas24, Kan06, Ass10, Col23, Lar13, Bro20b, Pre12, Ris13]. Warsaw [Pal14, Blu12, Del07, Gas24, Lot01, Lüt07, Muk06]. was [Bra12, Sun20, Zac20]. Washington [Rob16a, D’H18, Shu15]. way [Sch08b, Gno18]. ways [Gra10]. weaknesses [Gal19]. Wean [Sar15]. weapon [Her12]. Weapons [Bar17, Arn03, Cas11a, Fil23, Gri23, MV23, PP21, PP23]. Welcome [TL23]. Welfare [ZF19]. Wells [Gon19]. were [Fog20, Win09]. West [Ano15c, Rab09, Alb14, Ban21, Dro15, Eas15, Gij16, Gil10, Grü03, Gup19, Hir20, Hug02, Ian12, Lea23, Lot19, LK20, MM03, Nad04, Neh04, Shu12, Sur08, Töm19, Lak21]. western [Bel15, BB14, Boe19, Com20, Eas05, Gat22, Grü03, Lod10, Mad05, Nut14, Pac03, Pra12, Rea06, Tob11b]. Westernization’ [Neh04]. wheels [Ris11]. Which [Cas20, Per11]. whirlwind [SX10]. White [Ano07a, Low07, RS17, Sar09a, Gue10, O’M12, BR23, Gal19, Knü20, Sny12]. Who [O’M12, Cul03]. Whose [Tul12]. Wild [Aut21]. Wilford [Cha15]. Will [Fil06]. William [Mue24]. Wilson [Gil10]. win [Tro22]. winding [Rhé11]. Winner [Kom20, Hae15, Mav21, Tre20]. winter [Rub14]. wishful [Sai21]. Witch [Lib11]. Witch-hunts [Lib11]. Withdrawal [Ben05, Kal08]. within [Ger19a, Gno17, Sza17]. Wojtyła [Zar05]. wolf [Fai15]. women [Rei16]. Woods [Roy02]. Word [FS23]. words [Kel22]. Worker [Zar05, WD10]. workers [Gat22, Gu22a]. working [Pop10]. World [Gle18, Han15, Hil13a, Mag13, Mar16, Mit18a, Mit18b, Pea12, Pet14, RB24, Rom18b, Sny18, Ull11, Bag11, Cal19, Cas19, Cit12a, Cit12b, Doo21, Dum14b, Fri10, Gfe11, Hir20, Hor20, Kit12, Kli23, Lat11, Ler11, Pha12, Roe16, Sar09b, Shi20, Shi09, Wes10, Whi11, Ewi19, Hat13, Hay18, Hor16, Ked18, Lan22, Las24, Lüt17, Mag13, Mag17, Mat15, Pec14, Rei16, Rog12, Sto14, Woh06, vW11, Lor24]. Worldmaking [Per16]. writers [Sha03b]. Writing [Lin13]. Wyss [Tag20].

x [Kar21, Ped13]. Xavier [P´ 24]. Xiaoping [AC17, kM17]. xii [Cha15, Mue24, Tag20, Pal10]. xiii [Zub16]. xiv [Ces13]. xix [Fis21, Han17]. xl [Mag19]. xvi [Dro13]. xviii [Hil13a, Lop13, Rob16a]. xxii [Set13a]. xxiv [Cha15, Cox24]. XXVIII [Ons13].

Yale [Cox24]. Yalta [Dül12, Foo09]. year [Kra11, Wes10, Lou17b]. years [Cas15b, Cha14, Cul03, Dum10, ZCP^{+}21]. Yellow [Gin02]. Yemen [Lar14]. Yi [LS19]. Yom [Spe13]. York [Cha15, Cox13, Dee10, Han13, Han17, Lin13, Mag09, Man13, Sch20b, Sny13b, Zub16]. young [Dum14a]. Youth [Pea12, Hor16]. Yugoslav [Cav11, Sty05, UK14]. Yugoslav/Serbian [Cav11]. Yugoslavia [Cav11, Cra11, Dim08, Hay18, Mar23, Nie21, Woo13, Zac20].

zaboleli [Kal14]. Zambia [DeR07]. Zanier [Tag20]. Zedong [Ano15b, Far15b, SX14]. Zero [Ada06]. Zero-Hour [Ada06]. Zhang [Liu18]. Zhenbao [Kui00]. Zhivago [Kal11]. Zhou [SX10]. Ziemann [Ano14a]. Zimbabwe [CyJ17]. Zolov [Sch22]. zone [Rab09]. Zubok [Cox24]. Zulu [Mac22].

References

Abadi:2002:GBM

[Aba02] J. Abadi. Great Britain and the Maghreb in the epoch of Pan Arabism and Cold War. Cold War History, 2(2):125–160, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

      tandfonline.com/doi/abs/10.1080/713999956.

Albers:2017:SCS

[AC17] Martin Albers and Zhong Zhong Chen. Socialism, capitalism, and Sino– European relations in the Deng Xiaoping Era, 1978–1992. Cold War History, 17(2):115–119, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-

      7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
      1080/14682745.2017.1315937.

Adamsky:2006:ZHB

[Ada06] Dima P. Adamsky. ‘Zero-Hour for the Bears’: Inquiring into the Soviet decision to intervene in the Egyptian–Israeli War of Attrition, 1969–70. Cold War History, 6(1):113–136, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

      doi/full/10.1080/14682740500395501.

Adamson:2021:NRU

[Ada21] Matthew Adamson. Nuclear reach: uranium prospection and the global ambitions of the French nuclear programme, 1945–65. Cold War History, 21(3):319–336, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2020.1806238.

Aharon:2019:SIL

[Aha19] Eldad Ben Aharon. Superpower by invitation: late Cold War diplomacy and leveraging Armenian terrorism as a means to rapprochement in Israeli–Turkish relations (1980–1987). Cold War History, 19 (2):275–293, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1483342.

Ahlman:2013:BAC

[Ahl13] Jeffrey Ahlman. Battleground Africa: Cold War in the Congo, 1960– 1965. Cold War History, 13(3):435–436, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2013.819641.

Armstrong:2015:SAN

[AK15] Charles K. Armstrong and John Barry Kotch. Sino–American negotiations on Korea and Kissinger’s UN diplomacy. Cold War History, 15 (1):113–134, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.971765.

Albers:2014:BBD

[Alb14] Martin Albers. Business with Beijing, détente with Moscow: West Germany’s China policy in a global context, 1969–1982. Cold War History, 14(2):237–257, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.817989.

Alcalde:2018:WVI

[Alc18] Ángel Alcalde. War veterans, international politics, and the early Cold War, 1945–50. Cold War History, 18(4):409–427, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2018.1455663.

Aldrich:2001:RAL

[Ald01] R. Aldrich. Review article — liberation: Rolling back the frontiers of clandestine Cold War history? Cold War History, 1(2):127–138, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999920.

Alenius:2010:UPN

[Ale10] Kari Alenius. US policy in the Nordic–Baltic region: during the Cold War and after. Cold War History, 10(1):135, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903398938.

Allen:2010:KHJ

[All10] Sir Mark Allen. King Hussein of Jordan: a political life. Cold War History, 10(3):466–467, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2010.493719.

Allison:2024:BAC

[All24] Benjamin V. Allison. Beyond the arc of crisis: Jimmy Carter and the Arab ‘radicals,’ 1978–79. Cold War History, 24(1):67–86, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2253737.

Alvandi:2014:SDK

[Alv14] Roham Alvandi. The Shah’s détente with Khrushchev: Iran’s 1962 missile base pledge to the Soviet Union. Cold War History, 14(3): 423–444, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.890591.

Anonymous:2000:EB

[Ano00a] Anonymous. Editorial board. Cold War History, 1(1):??, 2000. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999908.

Anonymous:2000:EM

[Ano00b] Anonymous. End matter. Cold War History, 1(1):141–146, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999910.

Anonymous:2001:EBa

[Ano01a] Anonymous. Editorial board. Cold War History, 1(2):??, 2001. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999916.

Anonymous:2001:EBb

[Ano01b] Anonymous. Editorial board. Cold War History, 1(3):??, 2001. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999927.

Anonymous:2001:EBc

[Ano01c] Anonymous. Editorial board. Cold War History, 2(1):??, 2001. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999944.

Anonymous:2001:EMa

[Ano01d] Anonymous. End matter. Cold War History, 1(2):139–146, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999923.

Anonymous:2001:EMb

[Ano01e] Anonymous. End matter. Cold War History, 1(3):175–180, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999926.

Anonymous:2001:EMc

[Ano01f] Anonymous. End matter. Cold War History, 2(1):176–178, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999942.

Anonymous:2002:EBa

[Ano02a] Anonymous. Editorial board. Cold War History, 2(2):??, 2002. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999952.

Anonymous:2002:EBb

[Ano02b] Anonymous. Editorial board. Cold War History, 2(3):??, 2002. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999958.

Anonymous:2002:EBc

[Ano02c] Anonymous. Editorial board. Cold War History, 3(1):??, 2002. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999970.

Anonymous:2002:EMa

[Ano02d] Anonymous. End matter. Cold War History, 2(2):236–242, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999950.

Anonymous:2002:EMb

[Ano02e] Anonymous. End matter. Cold War History, 2(3):139–146, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999966.

Anonymous:2002:EMc

[Ano02f] Anonymous. End matter. Cold War History, 3(1):173–178, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999977.

Anonymous:2003:BR

[Ano03a] Anonymous. Book reviews. Cold War History, 3(3):167–172, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391688.

Anonymous:2003:EB

[Ano03b] Anonymous. Editorial board. Cold War History, 3(2):??, 2003. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999984.

Anonymous:2003:EM

[Ano03c] Anonymous. End matter. Cold War History, 3(2):174–178, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999986.

Anonymous:2003:IV

[Ano03d] Anonymous. Index to volume 3. Cold War History, 3(3):173– 175, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391698.

Anonymous:2004:BR

[Ano04] Anonymous. Book reviews. Cold War History, 4(2):192–202, 2004. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740412331391865.

Anonymous:2005:RAF

[Ano05] Anonymous. RETRACTED ARTICLE: ’Fraternal Socialism’: The international reconstruction of North Korea, 1953–62. Cold War History, 5(2):161–187, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
1462740500061160.

Anonymous:2007:IWP

[Ano07a] Anonymous. Introduction: White power, black nationalism and the Cold War in Southern Africa. Cold War History, 7(2):165– 168, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701284090.

Anonymous:2007:MSA

[Ano07b] Anonymous. Map of Southern Africa. Cold War History, 7 (2):163, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701374065.

Anonymous:2009:BR

[Ano09a] Anonymous. Books received. Cold War History, 9(4):545– 546, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903268446.

Anonymous:2009:MSR

[Ano09b] Anonymous. In memoriam: Saki Ruth Dockrill (1952–2009). Cold War History, 9(4):547–549, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903292453.

Anonymous:2010:BR

[Ano10a] Anonymous. Books received. Cold War History, 10(4):585– 588, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.513516.

Anonymous:2010:EB

[Ano10b] Anonymous. Editorial board. Cold War History, 10(4):??, 2010. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/14682745.2010.540145.

Anonymous:2010:O

[Ano10c] Anonymous. Obituary. Cold War History, 10(2):285, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.481228.

Anonymous:2011:BR

[Ano11a] Anonymous. Books received. Cold War History, 11(4):687– 688, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.617486.

Anonymous:2011:EB

[Ano11b] Anonymous. Editorial board. Cold War History, 11(4):??, 2011. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/14682745.2011.628188.

Anonymous:2011:E

[Ano11c] Anonymous. Erratum. Cold War History, 11(1):131, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.561131.

Anonymous:2011:MIV

[Ano11d] Anonymous. In memoriam: [Ilya V. Gaiduk, 1961–2011]. Cold War History, 11(4):487, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2011.630271.

Anonymous:2012:EB

[Ano12a] Anonymous. Editorial board. Cold War History, 12(4):??, 2012. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/14682745.2012.743762.

Anonymous:2012:N

[Ano12b] Anonymous. Note. Cold War History, 12(4):571, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2012.743771.

Anonymous:2013:EB

[Ano13] Anonymous. Editorial board. Cold War History, 13(4):??, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2013.856198.

Anonymous:2014:CNH

[Ano14a] Anonymous. Corrigendum: Nehring, H., & Ziemann, B. (2012) ‘Do all paths lead to Moscow? The NATO dual-track decision and the peace movement — a critique.’ Cold War History 12, no. 1 (2012). DOI: 10.1080/14682745.2011.625160. Cold War History, 14 (3):454, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.927688. See [NZ12].

Anonymous:2014:EB

[Ano14b] Anonymous. Editorial board. Cold War History, 14(4):??, 2014. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2014.966591.

Anonymous:2015:BR

[Ano15a] Anonymous. Books received. Cold War History, 15(2):265– 271, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1018113.

Anonymous:2015:CFM

[Ano15b] Anonymous. Corrigendum: Fardella, E. M. (2014). Article, Mao Zedong and the 1962 Cuban Missile Crisis, Cold War History. https://doi.org/10.1080/14682745.2014.971017. Cold War History, 15 (3):??, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.1000035. See [Far15b].

Anonymous:2015:EDM

[Ano15c] Anonymous. Erratum: Drogan M. 2014. The nuclear nation and the German question: an American reactor in West Berlin. Cold War History. https://doi.org/10.1080/14682745.2014.959500. Cold War History, 15(3):??, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.978523. See [Dro15].

Anonymous:2015:NNF

[Ano15d] Anonymous. Note on new format. Cold War History, 15(2): 253, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1017991.

Anonymous:2016:BR

[Ano16a] Anonymous. Books received. Cold War History, 16(1):121– 124, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1123137.

Anonymous:2016:FRR

[Ano16b] Anonymous. Federico Romero in response to Pierre Grosser. Cold War History, 16(3):327, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1155825. See [Gro15].

Anonymous:2017:BR

[Ano17a] Anonymous. Books received. Cold War History, 17(1):107– 109, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1259732.

Anonymous:2017:EB

[Ano17b] Anonymous. Editorial board. Cold War History, 17(4):??, 2017. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2017.1382770.

Anonymous:2018:BR

[Ano18] Anonymous. Books received. Cold War History, 18(1):109– 110, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1419591.

Anonymous:2020:SRF

[Ano20] Anonymous. Statement of retraction: “‘Fraternal Socialism”: The international reconstruction of North Korea, 1953–62’. Cold War History, 20(2):253, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
14682745.2020.1724643.

Anonymous:2021:CSC

[Ano21] Anonymous. Call for submissions: Conversations on Cold War history. Cold War History, 21(4):601–602, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1988221.

Apor:2020:SFE

[Apo20] Balázs Apor. Stalin and the fate of Europe: the post-war struggle for sovereignty. Cold War History, 20(2):246–248, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2020.1746007.

Appelqvist:2008:RUE

[App08] Örjan Appelqvist. Rediscovering uncertainty: early attempts at a pan-European post-war recovery. Cold War History, 8(3):327–352, 2008. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740802222130.

Arnold:2003:HNW

[Arn03] Lorna Arnold. The history of nuclear weapons: The Frisch–Peierls memorandum on the possible construction of atomic bombs of February 1940. Cold War History, 3(3):111–126, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740312331391658.

Arnson:2013:LAC

[Arn13] Cynthia J. Arnson. Latin America’s Cold War. Cold War History, 13 (1):133–134, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.762234.

Arnhold:2021:NNA

[Arn21] Valerie Arnhold. Normalisation of nuclear accidents after the Cold War. Cold War History, 21(3):261–281, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1806239.

Ashton:2007:R

[Ash07] Nigel J. Ashton. Reviews. Cold War History, 7(2):333–348, 2007. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740701284397.

Ashton:2009:MKB

[Ash09] Nigel J. Ashton. Macmillan, Khrushchev and the Berlin Crisis, 1958– 1960. Cold War History, 9(4):538–539, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740903268412.

Asinovskiy:2024:RUT

[Asi24] Dmitry Asinovskiy. Rostislav Ulianovskii, the Tudeh Party of Iran and Soviet attempts to set Iran on a non-capitalist path of development (1979–83). Cold War History, 24(1):45–65, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2217753.

Asselin:2010:MSV

[Ass10] Pierre Asselin. Making sense of the Vietnam wars: local, national, and transnational perspectives. Cold War History, 10(1):133–134, 2010. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740903398953.

Asselin:2011:DRV

[Ass11] Pierre Asselin. The Democratic Republic of Vietnam and the 1954 Geneva Conference: a revisionist critique. Cold War History, 11(2): 155–195, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903244934.

Asselin:2012:EAL

[Ass12] Pierre Asselin. Explaining America’s lost war. Cold War History, 12 (2):367–368, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.686584.

Aunesluoma:2006:R

[Aun06] Juhana Aunesluoma. Reviews. Cold War History, 6(2):261– 277, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600650268.

Autiello:2021:TWD

[Aut21] Nicholas Anthony Autiello. Taming the wild dragon: John F. Kennedy and the Republic of China, 1961–63. Cold War History, 21 (1):71–89, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1550077.

Badalassi:2018:NTM

[Bad18] Nicolas Badalassi. “Neither too much nor too little” France, the USSR and the Helsinki CSCE. Cold War History, 18(1):1–17, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2017.1356821.

Baginska:2011:RFC

[Bag11] Joanna Baginska. The red flag: communism and the making of the modern world. Cold War History, 11(2):276, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.569151.

Baines:2018:AVA

[Bai18] Gary Baines. An African Volk: the apartheid regime and the search for survival. Cold War History, 18(3):381–384, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2018.1489997.

Banks:2005:BCC

[Ban05] Alexander J. Banks. Britain and the Cambodian Crisis of spring 1970. Cold War History, 5(1):87–105, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/1468274042000283153.

Bandarra:2021:BLW

[Ban21] Leonardo Bandarra. From Bonn with love: West German interests in the 1975 nuclear agreement with Brazil. Cold War History, 21(3): 337–355, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2020.1819245.

Barnett:2015:NPA

[Bar15] Nicholas Barnett. ‘No protection against the H-bomb’: press and popular reactions to the Coventry civil defence controversy, 1954. Cold War History, 15(3):277–300, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.968558.

Baron:2017:IWS

[Bar17] Nick Baron. Innocent weapons: The Soviet and American politics of childhood in the Cold War. Cold War History, 17(2):202–204, 2017. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2017.1287240.

Batovic:2009:SUJ

[Bat09] Ante Batovic. The Soviet Union and the June 1967 Six Day War. Cold War History, 9(4):529–530, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903268420.

Baudet:2016:SAT

[Bau16] Floribert Baudet. ‘A statement against the totalitarian countries of Europe’: human rights and the early Cold War. Cold War History, 16 (2):125–140, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1128898.

Bauerkamper:2023:RRA

[Bau23] Arnd Bauerkämper. Revolutionaries for the right: anticommunist internationalism and paramilitary warfare in the Cold War. Cold War History, 23(2):327–328, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2022.2063752.

Blavoukos:2014:CBW

[BB14] Spyros Blavoukos and Dimitris Bourantonis. Calling the bluff of the Western powers in the United Nations disarmament negotiations, 1954– 55. Cold War History, 14(3):359–376, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2013.871261.

Bessner:2024:BHC

[BBF24] Daniel Bessner, Michael Brenes, and Michael Franczak. A brief history of Cold War liberalism. Cold War History, 24(2):299–308, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2024.2310155.

Becker:2023:CCA

[Bec23] Marc Becker. The CIA and creole anticommunism in Cold War Ecuador. Cold War History, 23(3):363–387, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2167981.

Bekes:2006:HRD

[Bék06] Csaba Békés. The 1956 Hungarian Revolution and the declaration of neutrality. Cold War History, 6(4):477–500, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740600979261.

Bell:2015:MWC

[Bel15] Aaron T. Bell. A matter of western civilisation: transnational support for the Salvadoran counterrevolution, 1979–1982. Cold War History, 15 (4):511–531, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1019870.

Bellotti:2023:RSD

[Bel23] Alissa Bellotti. The rock star and the dictator: Udo Lindenberg’s East German celebrity diplomacy. Cold War History, 23(1):83–102, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2022.2122446.

Benvenuti:2005:BMW

[Ben05] Andrea Benvenuti. The British military withdrawal from Southeast Asia and its impact on Australia’s Cold War strategic interests. Cold War History, 5(2):189–210, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740500061178.

Berghahn:2010:DAA

[Ber10] Volker R. Berghahn. The debate on ‘Americanization’ among economic and cultural historians. Cold War History, 10(1):107–130, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903388566.

Best:2012:WVW

[Bes12] Antony Best. ‘We are virtually at war with Russia’: Britain and the Cold War in East Asia, 1923–40. Cold War History, 12(2): 205–225, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569436.

Bidgood:2024:NRR

[Bid24] Sarah Bidgood. No risk (reduction), no reward: Re-examining the 1973 US–Soviet Agreement on the Prevention of Nuclear War. Cold War History, 24(2):209–238, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2023.2239727.

Bjork:2021:CBP

[Bjo21] Jim Bjork. Catholics on the barricades: Poland, France and the ‘Revolution’, 1891–1956. Cold War History, 21(4):533–535, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1944518.

Bjoernsson:2023:NAC

[Bjo23] Iben Bjoernsson. Negotiating Armageddon: civil defence in NATO and Denmark 1949–59. Cold War History, 23(2):217–238, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2022.2123915.

Burr:2003:NSN

[BK03] W. Burr and J. Kimball. Nixon’s secret nuclear alert: Vietnam War diplomacy and the Joint Chiefs of Staff readiness test, October 1969. Cold War History, 3(2):113–156, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/713999987.

Black:2022:IPP

[Bla22] Megan A. Black. Interior’s proper place: response to Tyler Priest. Cold War History, 22(4):539–543, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2123297. See [Pri22].

Bloem:2005:R

[Blo05] Filip Bloem. Reviews. Cold War History, 5(3):377–394, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740500222168.

Bluth:2001:RIA

[Blu01] C. Bluth. Reconciling the irreconcilable: Alliance politics and the paradox of extended deterrence in the 1960s. Cold War History, 1(2):73–102, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/713999919.

Bluth:2012:ACP

[Blu12] Christoph Bluth. Arms control as a part of strategy: the Warsaw Pact in MBFR negotiations. Cold War History, 12(2):245–268, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.557364.

Boel:2019:WJS

[Boe19] Bent Boel. Western journalism in the Soviet bloc during the Cold War: themes, approaches, theses. Cold War History, 19(4):593–614, 2019. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2019.1670421.

Booth:2016:MCW

[Boo16] William A. Booth. Mexico’s Cold War: Cuba, the United States, and the legacy of the Mexican Revolution. Cold War History, 16(3): 333–335, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1187718.

Boquerat:2005:ICP

[Boq05] Gilles Boquérat. India’s commitment to peaceful coexistence and the settlement of the Indochina War. Cold War History, 5(2):211– 234, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500062119.

Borhi:2001:ECS

[Bor01] L. Borhi. Empire by coercion: The Soviet Union and Hungary in the 1950s. Cold War History, 1(2):47–72, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/713999917.

Borhi:2004:WHC

[Bor04] László Borhi. ‘We Hungarian communists are realists’: János Kádár’s foreign policy in the light of Hungarian–US relations, 1957–67. Cold War History, 4(2):1–32, 2004. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740412331391795.

Borocz:2018:PSH

[Bör18] József Böröcz. Performing socialist Hungary in China: ‘modern, Magyar, European’. Cold War History, 18(3):257–274, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2018.1425293.

Borowy:2020:BRS

[Bor20] Iris Borowy. Book review: Sara Lorenzini, Global development: a Cold War history (Princeton, Princeton University Press, 2019). Cold War History, 20(4):523–526, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1787593.

Bousquet:2008:CAW

[Bou08] Antoine Bousquet. Cyberneticizing the American war machine: science and computers in the Cold War. Cold War History, 8(1):

77–102, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701791359.

Boyce:2014:BRL

[Boy14] Robert Boyce. Book review: Laurent Warlouzet, Le choix de la CEE par la France: L’Europe économique en débat de Mende‘s France à de Gaulle (1955–1969) (Paris: Comité pour l’histoire économique et financière de la France Ministère de l’Économie, des Finances et de l’Industrie, Ministère du Budget, des Comptes publics et de la Réforme de l’État, 2011), ix + 569 pp. Cold War History, 14(2): 285–287, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.893685.

Boyce:2019:PSI

[Boy19] Robert Boyce. The practice of socialist internationalism: European socialists and international politics, 1914–1960. Cold War History, 19 (2):320–322, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1601410.

Bozo:2007:MFE

[Boz07] Frédéric Bozo. Mitterrand’s France, the end of the Cold War, and German unification: a reappraisal. Cold War History, 7(4): 455–478, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701621689.

Bozek:2013:BES

[Boz13] Felix Jack Grenfell Bozek. Britain, European security and freer movement: the development of Britain’s CSCE policy 1969–1972. Cold War History, 13(4):439–461, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.756870.

Brawley:2023:SWA

[BR23] Sean Brawley and Mathew Radcliffe. Selling White Australia: the Asian Visits Fund and assimilation as a foundational concept in Australian Cold War public diplomacy. Cold War History, 23(1):103– 119, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2059072.

Brands:2007:NPD

[Bra07] Hal Brands. Non-proliferation and the dynamics of the Middle Cold War: The superpowers, the MLF, and the NPT. Cold War History, 7 (3):389–423, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701474857.

Brands:2012:SHU

[Bra12] Hal Brands. Saddam Hussein, the United States, and the invasion of Iran: was there a green light? Cold War History, 12(2): 319–343, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.564612.

Brain:2016:AAG

[Bra16] Stephen Brain. The appeal of appearing green: Soviet–American ideological competition and Cold War environmental diplomacy. Cold War History, 16(4):443–462, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.971015.

Brand:2021:IWP

[Bra21] Melanie Brand. Intelligence, warning, and policy: the Johnson administration and the 1968 Soviet invasion of Czechoslovakia. Cold War History, 21(2):197–214, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2020.1752675.

Brown:2007:PEC

[Bro07] Archie Brown. Perestroika and the end of the Cold War. Cold War History, 7(1):1–17, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740701197631. See comments [Pat07].

Brown:2010:GMT

[Bro10] Archie Brown. Gorbachev: man of the twentieth century? Cold War History, 10(4):571–572, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2010.513503.

Brogi:2011:STI

[Bro11] Alessandro Brogi. Stalin and Togliatti: Italy and the origins of the Cold War. Cold War History, 11(4):682–683, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2011.617435.

Brogi:2012:ACB

[Bro12] Alessandro Brogi. Ambassador Clare Boothe Luce and the evolution of psychological warfare in Italy. Cold War History, 12(2): 269–294, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.558611.

Brock:2014:MPO

[Bro14] Maria Brock. Mr Putin: Operative in the Kremlin. Cold War History, 14(2):289–291, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.893686.

Brockett:2020:DSU

[Bro20a] Charles D. Brockett. Disturbing secrets: US–Costa Rican relations during the Nixon administration. Cold War History, 20(3): 253–271, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1673372.

Brown:2020:CWC

[Bro20b] Lauren Erin Brown. Cold War, Culture Wars, War on Terror: the NEA and the art of public diplomacy. Cold War History, 20(4): 379–397, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1706165.

Buffet:2009:DLC

[Buf09a] Cyril Buffet. ‘Declaration of love on celluloid’: The depiction of the Berlin Wall in a GDR film, 1961–62. Cold War History, 9(4): 469–487, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903198205.

Buffet:2009:EP

[Buf09b] Cyril Buffet. Editor’s preface. Cold War History, 9(4):453– 454, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903198171.

Burr:2001:SAR

[Bur01] W. Burr. Sino–American relations, 1969: The Sino–Soviet border war and steps towards rapprochement. Cold War History, 1(3):73–112, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999930.

Burton:2019:NGS

[Bur19] Eric Burton. Navigating global socialism: Tanzanian students in and beyond East Germany. Cold War History, 19(1):63–83, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2018.1485146.

Burnham:2022:SLS

[Bur22] Thomas Burnham. Supping with a long spoon in the Indian Ocean: the negotiation of the 1972 Agreement on Economic and Technical Cooperation between Mauritius and the People’s Republic of China. Cold War History, 22(3):245–263, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1997990.

Busch:2001:SWB

[Bus01] P. Busch. Supporting the war: Britain’s decision to send the Thompson Mission to Vietnam, 1960–61. Cold War History, 2(1):69–94, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999936.

Brooks:2007:CEC

[BW07] Stephen G. Brooks and William C. Wohlforth. Clarifying the end of Cold War debate. Cold War History, 7(3):447–454, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740701474907.

Byrne:2017:DDU

[Byr17] Malcolm Byrne. Dealing with dictators: the United States, Hungary, and East Central Europe, 1942–1989. Cold War History, 17 (4):443–445, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1382975.

Byrnes:2019:DEE

[Byr19] Melissa K. Byrnes. Diplomacy at the end of empire: evolving French perspectives on Portuguese colonialism in the 1950s and 1960s. Cold War History, 19(4):477–491, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2019.1597857.

Calandri:2019:IDW

[Cal19] Elena Calandri. Italy, the developing world, and aid policy, 1969–1979: the ‘historic compromise’ and Italian foreign policy. Cold War History, 19(3):363–381, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1556640.

Campbell:2008:CMC

[Cam08] Edwina S. Campbell. A comment on Michael Cox’s ‘another transatlantic split? American and European narratives and the end of the Cold War’. Cold War History, 8(1):103–113, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740701791367. See [Cox07].

Campus:2015:MNC

[Cam15] Leonardo Campus. Missiles have no colour: African Americans’ reactions to the Cuban Missile Crisis. Cold War History, 15(1): 49–72, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.904291.

Cary:2011:ODB

[Car11] Noel D. Cary. Olympics in divided Berlin? Popular culture and political imagination at the Cold War frontier. Cold War History, 11 (3):291–316, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003686131.

Casey:2011:AHU

[Cas11a] Steven Casey. After Hiroshima: the United States, race and nuclear weapons in Asia, 1945–1965. Cold War History, 11(3):473– 474, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.599975.

Casey:2011:WKT

[Cas11b] Steven Casey. The war for Korea, 1950–1951: they came from the north. Cold War History, 11(2):281–282, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2011.569156.

Castano:2015:PTD

[Cas15a] David Castaño. ‘A practical test in the détente’: International support for the Socialist Party in the Portuguese Revolution (1974–1975). Cold War History, 15(1):1–26, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.932349.

Castiglioni:2015:NLC

[Cas15b] Claudia Castiglioni. No longer a client, not yet a partner: the US–Iranian alliance in the Johnson years. Cold War History, 15(4): 491–509, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1019868.

Casey:2019:AEA

[Cas19] Steven Casey. The age of Eisenhower: America and the world in the 1950s. Cold War History, 19(4):587–588, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2019.1678816.

Casals:2020:WBY

[Cas20] Marcelo Casals. Which borders have not yet been crossed? A supplement to Gilbert Joseph’s historiographical balance of the Latin American Cold War. Cold War History, 20(3):367–372, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2020.1762311. See

[Jos19] and response [Jos20].

Cavoski:2011:OBB

[Cav11] Jovan Cavoski. Overstepping the Balkan boundaries: The lesser known history of Yugoslavia’s early relations with Asian countries (new evidence from Yugoslav/Serbian archives). Cold War History, 11(4): 557–577, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003704223.

Celik:2022:RCW

[Çel22] Haydar Seçkin Çelik. Resurgence of the Cold War state of mind: the debate on constitutional tolerance of socialism vis-à-vis the emerging left in Turkey (1967–1971). Cold War History, 22(4):499–520, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2100354.

Cesari:2013:BRS

[Ces13] Laurent Cesari. Book review: Seth Jacobs, The Universe Unraveling: American Foreign Policy in Cold War Laos (Ithaca, NY and London: Cornell University Press, 2012), xiv + 312 pp. Cold War History, 13 (2):280–281, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.789662.

Chamberlin:2012:SEP

[Cha12] Paul Thomas Chamberlin. Schönau and the Eagles of the Palestinian Revolution: refugees, guerillas, and human rights in the global 1970s. Cold War History, 12(4):595–614, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.665368.

Chaudhuri:2014:WPM

[Cha14] Rudra Chaudhuri. War and peace in Modern India: a strategic history of the Nehru years. Cold War History, 14(4):705–706, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.955690.

Chamberlin:2015:BRH

[Cha15] Paul Thomas Chamberlin. Book reviews: Hugh Wilford, America’s Great Game: The CIA’s Secret Arabists and the Shaping of the Modern Middle East (New York: Basic Books, 2013), xxiv + 342 pp. Roham Alvandi, Nixon, Kissinger, and the Shah: The United States and Iran in the Cold War (New York: Oxford University Press, 2014), xii + 255 pp. Cold War History, 15(4):586–589, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1086473.

Chen:2014:DME

[Che14] Zhong Zhong Chen. Defying Moscow: East German–Chinese relations during the Andropov–Chernenko interregnum, 1982–1985. Cold War History, 14(2):259–280, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.876009.

Chen:2019:EGP

[Che19] Tao Chen. East German pragmatism, China’s policy of differentiation, and Soviet miscalculation: Hermann Matern’s 1961 trip to China revisited. Cold War History, 19(1):85–99, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2018.1483341.

Chen:2021:PPT

[Che21] Kuan-Jen Chen. Projecting Poseidon’s Trident: America’s East Asia and the shifting contours of 1950s post-war naval policy. Cold War History, 21(4):391–410, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2020.1752676.

Cheng:2022:CAG

[Che22] Xiaohe Cheng. The Chinese advisory groups in the first Indochina War: their formation, evolution, and disbandment. Cold War History, 22(2):195–213, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1939309.

Chiampan:2014:BRU

[Chi14] Andrea Chiampan. Book review: Umberto Tulli, Tra diritti umani e distensione: L’amministrazione Carter e il dissenso in Urss (Milano: Franco Angeli, 2013), 256 pp. Cold War History, 14(3): 452–453, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.917800.

Chourchoulis:2012:NDN

[Cho12a] Dionysios Chourchoulis. A nominal defence? NATO threat perception and responses in the Balkan area, 1951–1967. Cold War History, 12 (4):637–657, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.671297.

Chourchoulis:2012:STC

[Cho12b] Dionysios Chourchoulis. Stalin and the Turkish crisis of the Cold War, 1945–1953. Cold War History, 12(4):711–713, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2012.737561.

Choi:2013:LFC

[Cho13] Lyong Choi. Literature and film in Cold War South Korea: Freedom’s frontier. Cold War History, 13(3):432–433, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.819646.

Choi:2014:FNC

[Cho14] Lyong Choi. The first nuclear crisis in the Korean peninsula, 1975–76. Cold War History, 14(1):71–90, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.787412.

Ciglioni:2017:IPO

[Cig17] Laura Ciglioni. Italian public opinion in the Atomic Age: Mass-market magazines facing nuclear issues (1963–1967). Cold War History, 17 (3):205–221, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1291633.

Ciorciari:2014:CPP

[Cio14] John D. Ciorciari. China and the Pol Pot regime. Cold War History, 14(2):215–235, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.808624.

Citino:2012:CIUa

[Cit12a] Nathan J. Citino. The ‘crush’ of ideologies: The United States, the Arab world, and Cold War modernisation. Cold War History, 12 (1):89–110, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.506608.

Citino:2012:CIUb

[Cit12b] Nathan J. Citino. The ‘crush’ of ideologies: The United States, the Arab world, and Cold War modernisation. Cold War History, 12(2):379, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.691358.

Citino:2019:MEC

[Cit19] Nathan J. Citino. The Middle East and the Cold War. Cold War History, 19(3):441–456, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2019.1576677.

Clarkson:2008:HIP

[Cla08] Alexander Clarkson. Home and away: immigration and political violence in the Federal Republic of Germany, 1945–90. Cold War History, 8(1):1–21, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701791193.

Classen:2013:CAG

[Cla13] Christoph Classen. Captive audience? GDR radio in the mirror of listeners’ mail. Cold War History, 13(2):239–254, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2012.757136.

Clivio:2019:NNA

[Cli19] Carlotta Clivio. Neither for, nor against Mao: PCI-CCP interactions and the normalisation of Sino–Italian Relations, 1966–71. Cold War History, 19(3):383–400, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2018.1529758.

Colman:2012:LCC

[Col12] Jonathan Colman. Lost crusader? Chester L. Cooper and the Vietnam War, 1963–68. Cold War History, 12(3):429–449, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.573147.

Colbourn:2018:CTN

[Col18] Susan Colbourn. ‘Cruising toward nuclear danger’: Canadian antinuclear activism, Pierre Trudeau’s peace mission, and the transatlantic partnership. Cold War History, 18(1):19–36, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2017.1370456.

Coldiron:2023:CMW

[Col23] Katie Coldiron. Cuban memory wars: retrospective politics in revolution and exile. Cold War History, 23(2):328–330, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2022.2087945.

Comte:2020:WCW

[Com20] Emmanuel Comte. Waging the Cold War: the origins and launch of Western cooperation to absorb migrants from Eastern Europe, 1948–57. Cold War History, 20(4):461–481, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1756781.

Connolly:2005:AFS

[Con05] Chris Connolly. The American factor: Sino–American rapprochement and Chinese attitudes to the Vietnam War, 1968–72. Cold War History, 5(4):501–527, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500284887.

Connolly:2010:CWI

[Con10] Chris A. Connolly. Cold War island: Quemoy on the front line. Cold War History, 10(1):132–133, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903398979.

ConfortodeOliveira:2023:IMM

[Con23] Fernanda Conforto de Oliveira. The IMF as a ‘mantle of multilateral anonymity’: US–IMF–Brazil relations, 1956–9. Cold War History, 23(1):1–21, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2078312.

Cooper:2022:TRS

[Coo22] Austin R. Cooper. The Tunisian request: Saharan fallout, US assistance and the making of the International Atomic Energy Agency. Cold War History, 22(4):407–436, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2085252.

Corthorn:2013:CWB

[Cor13] Paul Corthorn. The Cold War and British debates over the boycott of the 1980 Moscow Olympics. Cold War History, 13(1):43–66, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2012.727799.

Costigliola:2000:CFE

[Cos00] F. Costigliola. ’I had come as a friend’: Emotion, cultural, and ambiguity in the formation of the Cold War, 1943–45. Cold War History, 1(1):103–128, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999913.

Cosse:2023:HRS

[Cos23] Isabella Cosse. Human rights and the status of children as victims in the Late Cold War. Cold War History, 23(3):339–361, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2167980.

Cox:2007:ATS

[Cox07] Michael Cox. Another transatlantic split? American and European narratives and the end of the Cold War. Cold War History, 7(1): 121–146, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/

14682740701218882. See comment [Cam08].

Cox:2010:DBP

[Cox10] Michael Cox. Documents on British policy overseas, series III, volume VII: German unification 1989–1990. Cold War History, 10(4): 581–583, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.513515.

Cox:2013:BRJ

[Cox13] Michael Cox. Book reviews: John Lewis Gaddis, George F. Kennan: An American Life (New York: The Penguin Press, 2011), 762 pp. Campbell Craig and Frederik Logevall, America’s Cold War: The Politics of Insecurity (Cambridge, MA: Harvard University Press, 2009), 426 pp. Jean Edward Smith, Eisenhower in War and Peace (New York: Random House, 2012), 950 pp. Heonik Kwon, The Other Cold War (New York: Columbia University Press, 2010), 204 pp. Cold War History, 13 (2):277–280, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.789663.

Cox:2014:LHS

[Cox14] Michael Cox. Learning from history? From Soviet collapse to the ‘new’ Cold War. Cold War History, 14(4):461–485, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2014.950241.

Cox:2015:LEF

[Cox15] Michael Cox. The Last Empire: The final days of the Soviet Union. Cold War History, 15(2):262–264, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1018480.

Cox:2018:SDC

[Cox18] Michael Cox. Safeguarding democratic capitalism: US foreign policy and national security, 1920–2015. Cold War History, 18(4):

521–522, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1533687.

Cox:2021:AIC

[Cox21] Michael Cox. ‘An alien ideology’: Cold War perceptions of the Irish Republican Left. Cold War History, 21(3):375–379, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2021.1906839.

Cox:2024:BRC

[Cox24] Michael Cox. Book review: Collapse: the fall of the Soviet Union. Vladislav M. Zubok, (New Haven: Yale University, 2021), xxiv + 560 pp. Cold War History, 24(2):351–353, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2022.2112133.

Chubariyan:2000:MLS

[CP00] A. Chubariyan and V. Pechatnov. Molotov ’the liberal’: Stalin’s 1945 criticism of his deputy. Cold War History, 1(1):129–140, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999911.

Crampton:2011:YSU

[Cra11] Richard J. Crampton. Yugoslavia and the Soviet Union in the early Cold War: reconciliation, comradeship, confrontation, 1953–1957. Cold War History, 11(2):280–281, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.569155.

Craig:2015:NSH

[Cra15] Campbell Craig. Nuclear statecraft: History and strategy in America’s Atomic Age. Cold War History, 15(2):258–262, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2015.1018477.

Craig:2016:TWC

[Cra16] Malcolm M. Craig. ‘I think we cannot refuse the order’: Britain, America, nuclear non-proliferation, and the Indian Jaguar deal, 1974–1978. Cold War History, 16(1):61–81, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1078314.

Creswell:2002:LHO

[Cre02] M. Creswell. ’With a little help from our friends’: How France secured an Anglo–American continental commitment, 1945–54. Cold War History, 3(1):1–28, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999975.

Crockatt:2008:BR

[Cro08] Richard Crockatt. Book reviews. Cold War History, 8(4):547– 567, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802373644.

Crockatt:2009:LCG

[Cro09] Richard Crockatt. Local consequences of the global Cold War. Cold War History, 9(4):542–543, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903268461.

Crockatt:2011:CHC

[Cro11] Richard Crockatt. The Cambridge history of the Cold War. Cold War History, 11(2):269–272, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.569147.

Cerny:2020:SPC

[CS20] Vladimír Cerný and Petr Suchý. Spies and peaceniks: Czechoslovak intelligence attempts to thwart NATO’s Dual-Track Decision. Cold War History, 20(3):273–291, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1724963.

Cull:2003:MWI

[Cul03] Nicholas J. Cull. ‘The man who invented truth’:the tenure of Edward R. Murrow as director of the United States Information Agency during the Kennedy years. Cold War History, 4(1):23–48, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740312331391724.

Culbert:2013:JEH

[Cul13] David Culbert. J. Edgar Hoover goes to the movies: The FBI and the origins of Hollywood’s Cold War. Cold War History, 13(3): 431–432, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.819642.

Cullen:2018:PCW

[Cul18] Poppy Cullen. ‘Playing Cold War politics’: the Cold War in Anglo– Kenyan relations in the 1960s. Cold War History, 18(1):37–54, 2018. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2017.1387774.

Choi:2017:NKZ

[CyJ17] Lyong Choi and Il young Jeong. North Korea and Zimbabwe, 1978– 1982: from the strategic alliance to the symbolic comradeship between Kim Il Sung and Robert Mugabe. Cold War History, 17(4): 329–349, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1328406.

Czernecki:2013:IOF

[Cze13] Igor Czernecki. An intellectual offensive: The Ford Foundation and the destalinization of the Polish social sciences. Cold War History, 13 (3):289–310, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.756473.

Daigle:2009:FDS

[Dai09] Craig Daigle. Foxbats over Dimona: the Soviets nuclear dilemma in the Six Day War. Cold War History, 9(4):527–529, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740903268487.

Daigle:2019:SAD

[Dai19] Craig Daigle. Sadat’s African dilemma: Libya, Ethiopia, and the making of the Camp David Accords. Cold War History, 19(2): 295–313, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1547710.

Dalsjo:2014:HRS

[Dal14] Robert Dalsjö. The hidden rationality of Sweden’s policy of neutrality during the Cold War. Cold War History, 14(2):175–194, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.765865.

Danaher:2022:BRD

[Dan22] David S. Danaher. Book review: David Galbreath Barton, Havel: unfinished revolution (Pittsburgh: University of Pittsburgh Press, 2020), 374 pp. Cold War History, 22(1):125–127, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1951780.

Davies:2003:SCE

[Dav03] Sarah Davies. Soviet cinema and the early Cold War: Pudovkin’s Admiral Nakhimov. Cold War History, 4(1):49–70, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740312331391734.

Davy:2009:HMS

[Dav09] Richard Davy. Helsinki myths: setting the record straight on the Final Act of the CSCE, 1975. Cold War History, 9(1):1–22, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740802490380.

Day:2022:HSB

[Day22] Christopher Day. Helmut Schmidt and British–German relations: a European misunderstanding. Cold War History, 22(2):242–244, 2022. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2021.1993524.

Deery:2006:DSC

[Dee06] Phillip Deery. ‘A divided soul’? The Cold War odyssey of O. John Rogge. Cold War History, 6(2):177–204, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740600650219.

Deery:2010:RHA

[Dee10] Phillip Deery. ‘Running with the hounds’: Academic McCarthyism and New York University, 1952–53. Cold War History, 10(4): 469–492, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903527692.

Deighton:2006:DBR

[Dei06] Anne Deighton. ‘A different 1956’: British responses to the Polish events, June–November 1956. Cold War History, 6(4):455– 475, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600979253.

Deighton:2013:DDI

[Dei13] Anne Deighton. Don and diplomat: Isaiah Berlin and Britain’s early Cold War. Cold War History, 13(4):525–540, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.773213.

Deletant:2007:TBR

[Del07] Dennis Deletant. ‘Taunting the Bear’: Romania and the Warsaw Pact, 1963–89. Cold War History, 7(4):495–507, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740701621796.

DellOmo:2020:TFG

[Del20] Augusta Lynn Dell’Omo. Tending the flock: German Lutherans, reconstruction, and prisoners of war. Cold War History, 20(2):123– 141, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1573814.

DeRoche:2007:NAR

[DeR07] Andy DeRoche. Non-alignment on the racial frontier: Zambia and the USA, 1964–68. Cold War History, 7(2):227–250, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740701284132.

DeVore:2011:UKL

[DeV11] Marc DeVore. The United Kingdom’s last hot war of the Cold War: Oman, 1963–75. Cold War History, 11(3):441–471, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2010.498823.

DHaeseleer:2018:DLS

[D’H18] Brian D’Haeseleer. ‘Drawing the line’ in El Salvador: Washington confronts insurgency in El Salvador, 1979–92. Cold War History, 18 (2):131–148, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1351430.

Dietl:2002:DAG

[Die02] R. Dietl. ’Une déception amoureuse’? Great Britain, the continent and European nuclear cooperation, 1953–57. Cold War History, 3(1):29–66, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/713999973.

Dietl:2003:TES

[Die03] Ralph Dietl. Towards a European security and defence policy. Cold War History, 3(3):144–166, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740312331391678.

Diedrich:2011:CCW

[Die11] Torsten Diedrich. A child of the Cold War — the state and society of the GDR dictatorship as a military-political result of the clash of systems. Cold War History, 11(4):601–624, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.571881.

Dimic:2008:HCW

[Dim08] Ljubodrag Dimić. Historiography on the Cold War in Yugoslavia: from ideology to science. Cold War History, 8(2):285–297, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740802018835.

Dimitrov:2010:SEE

[Dim10] Vesselin Dimitrov. The Sovietization of Eastern Europe: new perspectives on the postwar period. Cold War History, 10(2):274–275, 2010. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682741003711830.

Dippnall:2018:HAG

[Dip18] Stephen Dippnall. Hating America? Great Britain and the execution of Julius and Ethel Rosenberg. Cold War History, 18(1):55–71, 2018. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2017.1355910.

Dittmer:2011:TSH

[Dit11] Lowell Dittmer. Two suns in the heavens: the Sino–Soviet struggle for supremacy, 1962–1967. Cold War History, 11(2):278–279, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.569154.

deKeghel:2009:MEV

[dK09] Isabelle de Keghel. Meeting on the Elbe (Vstrecha na El’be): a visual representation of the incipient Cold War from a Soviet perspective.

Cold War History, 9(4):455–467, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903198197.

Do:2022:CEP

[Do22] Khue Dieu Do. Comrades at enmity: Pyongyang–Hanoi split after the fall of Saigon. Cold War History, 22(2):173–194, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2021.1923698.

Dobson:2012:RRR

[Dob12] Alan Dobson. The rebellion of Ronald Reagan. Cold War History, 12 (2):369–371, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.686587.

Dockrill:2008:BN

[Doc08] Saki Ruth Dockrill. Book notes. Cold War History, 8(1):115– 130, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701791391.

Doose:2021:GPD

[Doo21] Katja Doose. A global problem in a divided world: climate change research during the late Cold War, 1972–1991. Cold War History, 21 (4):469–489, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1885377.

Dragomir:2012:PTH

[Dra12] Elena Dragomir. The perceived threat of hegemonism in Romania during the second détente. Cold War History, 12(1):111–134, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.510837.

Dragomir:2013:RPA

[Dra13] Elena Dragomir. Romania’s participation in the Agricultural Conference in Moscow, 2–3 February 1960. Cold War History, 13(3): 331–351, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.768068.

Drinot:2011:CRI

[Dri11] Paulo Drinot. A century of revolution: insurgent and counterinsurgent violence during Latin America’s long Cold War. Cold War History, 11 (3):483–484, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.599985.

Drogan:2013:BRD

[Dro13] Mara Drogan. Book review: Derek C. Maus, Unvarnishing Reality: Subversive Russian and American Cold War Satire (Columbia, SC: University of South Carolina Press, 2011), xvi + 247 pp. Cold War History, 13(2):283–284, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.789664.

Drogan:2015:NNG

[Dro15] Mara Drogan. The nuclear nation and the German question: an American reactor in West Berlin. Cold War History, 15(3):301– 319, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.959500. See erratum [Ano15c].

Dujardin:2007:GBB

[Duj07] Vincent Dujardin. Go-between: Belgium and Détente, 1961–73. Cold War History, 7(1):95–116, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740701197706.

Dulffer:2008:CWH

[Dül08] Jost Dülffer. Cold War history in Germany. Cold War History, 8 (2):135–156, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802018652.

Dulffer:2012:YEA

[Dül12] Jost Dülffer. Yalta 1945: Europe and America at the crossroads. Cold War History, 12(1):175–176, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.655447.

Dumbrell:2010:PBC

[Dum10] John Dumbrell. Piercing the bamboo curtain: tentative bridgebuilding to China during the Johnson years. Cold War History, 10 (2):279–280, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003711848.

Dumancic:2014:CWC

[Dum14a] Marko Dumancić. The Cold War’s cultural ecosystem: angry young men in British and Soviet cinema, 1953–1968. Cold War History, 14 (3):403–422, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.887684.

Dumancic:2014:SSC

[Dum14b] Marko Dumancić. The socialist sixties: crossing borders in the second world. Cold War History, 14(4):708–709, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.955693.

Dwan:2001:JME

[Dwa01] R. Dwan. Jean Monnet and the European Defence Community, 1950– 54. Cold War History, 1(3):141–160, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/713999932.

Easter:2005:KIP

[Eas05] David Easter. ‘Keep the Indonesian pot boiling’: Western covert intervention in Indonesia, October 1965–March 1966. Cold War History, 5(1):55–73, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
1468274042000283144.

Easter:2015:ASM

[Eas15] David Easter. Active Soviet military support for Indonesia during the 1962 West New Guinea crisis. Cold War History, 15(2): 201–220, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.995173.

Ellison:2023:WU

[ECH^{+}23] James Ellison, Michael Cox, Jussi M. Hanhimäki, Hope M. Harrison, N. Piers Ludlow, Angela Romano, Kristina Spohr, and Vladislav Zubok. The war in Ukraine. Cold War History, 23(1):121–206, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2162329.

Edelman:2017:FHN

[Ede17] Robert Edelman. The five hats of Nina Ponomareva: sport, shoplifting and the Cold War. Cold War History, 17(3):223–239, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2016.1261114.

Efimova:2005:SRC

[Efi05] Larisa M. Efimova. Stalin and the revival of the Communist Party of Indonesia. Cold War History, 5(1):107–120, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/1468274042000283162.

Eisenman:2018:CAC

[Eis18] Joshua Eisenman. Comrades-in-arms: the Chinese Communist Party’s relations with African political organisations in the Mao era, 1949–76. Cold War History, 18(4):429–445, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1440549.

Elie:2010:AGU

[Eli10] Jérôme B. Elie. Americans at the gate: the United States and refugees during the Cold War. Cold War History, 10(1):137–138, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903398904.

Ellison:2006:DGA

[Ell06] James Ellison. Defeating the general: Anglo–American relations, Europe and the NATO Crisis of 1966. Cold War History, 6(1):85–111, 2006. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740500395469.

Ellison:2009:STF

[Ell09] James Ellison. The strategic triangle: France, Germany, and the United States in the shaping of the new Europe. Cold War History, 9(3): 447–448, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903135454.

Ellison:2010:HNR

[Ell10] James Ellison. Heath, Nixon and the rebirth of the Special Relationship: Britain, the US and the EC, 1969–74. Cold War History, 10 (3):462–464, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.493720.

Ellison:2011:EFP

[Ell11] James Ellison. European foreign policy during the Cold War: Heath, Brandt, Pompidou and the dream of political unity. Cold War History, 11(3):477–478, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.599980.

Ellison:2014:CLB

[Ell14] James Ellison. A companion to Lyndon B. Johnson. Cold War History, 14(4):711–712, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.955695.

Elli:2015:CBG

[Ell15] Mauro Elli. Callaghan, the British government and the N-bomb controversy. Cold War History, 15(3):321–339, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2014.971016.

Ellis:2016:NMS

[Ell16] Stephen Ellis. Nelson Mandela, the South African Communist Party and the origins of Umkhonto we Sizwe. Cold War History, 16(1): 1–18, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1078315.

Elli:2022:CAG

[Ell22a] Mauro Elli. Civil aviation and the globalization of the Cold War. Cold War History, 22(1):123–125, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1951772.

Elli:2022:NPJ

[Ell22b] Mauro Elli. ‘Nuclear power is not just economics’: atomic energy and economic development in the Karachi Nuclear Power Plant Project (Kanupp), 1955–1965. Cold War History, 22(4):383–405, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2059071.

Emery:2010:TCW

[Eme10] Christian Emery. The transatlantic and Cold War dynamics of Iran sanctions, 1979–80. Cold War History, 10(3):371–396, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903218854.

Ewing:2019:CPC

[Ewi19] Cindy Ewing. The Colombo Powers: crafting diplomacy in the Third World and launching Afro–Asia at Bandung. Cold War History, 19 (1):1–19, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1500553.

Fainberg:2015:UWS

[Fai15] Dina Fainberg. Unmasking the wolf in sheep’s clothing: Soviet and American campaigns against the enemy’s journalists, 1946–1953. Cold War History, 15(2):155–178, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.978762.

Falk:2005:R

[Fal05] Andrew J. Falk. Reviews. Cold War History, 5(4):551–559, 2005. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740500296170.

Fardella:2013:MSK

[Far13] Enrico Fardella. Mao, Stalin and the Korean War: Trilateral Communist relations in the 1950s. Cold War History, 13(3):433–434, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2013.819643.

Farbol:2015:CCW

[Far15a] Rosanna Farbøl. Commemoration of a cold war: the politics of history and heritage at Cold War memory sites in Denmark. Cold War History, 15(4):471–490, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1028532.

Fardella:2015:MZC

[Far15b] Enrico Maria Fardella. Mao Zedong and the 1962 Cuban Missile Crisis. Cold War History, 15(1):73–88, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.971017. See corrigendum [?, ?].

Fardella:2017:SPC

[Far17] Enrico Fardella. A significant periphery of the Cold War: Italy– China bilateral relations, 1949–1989. Cold War History, 17(2):181– 197, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1093847.

Faraldo:2022:SLE

[Far22] Jose M. Faraldo. The story of Laura. Eastern Bloc surveillance of Spain in the Late Cold War (1967–1990). Cold War History, 22 (1):1–18, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1938551.

Fenemore:2020:VKU

[Fen20] Mark Fenemore. Victim of kidnapping or an unfortunate defector? The strange case of Otto John. Cold War History, 20(2):143–160, 2020. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2019.1689390.

Fortmann:2014:GOS

[FH14] Michel Fortmann and David G. Haglund. Of ghosts and other spectres: the Cold War’s ending and the question of the next ‘hegemonic’ conflict. Cold War History, 14(4):515–532, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950244.

Field:2012:BRC

[Fie12] Thomas C. Field Jr. Book review: Cuba y Estados Unidos, 1933–1959: del compromiso nacionalista al conflicto, by Vanni Pettina, Madria, Catarate, 2011, 279 pp. Cold War History, 12(3):563–565, 2012. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2012.702083.

Filitov:2006:GWB

[Fil06] Alexey Filitov. ‘Germany will be a bourgeois–democratic republic’: The new evidence from the personal file of Georgiy Malenkov. Cold War History, 6(4):549–557, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740600980806.

Filipink:2023:PFN

[Fil23] Richard M. Filipink. Political fallout: nuclear weapons testing and the making of a global environmental crisis. Cold War History, 23 (4):549–551, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2113010.

Fischer:2009:MSG

[Fis09] Thomas Fischer. ‘A mustard seed grew into a bushy tree’: The Finnish CSCE initiative of 5 May 1969. Cold War History, 9(2): 177–201, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802170982.

Fischer:2010:MBC

[Fis10] Beth A. Fischer. The military balance in the Cold War: US perceptions and policy, 1976–85. Cold War History, 10(2):281–282, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682741003711855.

Fischer:2021:BRC

[Fis21] Benjamin B. Fischer. Book review: Christian Ostermann, Between containment and rollback: the United States and the Cold War in Germany (Stanford: Stanford University Press, 2021), xix + 392 pp. Cold War History, 21(4):535–538, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1951761.

Fiedler:2015:SPS

[FM15] Anke Fiedler and Michael Meyen. The steering of the press in the socialist states of Eastern Europe: the German Democratic Republic (GDR) as a case study. Cold War History, 15(4):449–470, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2015.1028531.

Foglesong:2020:WRR

[Fog20] David Scott Foglesong. When the Russians really were coming: citizen diplomacy and the end of Cold War enmity in America. Cold War History, 20(4):419–440, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2020.1735368.

Foo:2009:FBC

[Foo09] Yee Wah Foo. Fu Bingchang, Chiang Kai-shek and Yalta. Cold War History, 9(3):389–409, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740902978995.

Footitt:2011:AFF

[Foo11] Hilary Footitt. American forces in France: Communist representations of US deployment. Cold War History, 11(1):85–98, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2011.545601.

Footitt:2013:HDC

[Foo13] Hilary Footitt. ‘A hideously difficult country’: British propaganda to France in the early Cold War. Cold War History, 13(2):153– 169, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.746663.

Forsberg:2019:IUS

[For19] Carl Forsberg. Iraq, the United States, and the long shadow of the Cold War. Cold War History, 19(4):457–476, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2019.1594784.

Forster:2021:TPP

[For21] Elisabeth Forster. Threatened by peace: the PRC’s peacefulness rhetoric and the ‘China’ representation question in the United Nations (1949–71). Cold War History, 21(4):411–427, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2019.1678028.

Franczak:2018:HRB

[Fra18] Michael Franczak. Human rights and basic needs: Jimmy Carter’s North–South dialogue, 1977–81. Cold War History, 18(4):447–464, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/14682745.
2018.1468437.

Franczak:2020:RFO

[Fra20] Michael Franczak. The rise and fall of OPEC in the twentieth century. Cold War History, 20(2):243–246, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1746008.

Freedman:2002:KBC

[Fre02] L. Freedman. Kennedy, Bush and crisis management. Cold War History, 2(3):1–14, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999964.

Freedman:2014:SPL

[Fre14] Lawrence Freedman. Stephen Pinker and the long peace: alliance, deterrence and decline. Cold War History, 14(4):657–672, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.950243.

Friedman:2010:SPD

[Fri10] Jeremy Friedman. Soviet policy in the developing world and the Chinese challenge in the 1960s. Cold War History, 10(2):247–272, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.481426.

Friedman:2017:SSA

[Fri17] Jeremy Friedman. The Sino–Soviet alliance: an international history. Cold War History, 17(1):104–106, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1226827.

Fujiangchuan:2023:WWD

[FS23] Rao Fujiangchuan and Yifan Shi. Word war during Patrick Hurley’s 1944 mission to China. Cold War History, 23(4):537–547, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2202350.

Fulbrook:2011:BBW

[Ful11] Mary Fulbrook. Behind the Berlin Wall: East Germany and the frontiers of power. Cold War History, 11(3):480–481, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.599983.

Furby:2009:AGR

[Fur09] Daniel Furby. Anglo–German relations during the labour governments 1964–70: NATO strategy, détente and European integration. Cold War History, 9(4):539–540, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903268495.

Fan:2023:RSI

[FZ23] Hongwei Fan and Yizheng Zou. Rethinking the security issue in the China–Burma territorial dispute. Cold War History, 23(1):45– 59, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2086860.

Gaddis:2007:RPL

[Gad07] John Lewis Gaddis. Response to Painter and Lundestad. Cold War History, 7(1):117–120, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740701197771. See [Pai06, Lun06].

Gala:2019:EWD

[Gal19] Marilena Gala. ‘The essential weaknesses of the December 1979 “Agreement”’: the White House and the implementing of the dual-track decision. Cold War History, 19(1):21–38, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2018.1462339.

Garcia:2013:CMU

[Gar13] Oscar J. Martin Garcia. A complicated mission: The United States and Spanish students during the Johnson administration. Cold War History, 13(3):311–329, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.746664.

Garbagni:2023:BCH

[Gar23] Giulia Garbagni. ‘Better to be a chicken’s head than an ox’s tail’: Japanese envoy diplomacy in the mediation of Konfrontasi (1965). Cold War History, 23(3):389–409, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2157819.

Gasbarri:2018:SOB

[Gas18] Flavia Gasbarri. From the sands of the Ogaden to Black Hawk Down: The end of the Cold War in the Horn of Africa. Cold War History, 18(1):73–89, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1364729.

Gasztold:2024:WFW

[Gas24] Przemysław Gasztold. Warsaw and the Fedayeen: Wars in the Middle East, secret arms deals and Polish relations with the Palestine Liberation Organisation, 1967–1976. Cold War History, 24(2):161–186, 2024. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2023.2269100.

Gatejel:2022:BHA

[Gat22] Luminita Gatejel. Bargaining for humanitarian aid across the Iron Curtain: Western relief workers in Romania in the late 1970s. Cold War History, 22(1):41–57, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2021.1893307.

Gerlini:2010:WDU

[Ger10] Matteo Gerlini. Waiting for Dimona: The United States and Israel’s development of nuclear capability. Cold War History, 10(2): 143–161, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903162300.

Gerits:2016:BCB

[Ger16] Frank Gerits. Bandung as the call for a better development project: US, British, French and Gold Coast perceptions of the Afro–Asian Conference (1955). Cold War History, 16(3):255–272, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2016.1189412.

Gerits:2019:PCT

[Ger19a] Frank Gerits. The postcolonial cultural transaction: rethinking the Guinea crisis within the French cultural strategy for Africa, 1958–60. Cold War History, 19(4):493–509, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2019.1576170.

Germuska:2019:BBC

[Ger19b] Pál Germuska. Balancing between the COMECON and the EEC: Hungarian elite debates on European integration during the long 1970s. Cold War History, 19(3):401–420, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1544972.

Gerth:2019:RRA

[Ger19c] Matthew Gerth. Revolutionaries for the Right: anticommunist internationalism and paramilitary warfare in the Cold War. Cold War History, 19(2):317–320, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2019.1601412.

Gfeller:2011:EVA

[Gfe11] Aurélie Élisa Gfeller. A European voice in the Arab world: France, the superpowers and the Middle East, 1970–74. Cold War History, 11 (4):659–676, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.504207.

Gheorghe:2013:AMR

[Ghe13] Eliza Gheorghe. Atomic maverick: Romania’s negotiations for nuclear technology, 1964–1970. Cold War History, 13(3):373–392, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.776542.

Gijs:2016:FRP

[Gij16] Anne-Sophie Gijs. Fighting the red peril in the Congo. Paradoxes and perspectives on an equivocal challenge to Belgium and the West (1947– 1960). Cold War History, 16(3):273–290, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2016.1163340.

Gill:2009:ABO

[Gil09] David Gill. The atomic bomb and the origins of the Cold War. Cold War History, 9(3):442–443, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903135413.

Gill:2010:HWC

[Gil10] David James Gill. Harold Wilson’s Cold War: the Labour government and East–West politics, 1964–1970. Cold War History, 10(3): 461–462, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.493721.

Ginor:2002:UYA

[Gin02] I. Ginor. ’Under the yellow Arab helmet gleamed blue Russian eyes’: Operation Kavkaz and the War of Attrition, 1969–70. Cold War History, 3(1):127–156, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999972.

Givens:2022:APS

[Giv22] Seth A. Givens. All part of the same struggle: Berlin’s role in German– US relations during the Lyndon Johnson presidency. Cold War History, 22(3):287–304, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2036719.

Gleijeses:2007:CIN

[Gle07] Piero Gleijeses. Cuba and the independence of Namibia. Cold War History, 7(2):285–303, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740701284215.

Gleijeses:2016:CPO

[Gle16] Piero Gleijeses. The CIA’s paramilitary operations during the cold war: an assessment. Cold War History, 16(3):291–306, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2016.1177513.

Gleijeses:2018:CWW

[Gle18] Piero Gleijeses. The Cold War: a world history. Cold War History, 18(2):248–250, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1449571.

Gloriant:2018:ACW

[Glo18] Frédéric Gloriant. To adapt to the Cold War bipolar order? Or to challenge it? Macmillan and de Gaulle’s rift in the face of the Second Berlin Crisis. Cold War History, 18(4):465–483, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2018.1434509.

Gnoinska:2017:SFS

[Gno17] Margaret K. Gnoinska. ‘Socialist friends should help each other in crises’: Sino–Polish relations within the Cold War dynamics, 1980–1987. Cold War History, 17(2):143–159, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1247157.

Gnoinska:2018:PCW

[Gno18] Margaret K. Gnoinska. Promoting the ‘China Way’ of communism in Poland and beyond during the Sino–Soviet Split: the case of Kazimierz Mijal. Cold War History, 18(3):343–359, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2017.1362394.

Goldman:2009:TBJ

[Gol09] Zachary K. Goldman. Ties that bind: John F. Kennedy and the foundations of the American–Israeli alliance. Cold War History, 9 (1):23–58, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802170941.

Goncalves:2019:BPP

[Gon19] Stéphanie Gonçalves. Ballet, propaganda, and politics in the Cold War: the Bolshoi Ballet in London and the Sadler’s Wells Ballet in Moscow, October–November 1956. Cold War History, 19(2):171–186, 2019. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2018.1468436.

Goodman:2007:DDB

[Goo07] Michael S. Goodman. The dog that didn’t bark: The Joint Intelligence Committee and warning of aggression. Cold War History, 7 (4):529–551, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701621739.

Ginor:2006:FBA

[GR06] Isabella Ginor and Gideon Remez. Un-finnished business: Archival evidence exposes the diplomatic aspect of the USSR’s pre-planning for the Six Day War. Cold War History, 6(3):377–395, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740600795568.

Granville:2010:HSW

[Gra10] Johanna Granville. Hungary, 101: Seven ways to avoid a revolution and Soviet invasion of Romania. Cold War History, 10(1):81–106, 2010. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740902978938.

Granieri:2011:EAA

[Gra11] Ronald J. Granieri. Eisenhower and Adenauer: alliance maintenance under pressure, 1953–1960. Cold War History, 11(2):286–287, 2011. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2011.569161.

Graham:2020:IIH

[Gra20a] Simon Graham. Internationalising the intelligence history of the Prague Spring. Cold War History, 20(3):293–310, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2019.1697238.

Gray:2020:MFT

[Gra20b] Geoffrey Gray. ‘In my file, I am two different people’: Max Gluckman and A. L. Epstein, the Australian National University, and Australian Security Intelligence Organisation, 1958–60. Cold War History, 20(1):59–76, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1575367.

Greenwood:2009:USB

[Gre09] Sean Greenwood. The United States, Britain and the transatlantic crisis: rising to the Gaullist challenge, 1963–68. Cold War History, 9 (4):537–538, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903268503.

Griesdorf:2008:NRC

[Gri08] Michael Jeffrey Griesdorf. Norms as a rhetorical competition: Soviet– American confrontations over Cuba, 1970–85. Cold War History, 8 (3):299–326, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802222098.

Grimm:2023:PPP

[Gri23] Kevin E. Grimm. Perils, promises and perspectives: nuclear weapons, atomic energy and the International Confederation of Free Trade Unions in the early Cold War. Cold War History, 23(4):495–515, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2219614.

Grossheim:2005:RDR

[Gro05] Martin Grossheim. ‘Revisionism’ in the Democratic Republic of Vietnam: New evidence from the East German archives. Cold War History, 5(4):451–477, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500284804.

Grosser:2010:FBW

[Gro10] Pierre Grosser. The fall of the Berlin Wall: the revolutionary legacy of 1989. Cold War History, 10(4):578–579, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2010.513513.

Grosser:2015:LCC

[Gro15] Pierre Grosser. Looking for the core of the Cold War, and finding a mirage? Cold War History, 15(2):245–252, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2015.1023580. See

response [Ano16b].

Grunbacher:2003:SIW

[Grü03] Armin Grünbacher. Sustaining the island: Western aid to 1950s West Berlin. Cold War History, 3(3):1–23, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740312331391598.

Gu:2022:UPH

[Gu22a] Jikun Gu. Unknown page of history: the dispatch of Chinese workers to the USSR (1954–1957). Cold War History, 22(4):437–457, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2050698.

Guvenc:2022:LTT

[GU22b] Serhat Guvenc and Mesut Uyar. Lost in translation or transformation? The impact of American aid on the Turkish military, 1947–60. Cold War History, 22(1):59–77, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2020.1866551.

Guan:2002:RAJ

[Gua02] A. C. Guan. Review article — the Johnson Administration and ’confrontation’. Cold War History, 2(3):111–128, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999960.

Guderzo:2004:JEI

[Gud04] Max Guderzo. Johnson and European integration: a missed chance for transatlantic power. Cold War History, 4(2):89–114, 2004. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740412331391825.

Guelke:2010:CWS

[Gue10] Adrian Guelke. Cold War in Southern Africa: white power, black liberation. Cold War History, 10(3):465–466, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2010.493713.

Guelke:2024:CPS

[Gue24] Adrian Guelke. A comparative perspective on South Africa’s end of the Cold War. Cold War History, 24(2):323–328, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2024.2317468.

Gupta:2019:FIR

[Gup19] Amit R. Das Gupta. The fateful Indian recognition of West Germany, 1949. Cold War History, 19(1):101–117, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2018.1454901.

Gustafson:2012:ACI

[Gus12] Kristian Gustafson. Allende’s Chile and the inter-American Cold War. Cold War History, 12(2):377–378, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.686593.

Guth:2021:CCC

[Gut21] Stefan Guth. From close call to close contacts: transsystemic technodiplomacy and cooperation in the civilian uses of nuclear power, 1963–79. Cold War History, 21(4):491–508, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1775588.

Haeussler:2015:ESS

[Hae15] Mathias Haeussler. European Summer School 2013 Best Paper Prize Winner: a ‘Cold War European’? Helmut Schmidt and European integration, c.1945–1982. Cold War History, 15(4):427–447, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.989841.

Hall:2020:MUM

[Hal20] Jonny Hall. Making the unipolar moment: U.S. foreign policy and the rise of the post-Cold War order. Cold War History, 20(1): 118–119, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2020.1700605.

Hamilton:2021:MTR

[Ham21] Peter E. Hamilton. Migration in the time of revolution: China, Indonesia and the Cold War. Cold War History, 21(2):235–242, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1898524.

Hanhimaki:2008:CGR

[Han08] Jussi M. Hanhimäki. Conservative goals, revolutionary outcomes: the paradox of détente. Cold War History, 8(4):503–512, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740802373586.

Hanhimaki:2011:ERH

[Han11] Jussi M. Hanhimäki. The eccentric Realist: Henry Kissinger and the shaping of American foreign policy. Cold War History, 11(2): 274–275, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569150.

Hanhimaki:2013:BRT

[Han13] Jussi M. Hanhimäki. Book review: The Transatlantic Century: Europe and America, 1890–2010 (New York and Cambridge: Cambridge University Press, 2012), 392 pp. Cold War History, 13(4):563–564, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2013.849018.

Hanhimaki:2014:RGW

[Han14] Jussi M. Hanhimäki. The (really) good war? Cold War nostalgia and American foreign policy. Cold War History, 14(4):673–683, 2014. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2014.950245.

Hanhimaki:2015:WOR

[Han15] Jussi M. Hanhimäki. World order: Reflections on the character of nations and the course of history. Cold War History, 15(2): 255–258, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1018478.

Hanhimaki:2017:BRN

[Han17] Jussi M. Hanhimäki. Book reviews: Niall Ferguson, Kissinger, vol. 1: 1923–1968: The Idealist (New York: Penguin Press, 2015), xix + 986pp. and Greg Grandin, Kissinger’s Shadow: The Long Reach of America’s Most Controversial Statesman (New York: Metropolitan Books, 2015), 270 pp. Cold War History, 17(1):99–102, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2016.1226816.

Harrison:2000:DSW

[Har00] H. Harrison. Driving the Soviets up the Wall: a super-ally, a superpower, and the building of the Berlin Wall, 1958–61. Cold War History, 1(1):53–74, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999909.

Harrison:2008:TSC

[Har08] Hope M. Harrison. Teaching and scholarship on the Cold War in the United States. Cold War History, 8(2):259–284, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740802018819.

Harmer:2012:BCW

[Har12a] Tanya Harmer. Brazil’s Cold War in the southern cone, 1970–1975. Cold War History, 12(4):659–681, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.641953.

Harmer:2012:PFC

[Har12b] Tanya Harmer. Paulo Freire and the Cold War politics of literacy. Cold War History, 12(3):562–563, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.702081.

Harmer:2015:IOD

[Har15] Tanya Harmer. The ideological origins of the dirty war: Fascism, populism, and dictatorship in Twentieth Century Argentina. Cold War History, 15(3):417–420, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2015.1051363.

Harrison:2021:SES

[Har21] Mark Harrison. Security empire: the secret police in communist Eastern Europe. Cold War History, 21(1):115–118, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1866558.

Harper:2023:LMT

[Har23] John L. Harper. Il ‘lodo Moro’: terrorismo e ragion di stato, 1969– 1986 [The ‘lodo Moro’: terrorism and reason of state, 1969–1986]. Cold War History, 23(2):331–333, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2097589.

Hatzivassiliou:2013:ANP

[Hat13] Evanthis Hatzivassiliou. Out-of-area: NATO perceptions of the Third World, 1957–1967. Cold War History, 13(1):67–88, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2012.717532.

Haynes:2018:ESP

[Hay18] Mike Haynes. The economic struggle for power in Tito’s Yugoslavia: from World War II to non-alignment. Cold War History, 18(3): 379–381, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1489996.

Hatley-Broad:2006:R

[HB06] Barbara Hatley-Broad. Reviews. Cold War History, 6(4):559– 577, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600980822.

He:2019:RME

[HCL19] Yanqing He, Kee Cheok Cheong, and Ran Li. Revisiting the Malayan Emergency: the China factor in the Baling peace talks. Cold War History, 19(4):529–547, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2018.1555242.

Hein:2011:RAO

[Hei11] Laura Hein. Revisiting America’s occupation of Japan. Cold War History, 11(4):579–599, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2010.524210.

Heimann:2017:DRI

[Hei17] Gadi Heimann. Divide and rule: Israel’s tactics regarding the Jerusalem question and America’s response, 1949–1950. Cold War History, 17 (1):21–38, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1247156.

Hershberg:2003:WAI

[Her03] James G. Hershberg. The war in Afghanistan and the Iran–Contra affair: Missing links? Cold War History, 3(3):23–48, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740312331391608.

Hershberg:2007:HSC

[Her07] James G. Hershberg. ‘High-spirited confusion’: Brazil, the 1961 Belgrade Non-Aligned Conference, and the limits of an ‘independent’ foreign policy during the High Cold War. Cold War History, 7(3): 373–388, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701474840.

Hershberg:2012:FPC

[Her12] James G. Hershberg. A fiery peace in a Cold War: Bernard Schriever and the ultimate weapon. Cold War History, 12(2):371–373, 2012. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2012.686588.

Herrala:2019:CSA

[Her19] Meri Elisabet Herrala. Challenges for Soviet–American collaboration in the Cold War: the capitalisation of pianist Sviatoslav Richter for American musical markets. Cold War History, 19(2):187–208, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2018.1551369.

Heuser:2009:SAF

[Heu09] Beatrice Heuser. Soviet Air Force theory, 1918–1945. Cold War History, 9(3):441–442, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903135389.

Heuser:2014:LBQ

[Heu14] Beatrice Heuser. Looking back: a quarter of a century after the Cold War. Cold War History, 14(4):455–459, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2014.950240.

Hwang:2015:RTN

[hHC15] Jae ho Hwang and Lyong Choi. Re-thinking normalisation between the ROK and the PRC in the early 1990s: the South Korean perspective. Cold War History, 15(4):557–578, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1019869.

Hill:2013:BRT

[Hil13a] Matthew Alan Hill. Book review: Over There: Living with the U.S. Military Empire from World War Two to the Present (Durham, NC: Duke University Press, 2010), xviii + 453 pp. Cold War History, 13 (4):567–568, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.849021.

Hill:2013:RKR

[Hil13b] Matthew Alan Hill. The right kind of revolution: Modernization, development, and U.S. foreign policy from the Cold War to the present. Cold War History, 13(1):136–137, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.762235.

Hill:2021:BRA

[Hil21] Christopher Hill. Book review: Archie Brown, The human factor: Gorbachev, Reagan and Thatcher, and the end of the Cold War. Cold War History, 21(3):381–383, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1917047.

Hirsch:2020:CWA

[Hir20] Philipp Hirsch. The Cold War, the Arab world, and West Germany’s ‘Mediterranean moment’, 1967–73. Cold War History, 20(2): 161–178, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1690460.

Hof:2013:BRA

[Hof13] Tobias Hof. Book review: Alessandro Orsini, Anatomy of the Red Brigades: The Religious Mind-Set of Modern Terrorists (Ithaca and London: Cornell University Press, 2011), 317 pp. Cold War History, 13(2):286–287, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.789665.

Holloway:2016:SUC

[Hol16] David Holloway. The Soviet Union and the creation of the International Atomic Energy Agency. Cold War History, 16(2):177–193, 2016. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2015.1124265.

Hopkins:2001:PCW

[Hop01] M. Hopkins. The price of Cold War partnership: Sir Oliver Franks and the British military commitment in the Korean War. Cold War History, 1(2):28–46, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999922.

Hopkins:2008:TRC

[Hop08] Michael F. Hopkins. Teaching and research on the Cold War in the United Kingdom. Cold War History, 8(2):241–258, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740802018793.

Hornsby:2016:PSK

[Hor16] Robert Hornsby. The post-Stalin Komsomol and the Soviet fight for Third World youth. Cold War History, 16(1):83–100, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2015.1078316.

Horovitz:2020:WWS

[Hor20] Liviu Horovitz. When the world seemed new: George H. W. Bush and the end of the Cold War. Cold War History, 20(1):119–121, 2020. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2019.1695564.

Howell:2021:IES

[How21] Jordan Howell. Itineraries of expertise: science, technology and the environment in Latin America’s Long Cold War. Cold War History, 21 (3):384–386, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1934076.

Hughes:2002:WSS

[Hug02] R. Gerald Hughes. ’We are not seeking strength for its own sake’: The British Labour Party, West Germany and the Cold War, 1951–64. Cold War History, 3(1):67–94, 2002. CODEN ???? ISSN 1468-2745 (print),

1743-7962 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/713999976.

Hughes:2004:BPT

[Hug04] Geraint Hughes. British policy towards Eastern Europe and the impact of the ‘Prague Spring’, 1964–68. Cold War History, 4(2): 115–139, 2004. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740412331391835.

Hughes:2006:GRB

[Hug06] Geraint Hughes. ‘Giving the Russians a bloody nose’: Operation Foot and Soviet espionage in the United Kingdom, 1964–71. Cold War History, 6(2):229–249, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740600650235.

Hull:2012:PSA

[Hul12] Christopher Hull. Parallel spheres: Anglo–American cooperation over Cuba, 1959–61. Cold War History, 12(1):51–68, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.506510.

Humalajoki:2020:TBC

[Hum20] Reetta Humalajoki. Tearing down the ‘buckskin curtain’: domestic policy-making and Indigenous intellectuals in the Cold War United States and Canada. Cold War History, 20(2):223–242, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2019.1673738.

Hunt:2011:VW

[Hun11] David Hunt. Vietnam at war. Cold War History, 11(2):282– 283, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569157.

Husain:2005:CAU

[Hus05] Aiyaz Husain. Covert action and US Cold War strategy in Cuba, 1961– 62. Cold War History, 5(1):23–53, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/1468274042000283135.

Hyun:2022:EBA

[Hyu22] Sinae Hyun. In the eyes of the beholder: American and Thai perceptions of the highland minority during the Cold War. Cold War History, 22(2):153–171, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1928077.

Iandolo:2012:RFS

[Ian12] Alessandro Iandolo. The rise and fall of the ‘Soviet Model of Development’ in West Africa, 1957–64. Cold War History, 12(4):683– 704, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.649255.

Ingimundarson:2001:TPR

[Ing01] V. Ingimundarson. Targeting the periphery: The role of Iceland in East German foreign policy, 1949–89. Cold War History, 1(3):113–140, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999929.

Jacobson:2016:AMS

[Jac16] Zachary Jonathan Jacobson. On the ‘arcane modern science of Kremlinology’ or the case of the vanishing birthdays. Cold War History, 16 (2):141–158, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1128897.

Jagtiani:2022:FAF

[Jag22] Sharinee L. Jagtiani. ‘Foreign armies are functioning on Asian soil’: India, Indonesian decolonisation and the onset of the Cold War (1945– 1949). Cold War History, 22(3):305–324, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2021.2021888.

Jarquin:2018:RCS

[Jar18] Mateo Cayetano Jarquín. Red Christmases: the Sandinistas, indigenous rebellion, and the origins of the Nicaraguan civil war, 1981–82. Cold War History, 18(1):91–107, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2017.1347637.

Jersild:2012:GBR

[Jer12] Austin Jersild. The great betrayal: Russian memories of the ‘great friendship’. Cold War History, 12(1):159–169, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.508493.

Jersild:2018:SES

[Jer18] Austin Jersild. Socialist exhibits and Sino–Soviet relations, 1950–60. Cold War History, 18(3):275–289, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1481047.

Johnson:2006:EIF

[Joh06] Edward Johnson. Early indications of a freeze: Greece, Spain and the United Nations, 1946–47. Cold War History, 6(1):43–61, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740500395436.

Johnson:2023:TAF

[Joh23] Julianne Johnson. Teaching anticommunism: Fred Schwarz and American postwar conservatism. Cold War History, 23(4):554–556, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2022.2121357.

Jong:2020:CWS

[Jon20] Willem Pieter Theodoor (Wim) De Jong. Cold War social science in action. The Ford Foundation and liberal adult education in the United States (1945–60). Cold War History, 20(4):399–418, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2019.1673737.

Joseph:2019:BCR

[Jos19] Gilbert M. Joseph. Border crossings and the remaking of Latin American Cold War Studies. Cold War History, 19(1):141–170, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2019.1557824. See

supplement [Cas20] and response [Jos20].

Joseph:2020:CCB

[Jos20] Gilbert M. Joseph. The continuing challenge of border crossing: a response to Marcelo Casals’ commentary. Cold War History, 20(3): 373–377, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2020.1762312. See [Jos19, Cas20].

Jun:2011:STU

[Jun11] Niu Jun. Strait talk: United States–Taiwan relations and the crisis with China. Cold War History, 11(2):288–289, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2011.569165.

Jurdem:2017:MHT

[Jur17] Laurence R. Jurdem. ‘That mad hatters tea party on the East river’: Conservative journals of opinion and the United Nations 1964–1981. Cold War History, 17(1):39–59, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1255197.

Kadercan:2014:BMF

[Kad14] Burak Kadercan. Busting myths — five books on The Cuban Missile Crisis in its fiftieth anniversary. Cold War History, 14(3): 445–451, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.917802.

Kalinovsky:2008:OPN

[Kal08] Artemy Kalinovsky. Old politics, new diplomacy: the Geneva accords and the Soviet withdrawal from Afghanistan. Cold War History, 8 (3):381–404, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802222213.

Kalinovsky:2011:ZCL

[Kal11] Artemy Kalinovsky. Zhivago’s children: the last Russian intelligentsia. Cold War History, 11(2):277–278, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.569152.

Kalinovsky:2014:VKM

[Kal14] Artemy M. Kalinovsky. Virus A: kak my zaboleli vtorzheniem v Afghanistan [Virus A: How we caught the disease of intervening in Afghanistan]. Cold War History, 14(2):287–289, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.893688.

Kanet:2006:SQE

[Kan06] Roger E. Kanet. The superpower quest for empire: The Cold War and Soviet support for ‘wars of national liberation’. Cold War History, 6 (3):331–352, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600795469.

Kaplan:2003:DNA

[Kap03] Lawrence S. Kaplan. The development of the NATO archives. Cold War History, 3(3):103–106, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740312331391638.

Kapitonova:2014:VSL

[Kap14] Natalia Kapitonova. Visit of Soviet leaders Nikita Khrushchev and Nicholas Bulganin to Britain in April 1956. Cold War History, 14 (1):127–152, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.783343.

Kaplan:2019:GLL

[Kap19] Lawrence S. Kaplan. General Lyman L. Lemnitzer and NATO, 1948– 69: a deferential leader. Cold War History, 19(3):323–341, 2019. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2018.1530633.

Karamouzi:2011:USM

[Kar11] Eirini Karamouzi. The United States and the making of modern Greece: history and power, 1950–1974. Cold War History, 11(2): 287–288, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569163.

Karl:2016:RCR

[Kar16] Robert A. Karl. Reading the Cuban revolution from Bogotá, 1957–62. Cold War History, 16(4):337–358, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1218848.

Karamouzi:2021:BRK

[Kar21] Eirini Karamouzi. Book review: Kiran Klaus Patel, Project Europe: a history, (Cambridge: Cambridge University Press, 2020), x + 379 pp. Cold War History, 21(1):113–119, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1846700.

Katsakioris:2024:ASM

[Kat24] Constantin Katsakioris. The African Student Movement in the Soviet Union during the 1960s: Pan-Africanism and Communism in the shadow of nation-states. Cold War History, 24(1):109–129, 2024. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2023.2295358.

Kayaoglu:2009:SID

[Kay09] Barin Kayaoglu. Strategic imperatives, democratic rhetoric: The United States and Turkey, 1945–52. Cold War History, 9(3):321– 345, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740902981403.

Karamouzi:2019:TPA

[KC19] Eirini Karamouzi and Dionysios Chourchoulis. Troublemaker or peacemaker? Andreas Papandreou, the Euromissile Crisis, and the policy of peace, 1981–86. Cold War History, 19(1):39–61, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2018.1497014.

Kecskes:2024:RTC

[Kec24] Gusztáv Kecskés. Refugee transport and the Cold War: the Intergovernmental Committee for European Migration (ICEM) and the Hungarian refugees of 1956. Cold War History, 24(1):1–22, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2217761.

Kedar:2018:END

[Ked18] Claudia Kedar. Economic neutrality during the Cold War: the World Bank, the United States, and Pinochet’s Chile, 1973–1977. Cold War History, 18(2):149–167, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2017.1420056.

Kelley:2010:PPD

[Kel10] John Robert Kelley. Practicing public diplomacy: a Cold War odyssey. Cold War History, 10(1):138–139, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903398870.

Keller:2022:RNC

[Kel22] Renata Keller. Research note: the Cuban Missile Crisis and a war of words in Argentina. Cold War History, 22(4):521–528, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2097594.

Kenez:2008:BR

[Ken08] Peter A. Kenez. Book reviews. Cold War History, 8(3):405– 417, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802222221.

Kent:2016:BRJ

[Ken16] Neil Kent. Book review: Jonathan Haslam, A New History of Soviet Intelligence: Near and Distant Neighbours (Oxford: Oxford University Press, 2015), 400 pp. Cold War History, 16(3):329–331, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2016.1186262.

Khan:2011:CWC

[Kha11] Sulmaan Wasif Khan. Cold War co-operation: New Chinese evidence on Jawaharlal Nehru’s 1954 visit to Beijing. Cold War History, 11 (2):197–222, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.494300.

Khalil:2017:UVT

[Kha17] Osamah F. Khalil. Us Versus them: The United States, radical Islam and the rise of the Green Threat. Cold War History, 17(3): 323–325, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1341379.

Khoo:2005:RRI

[Kho05] Nicholas Khoo. Realism redux: Investigating the causes and effects of Sino–US rapprochement. Cold War History, 5(4):529–549, 2005. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740500320517.

Kieninger:2020:DBD

[Kie20] Stephan Kieninger. Diplomacy beyond deterrence: Helmut Schmidt and the economic dimension of Ostpolitik. Cold War History, 20(2): 179–196, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1607308.

Kim:2010:SCC

[Kim10] Donggil Kim. Stalin and the Chinese Civil War. Cold War History, 10(2):185–202, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003619447.

Kim:2012:PWR

[Kim12] Donggil Kim. Prelude to war? The repatriation of Koreans from the Chinese PLA, 1949–50. Cold War History, 12(2):227–244, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.558896.

Kim:2016:NIM

[Kim16] Donggil Kim. New insights into Mao’s initial strategic consideration towards the Korean War intervention. Cold War History, 16(3): 239–254, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1190339.

Kim:2023:BEN

[Kim23] Seung Woo Kim. A brief encounter: North Korea in the Eurocurrency market, 1973–80. Cold War History, 23(3):411–429, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2023.2184805.

Kinkela:2013:EHC

[Kin13] David Kinkela. Environmental histories of the Cold War. Cold War History, 13(3):428–429, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2013.819644.

Kinney:2024:PDG

[Kin24] Brandon Kinney. “Poor devils”: German contributions to American flood relief and the Early Cold War. Cold War History, 24(2): 257–275, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2282009.

Kirby:2020:BTC

[Kir20] James Kirby. Between two Chinas and two Koreas: African agency and non-alignment in 1970s Botswana. Cold War History, 20(1): 21–38, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1632291.

Kitchen:2012:DAC

[Kit12] Nicholas Kitchen. Dean Acheson and the creation of an American world order. Cold War History, 12(4):706–707, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2012.737563.

Kind-Kovacs:2013:VLL

[KK13] Friederike Kind-Kovács. Voices, letters, and literature through the Iron Curtain: exiles and the (trans)mission of radio in the Cold War. Cold War History, 13(2):193–219, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.746666.

Klimentov:2023:SIL

[Kli23] Vassily A. Klimentov. In search of Islamic legitimacy: the USSR, the Afghan communists and the Muslim world. Cold War History, 23 (2):283–305, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2103114.

Mark:2010:VWT

[kM10] Chi kwan Mark. Vietnam War tourists: US Naval visits to Hong Kong and British–American–Chinese relations, 1965–1968. Cold War History, 10(1):1–28, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740902837001.

Mark:2017:EDX

[kM17] Chi kwan Mark. To ‘educate’: Deng Xiaoping in capitalism: Thatcher’s visit to China and the future of Hong Kong in 1982. Cold War History, 17(2):161–180, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1094058.

Knusel:2020:WOR

[Knü20] Ariane Knüsel. ‘White on the outside but red on the inside’: Switzerland and Chinese intelligence networks during the Cold War. Cold War History, 20(1):77–94, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2019.1575368.

Kochavi:2008:JCB

[Koc08a] Noam Kochavi. Joining the conservative brotherhood: Israel, President Nixon, and the political consolidation of the ‘special relationship’, 1969–73. Cold War History, 8(4):449–480, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740802373560.

Kochavi:2008:RDN

[Koc08b] Noam Kochavi. Researching détente: new opportunities, contested legacy. Cold War History, 8(4):419–425, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740802373511.

Kolenovska:2018:SCC

[Kol18] Daniela Kolenovska. Sino–Czechoslovak cooperation on agricultural cooperatives: the twinning project. Cold War History, 18(3):291– 306, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1387775.

Komornicka:2020:WSR

[Kom20] Aleksandra Komornicka. Winner of the Saki Ruth Dockrill Memorial Prize ‘the unity of Europe is inevitable’: Poland and the European Economic Community in the 1970s. Cold War History, 20(4): 483–501, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2020.1757651.

Konieczna:2021:NTF

[Kon21] Anna Konieczna. Nuclear twins: French–South African strategic cooperation (1964–79). Cold War History, 21(3):283–300, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2020.1823968.

Kosicki:2011:MDR

[Kos11] Piotr H. Kosicki. Meeting the demands of reason: the life and thought of Andrei Sakharov. Cold War History, 11(3):481–482, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.599984.

Kramer:2011:BRP

[Kra11] Mark Kramer. Book review: Prager Frühling: das internationale Krisenjahr 1968. [Prague Spring: the international crisis year of 1968]. 2 vols., vol. 1: Beiträge, vol. 2: Dokumente, Cologne, Böhlau Verlag, 2008, 1296 and 1589 pp. Cold War History, 11(4):679–682, 2011. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2011.617432.

Kraus:2012:BRE

[Kra12] Charles Kraus. A border region ‘exuded with militant friendship’: Provincial narratives of China’s participation in the First Indochina War, 1949–1954. Cold War History, 12(3):495–514, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2011.627919.

Kraus:2014:SSS

[Kra14] Charles Kraus. To die on the steppe: Sino–Soviet–American relations and the Cold War in Chinese Central Asia, 1944–1952. Cold War History, 14(3):293–313, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2013.871262.

Krige:2015:EIP

[Kri15] John Krige. Euratom and the IAEA: the problem of self-inspection. Cold War History, 15(3):341–352, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.999046.

Kuisong:2000:SSB

[Kui00] Y. Kuisong. The Sino–Soviet border clash of 1969: From Zhenbao Island to Sino–American rapprochement. Cold War History, 1(1):21–52, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/713999906.

Kulczycki:2022:PWP

[Kul22] John J. Kulczycki. Poland 1945: war and peace. Cold War History, 22(2):237–239, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.2002244.

Kusters:2002:DKG

[Küs02] H. Küsters. Document — the Kohl–Gorbachev meetings in Moscow and in the Caucasus, 1990. Cold War History, 2(2):195–235, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999951.

Krol:2017:IOE

[KW17] Joanna Król and Teresa G. Wojcik. The “Ideological offensive” in education: the portrayal of the United States in secondary curricula and textbooks in Poland during the Stalinist Period (1948–1956). Cold War History, 17(3):299–319, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2017.1284202.

Leon-Aguinaga:2021:DUM

[LAGE21] Pablo León-Aguinaga and Lorenzo Delgado Gómez-Escalonilla. The deployment of US military assistance to Spain in the 1950s: limited modernisation and strategic dependence. Cold War History, 21 (1):55–70, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1492554.

Lahey:2013:TGR

[Lah13] Daniel James Lahey. The Thatcher government’s response to the Soviet invasion of Afghanistan, 1979–1980. Cold War History, 13 (1):21–42, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.721355.

Lajeunesse:2013:VPR

[Laj13] Adam Lajeunesse. A very practical requirement: under-ice operations in the Canadian Arctic, 1960–1986. Cold War History, 13(4): 507–524, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.727800.

Lak:2021:BRW

[Lak21] Martijn Lak. Book review: West Germany and Israel: foreign relations, domestic politics, and the Cold War, 1965–1974. Carole Fink (Cambridge: Cambridge University Press, 2019), vii–iv + 349 pp. Cold War History, 21(1):118–119, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1844692.

Lane:2008:I

[Lan08] Ann Lane. Introduction. Cold War History, 8(2):131–133, 2008. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740802018629.

Langstaff:2022:SJN

[Lan22] Alexander Langstaff. Soviet judgment at Nuremberg: a new history of the international military tribunal after World War II. Cold War History, 22(3):377–380, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2026282.

Laron:2010:PFS

[Lar10] Guy Laron. Playing with fire: The Soviet–Syrian–Israeli triangle, 1965– 1967. Cold War History, 10(2):163–184, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740902871869.

Larmer:2013:LCT

[Lar13] Miles Larmer. Local conflicts in a transnational war: the Katangese gendarmes and the Shaba wars of 1977–78. Cold War History, 13 (1):89–108, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.727801.

Laron:2014:NGH

[Lar14] Guy Laron. Nasser’s gamble: how intervention in Yemen caused the Six-Day War and the decline of Egyptian power. Cold War History, 14 (3):451–452, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.917801.

Lash:2024:TTW

[Las24] Christopher Lash. Taking Third World solidarity with a pinch of salt: Socialist Poland’s policies towards 1960s Mali. Cold War History, 24 (2):187–208, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2269850.

Latham:2009:BR

[Lat09] Michael E. Latham. Book reviews. Cold War History, 9(2):287– 300, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740902884664.

Latham:2011:GAM

[Lat11] Robert E. Latham. The great American mission: modernization and the construction of an American world order. Cold War History, 11 (2):272–273, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569148.

Lawson:2011:CWM

[Law11] Sean Lawson. Cold War military systems science and the emergence of a nonlinear view of war in the US military. Cold War History, 11 (3):421–440, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.494302.

Leake:2023:WBB

[Lea23] Elisabeth Leake. The West and the birth of Bangladesh: foreign policy in the face of mass atrocity. Cold War History, 23(3):467–469, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2022.2109289.

Lerner:2011:FPL

[Ler11] Mitchell Lerner. The foreign policy of Lyndon B. Johnson: the United States and the world, 1963–69. Cold War History, 11(3): 475–477, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.599979.

Levey:2008:AAU

[Lev08] Zach Levey. Anatomy of an airlift: United States military assistance to Israel during the 1973 war. Cold War History, 8(4):481–501, 2008. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740802373552.

Lever:2013:DPV

[Lev13a] Paul Lever. The dissent papers: The voices of diplomats in the Cold War and beyond. Cold War History, 13(3):427–428, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2013.819645.

Levering:2013:MME

[Lev13b] Ralph Levering. Mental maps in the early Cold War era, 1945–1968. Cold War History, 13(1):131–132, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.762236.

Lever:2014:CWG

[Lev14] Paul Lever. The Cold War: the golden age of arms control. Cold War History, 14(4):501–513, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950246.

Li:2011:IDM

[Li11] Mingjiang Li. Ideological dilemma: Mao’s China and the Sino–Soviet split, 1962–63. Cold War History, 11(3):387–419, 2011. CODEN ????

ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.498822.

Liberati:2011:WHC

[Lib11] Luigi Bruti Liberati. Witch-hunts and Corriere della Sera. A conservative perception of American political values in Cold War Italy: The 1950s. Cold War History, 11(1):69–83, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2011.545599.

Lien:2005:CQN

[Liê05] Trân Thi Liên. The Catholic question in North Vietnam: From Polish sources, 1954–56. Cold War History, 5(4):427–449, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740500284747.

Link:2013:BRC

[Lin13] Miles Link. Book review: Cold War Literature: Writing the Global Conflict (London and New York: Routledge, 2009 [pbk]), 272 pp. Cold War History, 13(4):564–565, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.849023.

Liu:2018:DBS

[Liu18] Lei Liu. ‘Dog-beating stick’: General Zhang Aiping’s contribution to the modernisation of China’s nuclear force and strategy since 1977. Cold War History, 18(4):485–501, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1434507.

Lounasmeri:2020:CBE

[LK20] Lotta Lounasmeri and Jukka Kortti. Campaigning between East and West: Finland and the Cold War in the presidential campaign films of urho kekkonen. Cold War History, 20(3):329–348, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2018.1532996.

Locher:2009:KJN

[Loc09a] Anna Locher. Kennedy, Johnson and NATO: Britain, America and the dynamics of alliance, 1962–68. Cold War History, 9(3): 446–447, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903135447.

Locher:2009:NAT

[Loc09b] Anna Locher. The North Atlantic Treaty Organization, 1948–1957: community or alliance? Cold War History, 9(3):444–445, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903135439.

Lodevole:2010:WCE

[Lod10] Matteo Lodevole. The Western Communists and the European military build-up, 1949–50: a preventive strategy. Cold War History, 10 (2):203–228, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903162383.

Long:2024:IUY

[Lon24] Stephen Long. ‘Inexpensive to us and yet very valuable to the impoverished Albanian people’: Covert foreign aid and the Anglo– American subversion of Albania, 1951–55. Cold War History, 24(2): 277–297, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2268537.

Lopes:2013:BRF

[Lop13] Rui Lopes. Book review: Foreign Intervention in Africa: From the Cold War to the War on Terror (Cambridge: Cambridge University Press, 2013), xviii + 267 pp. Cold War History, 13(4):568–570, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.849025.

Lopes:2022:RHC

[Lop22] Rui Lopes. Retconning the history of covert operations: spy comics at the end of the Cold War. Cold War History, 22(2):215–236, 2022. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2021.1933951.

Lorenzini:2009:GO

[Lor09] Sara Lorenzini. Globalising Ostpolitik. Cold War History, 9(2): 223–242, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802188166.

Lorenzini:2024:BRR

[Lor24] Sara Lorenzini. Book review: Ripe for revolution: building socialism in the Third World. Jeremy Friedman, ed., (Cambridge, MA:

Harvard University Press, 2022), 368 pp. Cold War History, 24(2): 356–358, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2171351.

Loth:2001:MPW

[Lot01] W. Loth. Moscow, Prague and Warsaw: Overcoming the Brezhnev Doctrine. Cold War History, 1(2):103–118, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/713999924.

Loth:2003:REG

[Lot03] W. Loth. Review essay — general views on the Cold War. Cold War History, 3(2):157–165, 2003. CODEN ???? ISSN 1468-2745 (print),

1743-7962 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/713999985.

Loth:2004:OSN

[Lot04] Wilfried Loth. The origins of Stalin’s note of 10 March 1952. Cold War History, 4(2):66–88, 2004. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740412331391815.

Loth:2009:EDE

[Lot09] Wilfried Loth. The emergence of détente in Europe: Brandt, Kennedy and the formation of Ostpolitik. Cold War History, 9(3): 448–449, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903135470.

Loth:2010:GQS

[Lot10] Wilfried Loth. The German question from Stalin to Khrushchev: The meaning of new documents. Cold War History, 10(2):229–245, 2010. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740903065438.

Lott:2019:EWW

[Lot19] Gaia Lott. East–West and West–West competition in the Great Lakes Region, 1975–85. Cold War History, 19(4):511–528, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2018.1559830.

Loureiro:2017:SSM

[Lou17a] Felipe P. Loureiro. Securing sex: morality and repression in the making of Cold War Brazil. Cold War History, 17(3):325–327, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2017.1341419.

Loureiro:2017:APP

[Lou17b] Felipe Pereira Loureiro. The Alliance for Progress and President João Goulart’s Three-Year Plan: the deterioration of U.S.–Brazilian relations in Cold War Brazil (1962). Cold War History, 17(1):61–79, 2017. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2016.1254620.

Lowry:2007:IAC

[Low07] Donal Lowry. The impact of anti-communism on white Rhodesian political culture, ca.1920s–1980. Cold War History, 7(2):169– 194, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701284108.

Laakkonen:2016:CWE

[LPT16] Simo Laakkonen, Viktor Pál, and Richard Tucker. The Cold War and environmental history: complementary fields. Cold War History, 16(4):377–394, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1248544.

Lankov:2019:PCR

[LS19] Andrei Lankov and Igor Selivanov. A peculiar case of a runaway ambassador: Yi Sang-Cho’s defection and the 1956 crisis in North Korea. Cold War History, 19(2):233–251, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1507022.

Lynn:2024:ECW

[LS24] Katalin Kádár Lynn and Mark Stout. Early Cold War intelligence paper mills: the case of the Association of Hungarian Veterans. Cold War History, 24(1):23–44, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2023.2191948.

Ludlow:2001:RRN

[Lud01] P. Ludlow. Research report — no longer a closed shop: Post-1945 research in the French archives. Cold War History, 2(1):158–163, 2001.

CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999937.

Ludlow:2009:BRM

[Lud09] Piers Ludlow. Book review: La mésentente apprivoisée: de Gaulle et les Allemands, 1963–1969 [Disagreement tamed: de Gaulle and the Germans, 1963–1969], by Benedikt Schoenborn, Paris, Presses Universitaires de France, 2007, 431 pp. Cold War History, 9(3):449–451, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903135488.

Ludlow:2011:SAU

[Lud11] Piers Ludlow. The strained alliance: US–European relations from Nixon to Carter. Cold War History, 11(3):478–480, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2011.599981.

Lundestad:2000:IOM

[Lun00] G. Lundestad. ’Imperial overstretch’, Mikhail Gorbachev, and the end of the Cold War. Cold War History, 1(1):1–20, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/713999912.

Lunak:2003:KBC

[Lun03] P. Lunák. Khrushchev and the Berlin Crisis: Soviet brinkmanship seen from inside. Cold War History, 3(2):53–82, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999988.

Lundestad:2006:CWA

[Lun06] Geir Lundestad. The Cold War according to John Gaddis. Cold War History, 6(4):535–542, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740600979303. See response [Gad07].

Luszczykiewicz:2023:PIT

[Lus23] Antonina Luszczykiewicz. Pirates and imperialists: Taiwan and the United States in the Polish communist press, 1953–1955. Cold War History, 23(4):517–535, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2023.2206650.

Luthi:2007:PRC

[Lüt07] Lorenz M. Lüthi. The People’s Republic of China and the Warsaw Pact Organization, 1955–63. Cold War History, 7(4):479–494, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740701621762.

Luthi:2009:OPD

[Lüt09] Lorenz M. Lüthi. The origins of proletarian diplomacy: The Chinese attack on the American Embassy in the Soviet Union, 4 March 1965. Cold War History, 9(3):411–426, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740902910501.

Luthi:2017:SCW

[Lüt17] Lorenz Lüthi. Shadow Cold War: The Sino–Soviet competition for the Third World. Cold War History, 17(2):200–201, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2017.1287241.

Macekura:2011:LGC

[Mac11] Stephen Macekura. The limits of the global community: The Nixon administration and global environmental politics. Cold War History, 11 (4):489–518, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.498821.

Macqueen:2022:SZP

[Mac22] Ian Macqueen. Shaka Zulu in the Polish People’s Republic (PRL): exploring South African–Polish links in the late Cold War. Cold War History, 22(3):265–286, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2022.2027913.

Maddrell:2005:WWD

[Mad05] Paul Maddrell. What we have discovered about the Cold War is what we already knew: Julius Mader and the Western secret services during the Cold War. Cold War History, 5(2):235–258, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740500062127.

Maguire:2009:DNY

[Mag09] Lori Maguire. The destruction of New York City: a recurrent nightmare of American Cold War cinema. Cold War History, 9(4):

513–524, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903198247.

Maguire:2013:OWB

[Mag13] Lori Maguire. One world, big screen: Hollywood, the Allies and World War II. Cold War History, 13(3):430–431, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.819647.

Maguire:2017:HMV

[Mag17] Lori Maguire. Hearts, minds, voices: US Cold War public diplomacy and the formation of the Third World. Cold War History, 17 (4):448–450, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1382979.

Maguire:2019:BRJ

[Mag19] Lori Maguire. Book review: Julian Jackson, A certain idea of France: the life of Charles de Gaulle (London: Allen Lane, 2018), xl + 887 pp. Cold War History, 19(2):315–317, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2019.1601411.

Maguire:2022:BRA

[Mag22] Lori Maguire. Book review: Une armée de diplomates: Les militaires américains et la France, 1944–1967. François Doppler-Speranza, (Strasbourg: Presses Universitaires de Strasbourg, 2021), 388 pp. Cold War History, 22(4):545–546, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2101975.

Major:2003:FPC

[Maj03] Patrick Major. Future perfect? Communist science fiction in the Cold War. Cold War History, 4(1):71–96, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740312331391744.

Major:2013:LBC

[Maj13] Patrick Major. Listening behind the curtain: BBC broadcasting to East Germany and its Cold War echo. Cold War History, 13(2): 255–275, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.746840.

Maj:2021:PDG

[Maj21] Ewa Maj. Poland’s double game strategy during the US–Iranian crises (1979–88). Cold War History, 21(2):159–177, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2019.1581767.

Manchanda:2009:WTS

[Man09] Arnav Manchanda. When truth is stranger than fiction: the Able Archer incident. Cold War History, 9(1):111–133, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740802490315.

Mandler:2013:BRM

[Man13] Peter Mandler. Book review: Mark Solovey and Hamilton Cravens (eds), Cold War Social Science: Knowledge Production, Liberal Democracy, and Human Nature (New York: Palgrave Macmillan, 2012). Cold War History, 13(2):281–283, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.789666.

Manchester:2021:CKW

[Man21] Margaret M. Manchester. Charlie Kersten’s war: a Catholic crusader goes to Congress. Cold War History, 21(2):121–138, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2020.1724100.

Maragkou:2009:FNS

[Mar09] Konstantina Maragkou. Favouritism in NATO’s southeastern flank: The case of the Greek colonels, 1967–74. Cold War History, 9(3): 347–366, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802490356.

Marr:2012:VDC

[Mar12] David G. Marr. Vietnam declassified: the CIA and counterinsurgency. Cold War History, 12(2):368–369, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.686585.

Martini:2016:WFP

[Mar16] Edwin A. Martini. World on fire: the politics of napalm in the Global Cold War. Cold War History, 16(4):463–481, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2015.1080242.

Marku:2017:CAC

[Mar17] Ylber Marku. China and Albania: the Cultural Revolution and Cold War relations. Cold War History, 17(4):367–383, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2017.1307179.

Martin:2019:BRG

[Mar19] Garret Martin. Book review: Georges-Henri Soutou, La Guerre Froide de la France, 1941–1990 (Paris: Tallandier, 2018), 592 pp. Cold War History, 19(4):590–592, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2019.1678818.

Marku:2021:SIC

[Mar21] Ylber Marku. Stories from the international communist movement: the Chinese front in Europe and the limits of the anti-revisionist struggle. Cold War History, 21(2):139–157, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2020.1803840.

Marmon:2022:ORC

[Mar22a] Brooks Marmon. Operation Refugee: the Congo Crisis and the end of humanitarian imperialism in Southern Rhodesia, 1960. Cold War History, 22(2):131–152, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2021.1933950.

Marmon:2022:RNN

[Mar22b] Brooks Marmon. Research notes: Negotiating South African ministerial archives (Defence and Foreign Affairs). Cold War History, 22 (3):359–362, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2032203.

Maricic:2023:BBY

[Mar23] Alan Maricić. Breaking down bipolarity: Yugoslavia’s foreign relations during the Cold War. Cold War History, 23(3):465–467, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2104444.

Mastny:2002:DLC

[Mas02] V. Mastny. Diplomacy and the legacy of the Cold War: Post-11 September. Cold War History, 2(3):15–28, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999959.

Mastny:2009:EEE

[Mas09] Vojtech Mastny. Eastern Europe and the early prospects for EC/EU and NATO membership. Cold War History, 9(2):203–221, 2009. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740802170834.

Mastny:2014:CWR

[Mas14] Vojtech Mastny. The Cold War in retrospect: too early to tell? Cold War History, 14(4):487–499, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950247.

Matlock:2010:SCP

[Mat10] Jack F. Matlock Jr. 1989: the struggle to create post-cold war Europe. Cold War History, 10(4):575–578, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2010.513512.

Matray:2011:KWS

[Mat11] James I. Matray. Korea’s war at 60: a survey of the literature. Cold War History, 11(1):99–129, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.545603.

Matz:2015:SUE

[Mat15] Johan Matz. Sweden, the USSR and the early Cold War 1944–47: declassified encrypted cables shed new light on Soviet diplomatic reporting about Sweden in the aftermath of World War II. Cold War History, 15(1):27–48, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.920825.

Mavrodin:2021:ESS

[Mav21] Corina Mavrodin. European Summer School 2017 Best Paper Prize Winner: Courting the non-aligned: Romania, petro-diplomacy and the global Cold War. Cold War History, 21(2):179–195, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2019.1675964.

Mawby:2013:MSG

[Maw13] Spencer Mawby. Mr. Smith goes to Vienna: Britain’s Cold War in the Caribbean 1951–1954. Cold War History, 13(4):541–561, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2012.756474.

Mazov:2003:UFI

[Maz03] Sergei Mazov. The USSR and the former Italian colonies, 1945–50. Cold War History, 3(3):49–78, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740312331391618.

Mazov:2007:SAG

[Maz07] Sergei Mazov. Soviet aid to the Gizenga Government in the former Belgian Congo (1960–61) as reflected in Russian archives. Cold War History, 7(3):425–437, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701474873.

McAlexander:2011:CMU

[McA11] Richard J. McAlexander. Couscous Mussolini: US perceptions of Gamal Abdel Nasser, the 1958 intervention in Lebanon and the origins of the US–Israeli special relationship. Cold War History, 11(3): 363–385, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.482960.

McDonnell:2012:DHR

[McD12] Timothy McDonnell. The dead hand: Reagan, Gorbachev and the untold story of the Cold War arms race. Cold War History, 12(2): 373–374, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.686589.

McKercher:2012:MSP

[McK12] Asa McKercher. ‘The most serious problem’? Canada–US relations and Cuba, 1962. Cold War History, 12(1):69–88, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.562202.

McMahon:2006:RA

[McM06] Robert McMahon. Review article. Cold War History, 6(2):251– 260, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600650243.

McMahon:2010:AME

[McM10] Robert J. McMahon. Apocalypse management: Eisenhower and the discourse of national insecurity. Cold War History, 10(2):277– 278, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003711863.

McPherson:2021:SBE

[McP21] Alan McPherson. Strange bedfellows at the end of the Cold War: the Letelier assassination, human rights, and state sovereignty. Cold War History, 21(1):1–18, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2019.1583212.

Mourlon-Druol:2010:FEL

[MD10] Emmanuel Mourlon-Druol. Filling the EEC leadership vacuum? The creation of the European Council in 1974. Cold War History, 10(3): 315–339, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003765430.

Melson:2007:AR

[Mel07] Charles D. Melson. Archives report. Cold War History, 7(2): 327–331, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701284371.

Meyskens:2021:ECW

[Mey21] Covell Meyskens. Experiencing the Cold War at Shanghai’s secret military industrial complex. Cold War History, 21(4):429–447, 2021. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2020.1842876.

Mikkonen:2020:IFG

[Mik20] Simo Mikkonen. Interference or friendly gestures? Soviet cultural diplomacy and Finnish elections, 1945–56. Cold War History, 20(3): 349–365, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1686613.

Miller:2012:TFA

[Mil12] Jamie Miller. Things fall apart: South Africa and the collapse of the Portuguese Empire, 1973–74. Cold War History, 12(2):183– 204, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.649946.

Milford:2022:RSM

[Mil22] Ismay Milford. Revolutionary state-making in Dar es Salaam: African liberation and the global Cold War, 1961–1974. Cold War History, 22 (3):380–382, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2030899.

Minami:2016:REE

[Min16] Kazushi Minami. Re-examining the end of Mao’s revolution: China’s changing statecraft and Sino–American relations, 1973–1978. Cold War History, 16(4):359–375, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1218473.

Mistry:2006:CPW

[Mis06] Kaeten Mistry. The case for political warfare: Strategy, organization and US involvement in the 1948 Italian election. Cold War History, 6 (3):301–329, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600795451.

Mistry:2012:CAC

[Mis12] Kaeten Mistry. Confronting America: the Cold War between the United States and the communists in France and Italy. Cold War History, 12 (1):179–180, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.655455.

Mitchell:2007:TCW

[Mit07] Nancy Mitchell. Tropes of the Cold War: Jimmy Carter and Rhodesia. Cold War History, 7(2):263–283, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740701284181.

Mitchell:2017:STR

[Mit17] Nancy Mitchell. A superpower transformed: The remaking of American foreign relations in the 1970s. Cold War History, 17(2):198– 200, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1287242.

Mitchell:2018:CWW

[Mit18a] Nancy Mitchell. The Cold War: a world history. Cold War History, 18(2):239–240, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1449567.

Mitter:2018:CWW

[Mit18b] Rana Mitter. The Cold War: a world history. Cold War History, 18 (2):245–247, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1449570.

Major:2003:EEW

[MM03] Patrick Major and Rana Mitter. East is East and West is West? Towards a comparative socio-cultural history of the Cold War. Cold War History, 4(1):1–22, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740312331391714.

Menetrey-Monchau:2005:MIE

[MM05] Cécile Menétrey-Monchau. The Mayaguez incident as an epilogue to the Vietnam War and its reflection of the post-Vietnam political equilibrium in Southeast Asia. Cold War History, 5(3):337–367, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740500222127.

Murashko:2001:DSNa

[MN01a] G. Murashko and A. Noskova. DOCUMENT — Stalin and the national-territorial controversies in Eastern Europe, 1945–47. Cold War History, 1(3):161–172, 2001. CODEN ???? ISSN 1468-2745 (print),

1743-7962 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/713999933.

Murashko:2001:DSNb

[MN01b] G. Murashko and A. Noskova. Document — Stalin and the nationalterritorial controversies in Eastern Europe, 1945–47 (part 2). Cold War History, 2(1):145–157, 2001. CODEN ???? ISSN 1468-2745 (print),

1743-7962 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/713999939.

MacQueen:2010:GMB

[MO10] Norrie MacQueen and Pedro Aires Oliveira. ‘Grocer meets butcher’: Marcello Caetano’s London visit of 1973 and the last days of Portugal’s Estado Novo. Cold War History, 10(1):29–50, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740902764551.

Moores:2002:NOS

[Moo02] S. Moores. ’Neutral on our side’: US policy towards Sweden during the Eisenhower Administration. Cold War History, 2(3):29–62, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999963.

Moro:2014:IAE

[Mor14a] Renato Moro. Ideologies in the age of extremes: Liberalism, Conservatism, Communism, Fascism 1914–1991. Cold War History, 14 (4):706–708, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.955692.

Morrison:2014:ICW

[Mor14b] John N. L. Morrison. Intelligence in the Cold War. Cold War History, 14(4):575–591, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.950248.

Moser:2023:IPH

[Mös23] Robin E. Möser. Inspectors for peace: a history of the International Atomic Energy Agency. Cold War History, 23(3):469–472, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2022.2112010.

Moulton:2015:BTO

[Mou15] Aaron Coy Moulton. Building their own Cold War in their own backyard: the transnational, international conflicts in the greater Caribbean basin, 1944–1954. Cold War History, 15(2):135–154, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2014.995172.

Mueller:2006:SAN

[Mue06] Wolfgang Mueller. Stalin and Austria: New evidence on Soviet policy in a secondary theatre of the Cold War, 1938–53/55. Cold War History, 6(1):63–84, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500395444.

Muehlenbeck:2013:ENE

[Mue13] Philip Muehlenbeck. Ethnicity, nationalism and the European Cold War. Cold War History, 13(3):436–437, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2013.819649.

Mueller:2024:BRF

[Mue24] Brian S. Mueller. Book review: Freedom on the offensive: human rights, democracy promotion, and US interventionism in the late Cold War. William Michael Schmidli, (Ithaca: Cornell University Press, 2022), xii + 324 pp. Cold War History, 24(2):353–355, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2158713.

Mukhina:2006:NRF

[Muk06] Irina Mukhina. New revelations from the former Soviet archives: The Kremlin, the Warsaw Uprising, and the coming of the Cold War. Cold War History, 6(3):397–411, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740600795584.

Muminov:2017:IRC

[Mum17] Sherzod Muminov. From imperial revenants to Cold War victims: ‘red repatriates’ from the Soviet Union and the making of the new Japan, 1949–1952. Cold War History, 17(4):425–442, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2017.1324849.

Musto:2021:QSC

[Mus21] Ryan A. Musto. ‘A question of survival’: Canada and the Rapacki Plan for the denuclearisation of Central Europe, 1957–59. Cold War History, 21(4):509–531, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1873278.

Mongin:2023:RAU

[MV23] Dominique Mongin and Maurice Vaïsse. Response to the article ‘Unfit for purpose: reassessing the development and deployment of French nuclear weapons (1956–74)’ by Mongin and Vaïsse. Cold War History, 23(3):459–463, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2207316. See [PP21, PP23].

Mytton:2011:ARB

[Myt11] Graham Mytton. Audience research at the BBC External Services during the Cold War: a view from the inside. Cold War History, 11 (1):49–67, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.545597.

Marchesi:2022:NLA

[MZ22] Aldo Marchesi and Michal Zourek. The New Latin American Left in a polarised Cold War: The story of Vivian Trías. Cold War History, 22(1):19–40, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1923697.

Nadzhafov:2004:BCW

[Nad04] Dzhahangir G. Nadzhafov. The begining of the Cold War between East and West: The aggravation of ideological confrontation. Cold War History, 4(2):140–174, 2004. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740412331391845.

Nafpliotis:2013:APM

[Naf13] Alexandros Nafpliotis. Andreas Papandreou: The making of a Greek democrat and political maverick. Cold War History, 13(1): 141–142, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.762238.

Nateras:2018:SRL

[Nat18] Gerardo Sánchez Nateras. The Sandinista revolution and the limits of the Cold War in Latin America: the dilemma of non-intervention during the Nicaraguan crisis, 1977–78. Cold War History, 18(2): 111–129, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1369046.

Navarro:2011:CTM

[Nav11] Aaron W. Navarro. Che’s travels: the making of a revolutionary in 1950s Latin America. Cold War History, 11(3):484–485, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.599986.

Nayudu:2018:VES

[Nay18] Swapna Kona Nayudu. ‘In the very eye of the storm’: India, the UN, and the Lebanon crisis of 1958. Cold War History, 18(2): 221–237, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1445997.

Nehring:2004:WNP

[Neh04] Holger Nehring. ‘Westernization’: a new paradigm for interpreting West European history in a Cold War context. Cold War History, 4 (2):175–191, 2004. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740412331391855.

Nemchenok:2010:SSA

[Nem10] Victor V. Nemchenok. In search of stability amid chaos: US policy toward Iran, 1961–63. Cold War History, 10(3):341–369, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903178579.

Newton:2011:SIH

[New11] Julie Newton. Superpower illusions: how myths and false ideologies led America astray — and how to return to reality. Cold War History, 11(4):685–686, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.617468.

Ngeow:2023:DGT

[Nge23] Chow Bing Ngeow. Dragon in the Golden Triangle: military operations of the People’s Liberation Army in Northern Burma, 1960–1961. Cold War History, 23(1):61–82, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2100355.

Nicolet:2002:DUP

[Nic02] C. Nicolet. The development of US plans for the resolution of the Cyprus Conflict in 1964: ’the limits of American power’. Cold War History, 3 (1):95–126, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999969.

Niebuhr:2021:YPA

[Nie21] Robert Niebuhr. Yugoslavia and political assassinations: the history and legacy of Tito’s campaign against the émigrés. Cold War History, 21(2):238–240, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1886896.

Nuti:2015:I

[NO15] Leopoldo Nuti and Christian Ostermann. Introduction. Cold War History, 15(3):273–276, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2015.1074442.

Noer:2012:ADC

[Noe12] Tom Noer. America, the UN and decolonization: Cold War conflict in the Congo. Cold War History, 12(4):708–709, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2012.737565.

Nuti:2002:RRS

[Nut02] L. Nuti. Research report — sources for the study of Italian foreign policy, 1861–1999. Cold War History, 2(3):93–110, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/713999961.

Nuti:2014:WAC

[Nut14] Leopoldo Nuti. Western anti-communism and the Interdoc network: Cold War Internationale. Cold War History, 14(4):709–711, 2014. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2014.955694.

Nuti:2017:NNS

[Nut17] Leopoldo Nuti. Nixon’s nuclear spectre: The secret alert of 1969, madman diplomacy, and the Vietnam War. Cold War History, 17 (4):445–448, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1382976.

Nehring:2012:DAP

[NZ12] Holger Nehring and Benjamin Ziemann. Do all paths lead to Moscow? The NATO dual-track decision and the peace movement — a critique. Cold War History, 12(1):1–24, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2011.625160. See corrigendum [Ano14a,

?].

Ochman:2006:SLP

[Och06] Ewa Ochman. The search for legitimacy in post-martial law Poland: The case of Claude Lanzmann’s Shoah. Cold War History, 6(4): 501–526, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600979246.

Oldham:2020:DLS

[Old20] Joseph Oldham. ‘Don’t let the side down, old boy’: interrogating the traitor in the ‘radical’ television dramas of John le Carré and Dennis Potter. Cold War History, 20(3):311–327, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2019.1638366.

Olesen:2008:UNP

[Ole08] Thorsten B. Olesen. Under the national paradigm: Cold War studies and Cold War politics in post-Cold War Norden. Cold War History, 8 (2):189–211, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802018686.

OMalley:2012:WKH

[O’M12] Alanna O’Malley. Who killed Hammarskjöld? The UN, the Cold War and white supremacy in Africa. Cold War History, 12(3): 567–568, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.702077.

Onate:2011:RAF

[Oña11] Andrea Oñate. The Red Affair: FMLN–Cuban relations during the Salvadoran Civil War, 1981–92. Cold War History, 11(2):133– 154, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2010.545566.

Onslow:2005:QTS

[Ons05a] Sue Onslow. A question of timing: South Africa and Rhodesia’s unilateral declaration of independence, 1964–65. Cold War History, 5 (2):129–159, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500062135.

Onslow:2005:RSA

[Ons05b] Sue Onslow. Republic of South Africa archives. Cold War History, 5(3):369–375, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500222150.

Onslow:2007:D

[Ons07] Sue Onslow. Documents. Cold War History, 7(2):305–325, 2007. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740701284355.

Onslow:2013:FRU

[Ons13] Sue Onslow. Foreign relations of the United States, 1969–1976, vol. XXVIII: Southern Africa. Cold War History, 13(1):137–138, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2013.762239.

Onslow:2016:BRS

[Ons16] Sue Onslow. Book review: Stephen A. Emerson, The Battle for Mozambique: The Frelimo–Renamo Struggle, 1977–1992 (Pinetown, South Africa, 30^{c}irc South, 2013), 240 pp. Cold War History, 16 (2):235–237, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1153849.

Onslow:2022:RNS

[Ons22] Sue Onslow. Research notes special collection: the Cold War in Southern Africa. Cold War History, 22(3):343–358, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2022.2104434.

Onslow:2024:RTE

[Ons24] Sue Onslow. Round table: The end of the Cold War and the end of apartheid, April 1994. Cold War History, 24(2):309–315, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2024.2317463.

Osornprasop:2007:AHC

[Oso07] Sutayut Osornprasop. Amidst the heat of the Cold War in Asia: Thailand and the American secret war in Indochina (1960–74). Cold War History, 7(3):349–371, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740701474832.

PerezdeArcos:2024:BRT

[P´ 24] Marina Pérez de Arcos. Book review: A través del Telón de Acero: Historia de las relaciones políticas entre España y la RDA (1973– 1990). Xavier María Ramos Diez-Astrain, (Madrid: Centro de Estudios Políticos y Constitucionales, M^{o} de la Presidencia, 2021), 331 pp. Cold War History, 24(1):153–154, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2121359.

Paczkowski:2003:MVW

[Pac03] Andrzej Paczkowski. Models of visits by Western politicians: Poland and Western diplomats in 1987. Cold War History, 3(3):127– 143, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391668.

Pagliarini:2021:PNH

[Pag21] Andre Pagliarini. Peripheral nerve: health and medicine in Cold War Latin America. Cold War History, 21(3):386–389, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1934077.

Painter:2006:PHC

[Pai06] David Painter. A partial history of the Cold War. Cold War History, 6(4):527–534, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/

14682740600979295. See response [Gad07].

Painter:2009:MPO

[Pai09] David S. Painter. The Marshall Plan and oil. Cold War History, 9 (2):159–175, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740902871851.

Palmer:2010:VSD

[Pal10] Andrew Palmer. Vatican secret diplomacy: Joseph P. Hurley and Pope Pius XII. Cold War History, 10(1):140–141, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740903398854.

Palmer:2014:NWP

[Pal14] Diego A. Ruiz Palmer. The NATO–Warsaw Pact competition in the 1970s and 1980s: a revolution in military affairs in the making or the end of a strategic age? Cold War History, 14(4):533–573, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.950250.

Parr:2006:SCF

[Par06] Helen Parr. Saving the community: The French response to Britain’s second EEC application in 1967. Cold War History, 6(4): 425–454, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600979287.

Paraskevov:2011:CNB

[Par11] Vasil Paraskevov. Conflict and necessity: British–Bulgarian relations, 1944–56. Cold War History, 11(2):241–268, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2010.504206.

Parker:2012:CMM

[Par12] Chad H. Parker. Controlling man-made malaria: Corporate modernisation and the Arabian American Oil Company’s malaria control program in Saudi Arabia, 1947–1956. Cold War History, 12(3): 473–494, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.621751.

Patman:2007:SRA

[Pat07] Robert G. Patman. Some reflections on Archie Brown and the end of the Cold War. Cold War History, 7(3):439–445, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682740701474899. See

[Bro07].

Path:2011:EFS

[Pat11] Kosal Path. The economic factor in the Sino–Vietnamese split, 1972– 75: an analysis of Vietnamese archival sources. Cold War History, 11 (4):519–555, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
01446193.2010.512497.

Patti:2015:OBN

[Pat15] Carlo Patti. The origins of the Brazilian nuclear programme, 1951–1955. Cold War History, 15(3):353–373, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.968557.

Patalano:2021:SFS

[Pat21] Alessio Patalano. ‘The silent fight’: submarine rearmament and the origins of Japan’s military engagement with the Cold War, 1955–76. Cold War History, 21(1):91–111, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2019.1615894.

Peacock:2012:PBC

[Pea12] Margaret Peacock. The perils of building Cold War consensus at the 1957 Moscow World Festival of Youth and Students. Cold War History, 12(3):515–535, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.645809.

Pechatnov:2001:EFS

[Pec01] V. Pechatnov. Exercise in frustration: Soviet foreign propaganda in the Early Cold War, 1945–47. Cold War History, 1(2):1–27, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999921.

Pechatnov:2014:HSC

[Pec14] Vladimir Pechatnov. How Soviet Cold Warriors viewed World War II: the inside story of the 1957 edition of the Big Three correspondence. Cold War History, 14(1):109–125, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2013.796735.

Pechlivanis:2017:BDD

[Pec17] Paschalis Pechlivanis. Between Détente and differentiation: Nixon’s visit to Bucharest in August 1969. Cold War History, 17(3):241– 258, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1267144.

Pedaliu:2001:BRP

[Ped01] E. Pedaliu. Britain and the reconstruction of the post-Fascist Italian Armed Forces, 1943–48. Cold War History, 2(1):39–68, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999941.

Pedaliu:2013:BRS

[Ped13] Effie G. H. Pedaliu. Book review: Sarah B. Snyder, Human Rights Activism and the End of the Cold War: a Transnational History of the Helsinki Network (Cambridge: Cambridge University Press, 2011), x + 293 pp. Cold War History, 13(4):572–574, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.849019.

Pedersen:2016:PPF

[Ped16] Rasmus Brun Pedersen. Past, present, and future: the role of the Cold War in legitimising Danish foreign policy activism. Cold War History, 16(1):101–120, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1078791.

Pedemonte:2022:FGC

[Ped22] Rafael Pedemonte. The first generation of Cuban students in the 1960s Soviet Union: shaping a revolutionary ‘culture of militancy’. Cold War History, 22(4):479–497, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2057472.

Peoples:2008:SST

[Peo08] Columba Peoples. Sputnik and ‘skill thinking’ revisited: technological determinism in American responses to the Soviet missile threat. Cold War History, 8(1):55–75, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740701791334.

Pero:2011:WCA

[Per11] Mario Del Pero. ‘Which Chile, Allende?’ Henry Kissinger and the Portuguese revolution. Cold War History, 11(4):625–657, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.494301.

Pero:2012:NKU

[Per12] Mario Del Pero. Nixon, Kissinger, and U.S. foreign policy making: the machinery of crisis. Cold War History, 12(4):710–711, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2012.737562.

Pero:2016:WAS

[Per16] Mario Del Pero. Worldmaking: The art and science of American diplomacy. Cold War History, 16(3):331–333, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2016.1187324.

Perlman:2017:CFF

[Per17] Susan McCall Perlman. Contesting France: French informants and American intelligence in the dawning Cold War. Cold War History, 17(1):81–98, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1205587.

Person:2019:NKR

[Per19] James F. Person. North Korea in 1956: reconsidering the August Plenum and the Sino–Soviet joint intervention. Cold War History, 19 (2):253–274, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1509849.

Pettina:2011:SCW

[Pet11] Vanni Pettinà. The shadows of Cold War over Latin America: the US reaction to Fidel Castro’s nationalism, 1956–59. Cold War History, 11 (3):317–339, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003686115.

Petersen:2012:FPI

[Pet12a] Nikolaj Petersen. ‘Footnoting’ as a political instrument: Denmark’s NATO policy in the 1980s. Cold War History, 12(2):295–317, 2012. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2011.558176.

Pettina:2012:BP

[Pet12b] Vanni Pettinà. The Bay of Pigs. Cold War History, 12(1):172– 173, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.655445.

Peter:2014:EIA

[Pet14] Matthias Peter. Europe in the international arena during the 1970s: Entering a different world. Cold War History, 14(2):283–285, 2014. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2014.893700.

Pharo:2012:GCC

[Pha12] Helge Pharo. The global 1989: continuity and change in world politics. Cold War History, 12(2):374–375, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.686591.

Piirimae:2009:BQD

[Pii09] Kaarel Piirimäe. The Baltic question during the Cold War. Cold War History, 9(4):540–542, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903268511.

Pince:2024:FNS

[Pin24] Yannick Pincé. French nuclear strategy: The making of a ‘consensus’ (1972–1988). Cold War History, 24(2):239–256, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2023.2282703.

Pittaway:2003:EDR

[Pit03] Mark Pittaway. The education of dissent: The reception of the Voice of Free Hungary, 1951–56. Cold War History, 4(1):97–116, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740312331391754.

Pleshakov:2013:RCW

[Ple13] Constantine Pleshakov. Russia’s Cold War: From the October Revolution to the fall of the Wall. Cold War History, 13(1):132–133, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2013.762240.

Pons:2002:DMB

[Pon02] S. Pons. Documents — meetings between the Italian Communist Party and the Communist Party of the Soviet Union, Moscow and Rome, 1978– 80. Cold War History, 3(1):157–166, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/713999974.

Popp:2010:AWR

[Pop10] Roland Popp. Accommodating to a working relationship: Arab Nationalism and US Cold War policies in the Middle East, 1958–60. Cold War History, 10(3):397–427, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682741003686107.

Popp:2015:GRS

[Pop15] Roland Popp. Group review: From sideshow to centre stage: The Middle East in the Middle Cold War. Cold War History, 15(4): 579–586, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1086474.

Pelopidas:2021:UPR

[PP21] Benoît Pelopidas and Sébastien Philippe. Unfit for purpose: reassessing the development and deployment of French nuclear weapons (1956–1974). Cold War History, 21(3):243–260, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2020.1832472. See

responses [PP23, MV23].

Pelopidas:2023:RAU

[PP23] Benoît Pelopidas and Sébastien Philippe. Response to the article ‘Unfit for purpose: reassessing the development and deployment of French nuclear weapons (1956–74)’ by Pelopidas and Philippe. Cold War History, 23(3):453–457, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2207314. See [PP21].

Prazmowska:2002:KPE

[Pra02] A. Prazmowska. The Kielce Pogrom 1946 and the emergence of Communist power in Poland. Cold War History, 2(2):101–124, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999953.

Prazmowska:2010:RSP

[Pra10] Anita J. Prazmowska. Rebellious satellite: Poland 1956. Cold War History, 10(3):458–459, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2010.493714.

Prados:2011:VWR

[Pra11] John Prados. A Vietnam War reader: a documentary history from American and Vietnamese perspectives. Cold War History, 11(2): 283–284, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569158.

Prazmowska:2012:SSW

[Pra12] Anita J. Prazmowska. Solidarity with Solidarity: Western European trade unions and the Polish crisis, 1980–1982. Cold War History, 12 (4):713–715, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.737567.

Prasad:2024:WIT

[Pra24] Binay Prasad. What it takes to recognise a new government? India’s diplomatic recognition and understanding of Castro’s Cuba. Cold War History, 24(1):131–152, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2023.2240155.

Preston:2009:RAF

[Pre09a] Andrew Preston. Religion and American foreign policy, 1945–1960: the soul of containment. Cold War History, 9(4):534–536, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903268529.

Preston:2009:SKW

[Pre09b] Andrew Preston. Selling the Korean War: propaganda, politics, and public opinion in the United States, 1950–1953. Cold War History, 9 (4):530–532, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903268537.

Preston:2012:FWH

[Pre12] Andrew Preston. Faith and war: how Christians debated the Cold and Vietnam Wars. Cold War History, 12(3):560–562, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2012.702078.

Preston:2013:PVA

[Pre13] Andrew Preston. Peripheral visions: American mainline Protestants and the global Cold War. Cold War History, 13(1):109–130, 2013. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2012.707648.

Prentice:2022:FNV

[Pre22] David L. Prentice. Finding nguyen Van thieu: the value of multinational, multi-archival research. Cold War History, 22(4):529– 532, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2124895.

Priest:2022:BRM

[Pri22] Tyler Priest. Book review: Megan Black, The global interior: mineral frontiers and American power (Cambridge: Harvard University Press, 2019), 248 pp. Cold War History, 22(4):533–538, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.2016469. See

response [Bla22].

Prott:2023:WTW

[Pro23] Volker Prott. ‘We have to tread warily’: East Pakistan, India and the pitfalls of foreign intervention in the Cold War. Cold War History, 23(1):23–44, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2102607.

Pucci:2020:EFS

[Puc20] Molly Pucci. Empire of friends: Soviet power and socialist internationalism in Cold War Czechoslovakia. Cold War History, 20(1): 115–117, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1695796.

Puma:2022:GSE

[Pum22] Jorge Puma. The gathering storm: Eduardo Frei’s revolution in liberty and Chile’s Cold War. Cold War History, 22(2):239–241, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1983936.

Quinn-Judge:2005:IDD

[QJ05] Sophie Quinn-Judge. The ideological debate in the DRV and the significance of the anti-party affair, 1967–68. Cold War History, 5(4): 479–500, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500284838.

Quinn-Judge:2009:BRG

[QJ09] Sophie Quinn-Judge. Book review: Les grandes puissances et le Laos, 1954–1964 [The Great Powers and Laos, 1954–1964], by Laurent Césari, Arras, Artois Presses Université, 2007, 374 pp. Cold War History, 9 (4):543–544, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903268545.

Quinn-Judge:2011:CCF

[QJ11] Sophie Quinn-Judge. China, Cambodia, and the five principles of peaceful coexistence. Cold War History, 11(4):684–685, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2011.617445.

Rabe:2009:CWH

[Rab09] Stephen G. Rabe. Cold War in a hot zone: the United States confronts labor and independence struggles in the British West Indies. Cold War History, 9(4):533–534, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903268552.

Radchenko:2003:MA

[Rad03] Sergey S. Radchenko. Mongolian archives. Cold War History, 3(3): 107–110, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391648.

Raine:2001:SCA

[Rai01] F. Raine. Stalin and the creation of the Azerbaijan Democratic Party in Iran, 1945. Cold War History, 2(1):1–38, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/713999940.

Rakove:2014:TRB

[Rak14] Robert B. Rakove. Two roads to Belgrade: the United States, Great Britain, and the first nonaligned conference. Cold War History, 14 (3):337–357, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.871528.

Raber:2024:PTF

[RB24] Richard Levi Raber and Lennart Bolliger. Piecing together a fragmentary history: African soldiers from decolonization to the post-Cold War world. Cold War History, 24(2):339–346, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2024.2317464.

Readman:2006:BPR

[Rea06] Kristina Spohr Readman. Between political rhetoric and Realpolitik calculations: Western diplomacy and the Baltic independence struggle in the Cold War endgame. Cold War History, 6(1):1–42, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740500395402.

Readman:2010:CDE

[Rea10] Kristina Spohr Readman. The crisis of détente in Europe: from Helsinki to Gorbachev, 1975–1985. Cold War History, 10(3):460–461, 2010. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2010.493716.

Readman:2012:SGP

[Rea12] Kristina Spohr Readman. The shock of the global: the 1970s in perspective. Cold War History, 12(3):559–560, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2012.702080.

Reid:2010:CWU

[Rei10] Brian Holden Reid. The Cold War U.S. Army: building deterrence for limited war. Cold War History, 10(3):464–465, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2010.493715.

Reinsone:2016:FSF

[Rei16] Sanita Reinsone. Forbidden and sublime forest landscapes: narrated experiences of Latvian national partisan women after World War II. Cold War History, 16(4):395–416, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.986108.

Rentola:2000:HAH

[Ren00] K. Rentola. From half-adversary to half-ally: Finland in Soviet policy, 1953–58. Cold War History, 1(1):75–102, 2000. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999907.

Renouard:2018:SMH

[Ren18] Joe Renouard. From Selma to Moscow: how human rights activists transformed U.S. foreign policy. Cold War History, 18(4): 525–528, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1533690.

Rey:2004:EOC

[Rey04] Marie-Pierre Rey. ‘Europe is our common home’: a study of Gorbachev’s diplomatic concept. Cold War History, 4(2):33–65, 2004. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740412331391805.

Reynolds:2007:R

[Rey07] David Reynolds. Reviews. Cold War History, 7(1):147–162, 2007. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740701197805.

Rey:2013:HRP

[Rey13] Marie-Pierre Rey. Human rights, perestroika and the end of the Cold War. Cold War History, 13(1):140–141, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2012.737558.

Reynolds:2023:OIA

[Rey23] Bradley Reynolds. Not one inch: America, Russia and the making of post-Cold War stalemate. Cold War History, 23(1):209–212, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2022.2077313.

Rheaume:2011:CNP

[Rhé11] Charles Rhéaume. Cautious neighbour policy: Canada’s helping hand in winding down the Vietnam War. Cold War History, 11(2): 223–239, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903527684.

Rietkerk:2014:BRR

[Rie14] Aaron Rietkerk. Book review: Roger Lipsey, Hammarskjöld: a Life (Ann Arbor: University of Michigan Press, 2013), 768 pp. Cold War History, 14(2):282–283, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2014.893701.

Riehle:2019:ECW

[Rie19] Kevin P. Riehle. Early Cold War evolution of British and US defector policy and practice. Cold War History, 19(3):343–361, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2018.1539079.

Risso:2007:EPO

[Ris07] Linda Risso. ‘Enlightening public opinion’: a study of NATO’s information policies between 1949 and 1959 based on recently declassified documents. Cold War History, 7(1):45–74, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740701197672.

Risso:2009:DMS

[Ris09] Linda Risso. ‘Don’t mention the Soviets!’ An overview of the short films produced by the NATO Information Service between 1949 and 1969. Cold War History, 9(4):501–512, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903198239.

Risso:2011:PWN

[Ris11] Linda Risso. Propaganda on wheels: The NATO travelling exhibitions in the 1950s and 1960s. Cold War History, 11(1):9–25, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.545595.

Risso:2013:RWB

[Ris13] Linda Risso. Radio Wars: Broadcasting in the Cold War. Cold War History, 13(2):145–152, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2012.757134.

Risso:2019:EAH

[Ris19] Linda Risso. Enduring alliance: a history of NATO and the postwar global order. Cold War History, 19(4):588–589, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2019.1678817.

Rizas:2009:MCB

[Riz09] Sotiris Rizas. Managing a conflict between allies: United States policy towards Greece and Turkey in relation to the Aegean dispute, 1974–76. Cold War History, 9(3):367–387, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740802170883.

Rizas:2012:UEC

[Riz12] Sotiris Rizas. The US embassy and the collapse of democracy in Greece, 1966–1969: the testimony of a diplomat. Cold War History, 12(4): 709–710, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.737559.

Roberts:2016:BRR

[Rob16a] Geoffrey Roberts. Book review: Raymond L. Garthoff, Soviet Leaders and Intelligence: Assessing the American Adversary During the Cold War (Washington, DC: Georgetown University Press, 2015), xviii + 142 pp. Cold War History, 16(2):233–234, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2016.1153850.

Robertson:2016:CWL

[Rob16b] Thomas Robertson. Cold War landscapes: towards an environmental history of US development programmes in the 1950s and 1960s. Cold War History, 16(4):417–441, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950238.

Roberts:2017:AEM

[Rob17] George Roberts. The assassination of Eduardo Mondlane: FRELIMO, Tanzania, and the politics of exile in Dar es Salaam. Cold War History, 17(1):1–19, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1246542.

Roehrlich:2016:CWD

[Roe16] Elisabeth Roehrlich. The Cold War, the developing world, and the creation of the International Atomic Energy Agency (IAEA), 1953–1957. Cold War History, 16(2):195–212, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1129607.

Rognoni:2012:ECW

[Rog12] Maria Stella Rognoni. The end of the Cold War in the Third World: new perspectives on regional conflict. Cold War History, 12 (3):568–569, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.702082.

Romero:2011:SMU

[Rom11] Federico Romero. For the soul of mankind: the United States, the Soviet Union, and the Cold War. Cold War History, 11(4): 677–678, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.617429.

Romano:2014:UCW

[Rom14a] Angela Romano. Untying Cold War knots: The EEC and Eastern Europe in the long 1970s. Cold War History, 14(2):153–173, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.791680.

Romero:2014:CWH

[Rom14b] Federico Romero. Cold War historiography at the crossroads. Cold War History, 14(4):685–703, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950249.

Roman:2018:SRB

[Rom18a] Meredith L. Roman. Soviet ‘renegades’, Black Panthers, and Angela Davis: the politics of dissent in the Soviet press, 1968–73. Cold War History, 18(4):503–519, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1465932.

Romero:2018:CWW

[Rom18b] Federico Romero. The Cold War: a world history. Cold War History, 18(2):240–243, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1449568.

Roth:2023:BAR

[Rot23] Sarah Roth. Burn after reading: Operation Focus and the fictional Nemzeti Ellenzéki Mozgalom [National Opposition Movement] in the lead-up to the 1956 Hungarian uprising. Cold War History, 23(2): 239–262, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2124972.

Roy:2002:BBW

[Roy02] R. Roy. The battle for Bretton Woods: America, Britain and the international financial crisis of October 1967–March 1968. Cold War History, 2(2):33–60, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999955.

Radchenko:2017:RWK

[RS17] Sergey Radchenko and Bernd Schaefer. ‘Red on white’: Kim Il Sung, Park Chung Hee, and the failure of Korea’s reunification, 1971–1973. Cold War History, 17(3):259–277, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1265508.

Rubinson:2014:GEN

[Rub14] Paul Rubinson. The global effects of nuclear winter: science and antinuclear protest in the United States and the Soviet Union during the 1980s. Cold War History, 14(1):47–69, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.759560.

Ruotsila:2005:BR

[Ruo05] Markku Ruotsila. Book reviews. Cold War History, 5(2):259– 279, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500062143.

Rutenberg:2013:DDA

[Rut13] Amy J. Rutenberg. Drafting for domesticity: American deferment policy during the Cold War, 1948–1965. Cold War History, 13(1): 1–20, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.728208.

Saito:2021:CFW

[Sai21] Yoshiomi Saito. ‘The country is full of wishful thinkers’: Britain’s Information Research Department and its post-war propaganda operations in Japan, 1948–70. Cold War History, 21(4):449–468, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2020.1832992.

Salmon:2010:MEC

[Sal10] Patrick Salmon. Mitterrand, the end of the Cold War, and German unification. Cold War History, 10(4):579–581, 2010. CODEN ????

ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.513514.

Salgado:2018:CHR

[Sal18] Alfonso Salgado. Communism and human rights in Pinochet’s Chile: the 1977 hunger strike against forced disappearance. Cold War History, 18(2):169–186, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1404988.

Samuel:2012:CIR

[Sam12] Robert Samuel. Conservative intellectuals and the Reagan–Gorbachev summits. Cold War History, 12(1):135–157, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.524209.

Sambaluk:2014:UPC

[Sam14] Nicholas Michael Sambaluk. US policymakers confront aerospace doctrine, 1957–59. Cold War History, 14(1):91–107, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.782536.

Sarantakes:2009:MVA

[Sar09a] Nicholas Evan Sarantakes. Moscow versus Los Angeles: the Nixon White House wages Cold War in the Olympic selection process. Cold War History, 9(1):135–157, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740802573508.

Sargent:2009:NWA

[Sar09b] Daniel Sargent. Nixon in the world: American foreign relations, 1969– 1977. Cold War History, 9(4):532–533, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740903268578.

Sarotte:2010:HTE

[Sar10] Mary E. Sarotte. Helsinki 1975 and the transformation of Europe. Cold War History, 10(2):275–276, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682741003711871.

Sargent:2013:CWI

[Sar13] Daniel Sargent. The Cold War and the international political economy in the 1970s. Cold War History, 13(3):393–425, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2013.789693.

Sarkar:2015:WTF

[Sar15] Jayita Sarkar. ‘Wean them away from French tutelage’: Franco–Indian nuclear relations and Anglo–American anxieties during the early Cold War, 1948–1952. Cold War History, 15(3):375–394, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2014.989840.

Sarkar:2021:DDP

[Sar21] Jayita Sarkar. From the dependable to the demanding partner: the renegotiation of French nuclear cooperation with India, 1974–80. Cold War History, 21(3):301–318, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2019.1694908.

Saull:2013:BRC

[Sau13] Richard Saull. Book review: Constructing a Post-War Order: The Rise of US Hegemony and the Origins of the Cold War (London: I. B. Tauris, 2011), 327 pp. Cold War History, 13(4):566–567, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.849022.

Saunders:2017:JCA

[Sau17] Chris Saunders. Jimmy Carter in Africa: Race and the Cold War. Cold War History, 17(1):102–104, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1226820.

Saunders:2024:TES

[Sau24] Chris Saunders. Two endings: semi-historiographical musings. Cold War History, 24(2):317–322, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2024.2317460.

Sayle:2009:AHC

[Say09] Timothy Andrews Sayle. Andropov’s Hungarian complex. Cold War History, 9(3):427–439, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740902764528.

Sayle:2016:GLP

[Say16] Timothy Andrews Sayle. ‘A great list of potential mistakes’: NATO, Africa, and British efforts to limit the global Cold War. Cold War History, 16(1):19–36, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2015.1078313.

Sbardellati:2009:MAR

[Sba09] John Sbardellati. ‘The Maltz Affair’ revisited: How the American Communist Party relinquished its cultural influence at the dawn of the Cold War. Cold War History, 9(4):489–500, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740903198221.

Scanlon:2011:VSF

[Sca11] Sandra Scanlon. Vietnam’s second front: domestic politics, the Republican Party, and the war. Cold War History, 11(2):285–286, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2011.569159.

Scala:2012:DDG

[Sca12] Stephen J. Scala. Divided, but not disconnected: German experiences of the Cold War. Cold War History, 12(1):176–178, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2012.655449.

Scanlon:2023:TCC

[Sca23] Sean Scanlon. A twentieth-century crusade: the vatican’s battle to remake Christian Europe. Cold War History, 23(4):551–553, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2022.2116779.

Schulze:2001:BR

[Sch01] K. Schulze. Book reviews. Cold War History, 2(1):164–175, 2001. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999945.

Schulze:2002:BRa

[Sch02a] K. Schulze. Book reviews. Cold War History, 2(3):129–135, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999962.

Schulze:2002:BRb

[Sch02b] K. Schulze. Book reviews. Cold War History, 3(1):167–172, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999971.

Schulze:2003:BR

[Sch03] K. Schulze. Book reviews. Cold War History, 3(2):166–173, 2003. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/713999983.

Schoenborn:2008:BBC

[Sch08a] Benedikt Schoenborn. Bargaining with the Bear: Chancellor Erhard’s bid to buy German reunification, 1963–64. Cold War History, 8(1):23–53, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701791318.

Schwartz:2008:LDT

[Sch08b] Thomas A. Schwartz. Legacies of détente: a three-way discussion. Cold War History, 8(4):513–525, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740802373594.

Schwartz:2011:PIC

[Sch11] Thomas A. Schwartz. Perforating the Iron Curtain: European détente, transatlantic relations, and the Cold War, 1965–1985. Cold War History, 11(4):678–679, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.617431.

Schmidli:2012:HRC

[Sch12] William Michael Schmidli. Human rights and the Cold War: the campaign to halt the Argentine ‘dirty war’. Cold War History, 12 (2):345–365, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569540.

Schoenborn:2014:CES

[Sch14] Benedikt Schoenborn. Chancellor Erhard’s silent rejection of de Gaulle’s plans: the example of monetary union. Cold War History, 14(3): 377–402, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.890592.

Schattenberg:2020:FAH

[Sch20a] Susanne Schattenberg. The final act: the Helsinki Accords and the transformation of the Cold War. Cold War History, 20(2): 248–250, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/

14682745.2020.1743032. Series: America in the world.

Schmidli:2020:BRP

[Sch20b] William Michael Schmidli. Book review: Petra Goedde, The politics of peace: a global Cold War History (New York: Oxford University Press, 2019), 320 pp. Cold War History, 20(4):521–523, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2020.1819304.

Schaefer:2022:BRE

[Sch22] Timo Schaefer. Book review: Eric Zolov, The last good neighbor: Mexico in the global sixties (Durham and London: Duke University Press, 2020), 424 pp. Cold War History, 22(1):127–130, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2021.1955703.

Schalbroeck:2023:PMH

[Sch23] Eva Schalbroeck. Protestant missionaries and humanitarianism in the DRC: the politics of aid in Cold War Africa. Cold War History, 23 (2):333–338, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2023.2190427.

Scott:2009:SVC

[Sco09] Carl-Gustaf Scott. Swedish Vietnam criticism reconsidered: Social Democratic Vietnam policy a manifestation of Swedish Ostpolitik? Cold War History, 9(2):243–266, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740802490273.

Scott:2010:TCD

[Sco10] Andrew Scott. Twentieth-century diplomacy: a case study of British practice, 1963–1976. Cold War History, 10(2):282–283, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682741003711897.

Seabra:2021:DSB

[Sea21] Pedro Seabra. ‘Despite the special bonds that tie us’: Portugal, Brazil, and the South Atlantic in the late Cold War. Cold War History, 21 (3):357–374, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2020.1832471.

See:2002:UTJ

[See02] J. See. An uneasy truce: John F. Kennedy and Soviet–American Détente, 1963. Cold War History, 2(2):161–194, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999948.

Seed:2003:DNR

[See03] David Seed. The debate over nuclear refuge. Cold War History, 4 (1):117–142, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391764.

Settje:2013:BRP

[Set13a] David E. Settje. Book review: Philip E. Muehlenbeck (ed.), Religion and the Cold War: a Global Perspective (Nashville: Vanderbilt University Press, 2012), xxii + 314 pp. Cold War History, 13 (2):285–286, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.789667.

Settje:2013:SIC

[Set13b] David E. Settje. The spiritual–industrial complex: America’s religious battle against Communism in the Early Cold War. Cold War History, 13(1):134–135, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.762241.

Sfikas:2014:AUI

[Sfi14] Thanasis D. Sfikas. ‘An almost unique isle in the sea of democratic Europe’: Greek Communists’ perceptions of international reality, 1944– 1949. Cold War History, 14(1):1–21, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2012.695349.

Share:2003:IFU

[Sha03a] M. Share. From ideological foe to uncertain friend: Soviet relations with Taiwan, 1943–82. Cold War History, 3(2):1–34, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/713999981.

Shaw:2003:SWM

[Sha03b] Tony Shaw. ‘Some writers are more equal than others’: George Orwell, the state and cold war privilege. Cold War History, 4(1): 143–170, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391774.

Shaw:2010:CWU

[Sha10] Tony Shaw. The Cold War and the United States Information Agency: American propaganda and public diplomacy, 1945–1989. Cold War History, 10(2):276–277, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682741003711905.

Shaw:2017:SRB

[Sha17] Alexander Nicholas Shaw. Sir Reader Bullard, Frank Roberts and the Azerbaijan crisis of 1945–46: Bevin’s officials, perceptions and the adoption of a Cold War mentality in British Soviet Policy. Cold War History, 17(3):279–297, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2016.1264389.

Shiwen:2009:SSS

[Shi09] Wang Shiwen. The Sino–Soviet split: Cold War in the communist world. Cold War History, 9(4):525–526, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740903268586.

Shillam:2020:SLG

[Shi20] Thomas William Shillam. Shattering the ‘looking-glass world’: the Congress for Cultural Freedom in South Asia, 1951–55. Cold War History, 20(4):441–459, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2019.1699067.

Shorten:2018:CWC

[Sho18] Richard Shorten. The Cold War as comparative political thought. Cold War History, 18(4):385–408, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1434508.

Shubin:2007:UHS

[Shu07] Vladimir Shubin. Unsung heroes: The Soviet military and the liberation of Southern Africa. Cold War History, 7(2):251–262, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740701284157.

Shubin:2012:DFC

[Shu12] Vladimir Shubin. A distant front in the Cold War: the USSR in West Africa and the Congo, 1956–1964. Cold War History, 12(3): 565–567, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.702079.

Shubin:2015:VFH

[Shu15] Vladimir Shubin. Visions of freedom: Havana, Washington, Pretoria and the struggle for Southern Africa, 1976–1991. Cold War History, 15 (3):421–424, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1051364.

Shubin:2024:SAE

[Shu24] Vladimir Shubin. South Africa and end of the Cold War. Cold War History, 24(2):335–338, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2024.2317467.

Simpson:2005:IBU

[Sim05] Brad Simpson. ‘Illegally and beautifully’: The United States, the Indonesian invasion of East Timor and the international community, 1974– 76. Cold War History, 5(3):281–315, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740500222028.

Simpson:2024:CCW

[Sim24] Thula Simpson. The consequences of the Cold War for the ANC. Cold War History, 24(2):347–350, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2024.2317466.

Sittmann:2018:ICI

[Sit18] Julia Sittmann. Illusions of care: Iraqi students between the Ba’thist State and the Stasi in socialist East Germany, 1958–89. Cold War History, 18(2):187–202, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2017.1415327.

Shen:2015:UOC

[SL15] Zhihua Shen and Julia Lovell. Undesired outcomes: China’s approach to border disputes during the Early Cold War. Cold War History, 15 (1):89–111, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.932350.

Smedley:2023:MTB

[Sme23] Stuart Smedley. Margaret Thatcher, British public opinion and German reunification, 1989–90. Cold War History, 23(3):431–451, 2023. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2023.2191950.

Smith:2001:HSI

[Smi01] M. L. Smith. Hear the silence: Investigating exclusion in Cold War international relations. Cold War History, 1(3):43–72, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999934.

Smith:2012:PCA

[Smi12] Mark B. Smith. Peaceful coexistence at all costs: Cold War exchanges between Britain and the Soviet Union in 1956. Cold War History, 12 (3):537–558, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.583239.

Smula:2006:PPS

[Smu06] Johann Smula. The party and the proletariat: Skoda 1948–53. Cold War History, 6(2):153–175, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740600650201.

Snyder:2010:FVR

[Sny10] Sarah Snyder. The foundation for Vienna: a reassessment of the CSCE in the mid-1980s. Cold War History, 10(4):493–512, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740903460357.

Snyder:2012:OWH

[Sny12] Sarah B. Snyder. An outsider in the White House: Jimmy Carter, his advisors, and the making of American foreign policy. Cold War History, 12(1):174–175, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.655446.

Snyder:2013:BCF

[Sny13a] Sarah B. Snyder. Beyond containment? The first Bush administration’s sceptical approach to the CSCE. Cold War History, 13(4): 463–484, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.758104.

Snyder:2013:BRF

[Sny13b] Sarah B. Snyder. Book review: Frédéric Bozo, Marie-Pierre Rey, N. Piers Ludlow, and Bernd Rother (eds.), Visions of the End of the Cold War in Europe, 1945–1990 (New York and Oxford: Berghahn Books, 2012), vii + 358 pp. Cold War History, 13(4):570–571, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.849024.

Snyder:2013:BSI

[Sny13c] Sarah B. Snyder. Brent Scowcroft: Internationalism and post-Vietnam War American foreign policy. Cold War History, 13(1): 139–140, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.762244.

Snyder:2018:CWW

[Sny18] Sarah B. Snyder. The Cold War: a world history. Cold War History, 18(2):243–245, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1449569.

Snyder:2021:EOS

[Sny21] Sarah B. Snyder. ‘Ending our support for the dictators’: Ed Koch, Uruguay, and human rights. Cold War History, 21(1):19–36, 2021. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2019.1705790.

Soares:2009:AHC

[Soa09] John Soares. Averell Harriman has changed his mind: the Seattle speech and the rhetoric of Cold War confrontation. Cold War History, 9(2):267–286, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802490430.

Somerville:2013:BRK

[Som13] Keith Somerville. Book review: Kristin Roth-Ely, Moscow Prime Time: How the Soviet Union Built the Media Empire That Lost the Cultural Cold War (Ithaca and London: Cornell University Press, 2011), ix + 315 pp. Cold War History, 13(4):571–572, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2013.849026.

Spektor:2010:EGA

[Spe10] Matias Spektor. Economists with guns: authoritarian development and US–Indonesian relations, 1960–1968. Cold War History, 10(1): 136–137, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903398912.

Spelling:2013:RRA

[Spe13] Alex Spelling. ‘Recrimination and reconciliation’: Anglo–American relations and the Yom Kippur War. Cold War History, 13(4):485– 506, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.679658.

Spinardi:2006:STC

[Spi06] Graham Spinardi. Science, technology, and the Cold War: The military uses of the Jodrell Bank Radio Telescope. Cold War History, 6 (3):279–300, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600795428.

Spohr:2015:GAS

[Spo15] Kristina Spohr. Germany, America and the shaping of post-Cold War Europe: a story of German international emancipation through political unification, 1989–90. Cold War History, 15(2):221–243, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2015.1018479.

Scott-Smith:2007:CPC

[SS07] Giles Scott-Smith. Confronting peaceful co-existence: Psychological warfare and the role of Interdoc, 1963–72. Cold War History, 7 (1):19–43, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701197664.

Scott-Smith:2009:AHP

[SS09] Giles Scott-Smith. American hegemony and the postwar reconstruction of science in Europe. Cold War History, 9(3):443–444, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903135421.

Stankova:2012:STE

[Sta12] Marietta Stankova. Stalinist terror in Eastern Europe: elite purges and mass repression. Cold War History, 12(2):376–377, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2012.686592.

Stanciu:2013:CMC

[Sta13] Cezar Stanciu. Crisis management in the Communist bloc: Romania’s policy towards the USSR in the aftermath of the Prague Spring. Cold War History, 13(3):353–372, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.746955.

Stanciu:2019:LCB

[Sta19] Cezar Stanciu. A lost chance for Balkan cooperation? The Romanian view on ‘regional micro-détente’, 1969–75. Cold War History, 19 (3):421–439, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1524878.

Stadler:2023:TSW

[Sta23] Liliane Stadler. The twilight struggle: what the Cold War teaches us about Great Power rivalry today. Cold War History, 23(1): 212–215, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2077988.

Stadler:2024:ACS

[Sta24] Liliane Stadler. Afghan crucible: the Soviet invasion and the making of modern Afghanistan. Cold War History, 24(1):157–160, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2158711.

Stenner:2016:BTE

[Ste16] David Stenner. ‘Bitterness towards Egypt’ — the Moroccan nationalist movement, revolutionary Cairo and the limits of anti-colonial solidarity. Cold War History, 16(2):159–175, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1100605.

Stewart:2018:VAW

[Ste18] Geoffrey C. Stewart. Vietnam’s American war: a history. Cold War History, 18(4):523–525, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2018.1533689.

Stokes:2003:CST

[Sto03] Doug Stokes. Countering the Soviet threat? An analysis of the justifications for US military assistance to El salvador, 1979–92. Cold War History, 3(3):79–102, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682740312331391628.

Stoddart:2010:BLG

[Sto10] Kristan Stoddart. The British labour government and the development of Chevaline, 1974–79. Cold War History, 10(3):287–314, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682741003679375.

Stocker:2012:DCR

[Sto12] James Stocker. Diplomacy as counter-revolution? The ‘moderate states’, the Fedayeen and State Department initiatives towards the Arab–Israeli conflict, 1969–1970. Cold War History, 12(3):407– 428, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.621666.

Stocker:2014:OSC

[Sto14] James Stocker. Origins of the Suez Crisis: Postwar development diplomacy and the struggle over Third World industrialization (1945–1956). Cold War History, 14(2):281–282, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.893702.

Strupp:2023:RAI

[Str23] Christoph Strupp. Reconnecting across the Iron Curtain: Hamburg’s policy of the Elbe. Cold War History, 23(2):263–282, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2022.2106220.

Stykalin:2001:HCS

[Sty01] A. Stykalin. The Hungarian Crisis of 1956: The Soviet role in the light of new archival documents. Cold War History, 2(1):113–144, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999938.

Stykalin:2005:SYR

[Sty05] A. S. Stykalin. Soviet–Yugoslav relations and the case of Imre Nagy. Cold War History, 5(1):3–22, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/1468274042000283126.

Sun:2020:NCW

[Sun20] Jodie Yuzhou Sun. ‘Now the cry was Communism’: the Cold War and Kenya’s relations with China, 1964–70. Cold War History, 20 (1):39–58, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1602120.

Sun:2022:SCA

[Sun22] Jodie Yuzhou Sun. Supplied cash and arms but losing anyway: Chinese support of the Lumumbist insurgencies in the Congo Crisis (1959–65). Cold War History, 22(4):459–478, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2050699.

Suri:2006:CWD

[Sur06] Jeremi Suri. The Cold War, decolonization, and global social awakenings: Historical intersections. Cold War History, 6(3):353– 363, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600795519.

Suri:2008:DHR

[Sur08] Jeremi Suri. Détente and human rights: American and West European perspectives on international change. Cold War History, 8(4): 527–545, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802373628.

Shen:2010:WCZ

[SX10] Zhihua Shen and Yafeng Xia. The whirlwind of China: Zhou Enlai’s shuttle diplomacy in 1957 and its effects. Cold War History, 10 (4):513–535, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903167978.

Shen:2014:LTA

[SX14] Zhihua Shen and Yafeng Xia. Leadership transfer in the Asian revolution: Mao Zedong and the Asian Cominform. Cold War History, 14 (2):195–213, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.803471.

Sylvest:2022:PEN

[Syl22] Casper Sylvest. Pre-enacting the next war: the visual culture of Danish civil defence in the early nuclear age. Cold War History, 22 (1):79–102, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1874932.

Szalontai:2005:PEC

[Sza05] Balazs Szalontai. Political and economic crisis in North Vietnam, 1955– 56. Cold War History, 5(4):395–426, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740500284630.

Szalontai:2017:SWL

[Sza17] Balázs Szalontai. Solidarity within limits: Interkit and the evolution of the Soviet Bloc’s Indochina policy, 1967–1985. Cold War History, 17 (4):385–403, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1319818.

Szalontai:2023:MKN

[Sza23] Balázs Szalontai. Magyar-kínai kapcsolatok 1949–1989. Cold War History, 23(1):207–209, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2022.2077307.

Tagirova:2017:SPD

[Tag17] Alsu Tagirova. Soviet public diplomacy in China: “Small steps” towards bilateral rapprochement (1978–1985). Cold War History, 17(4): 405–423, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1307180.

Tagirova:2020:BRE

[Tag20] Alsu Tagirova. Book review: Edited by Janick Marina Schaufelbuehl, Marco Wyss, and Valeria Zanier. Europe and China in the Cold War: exchanges beyond the bloc logic and the Sino–Soviet split, (Leiden: Brill, 2019), xii + 242 pp. Cold War History, 20(4):527–529, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2020.1819315.

Tambor:2010:RSG

[Tam10] Molly Tambor. Red Saints: Gendering the Cold War, Italy 1943–1953. Cold War History, 10(3):429–456, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2010.494299.

Tang:2014:CWR

[Tan14] James T. H. Tang. Has the Cold War returned to East Asia? Cold War History, 14(4):647–656, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950242.

Tardelli:2012:LWH

[Tar12] Luca Tardelli. Leaders at war: how presidents shape military interventions. Cold War History, 12(1):171–172, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2012.655444.

Taubman:2009:GGS

[Tau09] William Taubman. Gorbachev’s gamble: Soviet foreign policy and the end of the Cold War. Cold War History, 9(4):526–527, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682740903268594.

Thompson:2005:BR

[Tho05] John A. Thompson. Book reviews. Cold War History, 5(1):121– 128, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
135145042000283111.

Thunemann:2021:SPR

[Thu21] Fabian Thunemann. Stalin: passage to revolution. Cold War History, 21(4):538–540, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2021.1941887.

Traverso:2023:WSC

[TL23] Diego Barría Traverso and Eduardo Carreño Lara. Welcome to Santiago. commercial aviation relations between Chile and the socialist countries, 1970–1973. Cold War History, 23(4):473–494, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2023.2206649.

Tobin:2002:RUV

[Tob02] K. Tobin. The reduction of urban vulnerability: Revisiting 1950s American suburbanization as civil defence. Cold War History, 2(2):1–32, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/713999949.

Tobia:2011:AAV

[Tob11a] Simona Tobia. Advertising America: VOA and Italy. Cold War History, 11(1):27–47, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2011.545596.

Tobia:2011:IEA

[Tob11b] Simona Tobia. Introduction: Europe Americanized? Popular reception of Western Cold War propaganda in Europe. Cold War History, 11(1):1–7, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.545593.

Tobia:2013:DRB

[Tob13] Simona Tobia. Did the RAI buy it? The role and limits of American broadcasting in Italy in the Cold War. Cold War History, 13 (2):171–191, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.746665.

Tobin:2016:PPP

[Tob16] Kathleen A. Tobin. People, not property: population issues and the neutron bomb. Cold War History, 16(3):307–325, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2016.1191470.

Tokic:2022:CWE

[Tok22] Mate Nikola Tokić. Cold War exiles and the CIA: plotting to free Russia. Cold War History, 22(3):375–377, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2021.2016053.

Tommel:2019:GQJ

[Töm19] Till Florian Tömmel. The German question in Jakarta Indonesia in West Germany’s foreign policy, 1955–65. Cold War History, 19 (1):119–140, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1504777.

Toth:2020:RNM

[Tót20] György Tóth. ‘Red’ nations: Marxists and the Native American sovereignty movement of the late Cold War. Cold War History, 20 (2):197–221, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1645126.

Trauschweizer:2006:TCC

[Tra06] Ingo Wolfgang Trauschweizer. Tanks at Checkpoint Charlie: Lucius Clay and the Berlin Crisis, 1961–62. Cold War History, 6(2): 205–228, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600650227.

Trauschweizer:2021:BCM

[Tra21] Ingo Trauschweizer. Berlin commander: Maxwell Taylor at the Cold War’s frontlines, 1949–51. Cold War History, 21(1):37–53, 2021. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/14682745.2019.1581766.

Trentin:2008:TNT

[Tre08] Massimiliano Trentin. ‘Tough negotiations’. the two Germanys in Syria and Iraq, 1963–74. Cold War History, 8(3):353–380, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740802222155.

Trecker:2020:ESS

[Tre20] Max Trecker. European Summer School 2016 Best Paper Prize Winner: Circle of debt: how the crisis of the Global South in the 1980s affected the socialist East. Cold War History, 20(1):1–19, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2019.1621846.

Tromly:2022:WCC

[Tro22] Benjamin Tromly. ‘To win the confidence of these curiously twisted and disoriented people’: the Ford Foundation’s Free Russia Fund, George F. Kennan and refugees from the Soviet Union. Cold War History, 22 (3):325–341, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2036720.

Tsvetkova:2024:PSC

[Tsv24] Natalia Tsvetkova. Professors and students in the cultural Cold War: The case of Ethiopia. Cold War History, 24(1):87–107, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2023.2231871.

Tubilewicz:2005:TSU

[Tub05] Czeslaw Tubilewicz. Taiwan and the Soviet Union during the Cold War: Enemies or ambiguous friends? Cold War History, 5(1): 75–86, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
1468274042000339179.

Tulli:2012:WRH

[Tul12] Umberto Tulli. ‘Whose rights are human rights?’ The ambiguous emergence of human rights and the demise of Kissingerism. Cold War History, 12(4):573–593, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.654491.

Tunc:2015:EST

[Tun15] Tanfer Emin Tunc. Eating in survival town: Food in 1950s atomic America. Cold War History, 15(2):179–200, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2014.950239.

Twigge:2016:AMP

[Twi16] Stephen Twigge. The Atomic Marshall Plan: Atoms for Peace, British diplomacy and civil nuclear power. Cold War History, 16(2): 213–230, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1117450.

Tyszkiewicz:2023:HRJ

[Tys23] Jakub Tyszkiewicz. Human rights and the Jimmy Carter administration’s policy towards Poland, 1977–80. Cold War History, 23(2):

307–325, 2023. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2022.2102606.

Unkovski-Korica:2014:YCS

[UK14] Vladimir Unkovski-Korica. The Yugoslav Communists’ Special Relationship with the British Labour Party 1950–1956. Cold War History, 14(1):23–46, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.765864.

Ulianova:2014:CBR

[Uli14] Olga Ulianova. Corvalán for Bukovsky: a real exchange of prisoners during an imaginary war. the Chilean dictatorship, the Soviet Union, and US mediation, 1973–1976. Cold War History, 14(3):315–336, 2014. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2013.793310.

Ullrich:2011:PPB

[Ull11] Weston Ullrich. Preventing ‘peace’: The British Government and the Second World Peace Congress. Cold War History, 11(3):341– 362, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682741003686123.

Ulunian:2003:SCW

[Ulu03] A. A. Ulunian. Soviet Cold War perceptions of Turkey and Greece, 1945–58. Cold War History, 3(2):35–52, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/713999982.

Urbansky:2018:KRN

[UT18] Sören Urbansky and Max Trecker. Kremlinology revisited: the nuances of reporting on China in the Eastern bloc press. Cold War History, 18 (3):307–324, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2018.1458093.

vanAlstein:2009:MHB

[vA09] Maarten van Alstein. The meaning of hostile bipolarization: Interpreting the origins of the Cold War. Cold War History, 9(3):301–319, 2009. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740902981395.

Vaisse:2012:GGI

[Vaï12] Maurice Vaïsse. Globalizing de Gaulle: international perspectives on French foreign policies, 1958–1969. Cold War History, 12(1): 180–182, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.655456.

Vamos:2018:HMC

[Vám18] Péter Vámos. A Hungarian model for China? Sino–Hungarian relations in the era of economic reforms, 1979–89. Cold War History, 18 (3):361–378, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1402887.

Vanke:2001:DGA

[Van01a] J. Vanke. DOCUMENT — De Gaulle’s atomic defence policy in 1963. Cold War History, 1(2):119–126, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/713999918.

Vanke:2001:IUD

[Van01b] J. Vanke. An impossible union: Dutch objections to the Fouchet Plan, 1959–62. Cold War History, 2(1):95–112, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999943.

Varsori:2003:BUI

[Var03] A. Varsori. Britain and US involvement in the Vietnam War during the Kennedy Administration, 1961–63. Cold War History, 3(2):83–112, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/713999980.

Varsori:2008:CWH

[Var08] Antonio Varsori. Cold War history in Italy. Cold War History, 8 (2):157–187, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802018678.

Varsori:2012:IEC

[Var12] Antonio Varsori. Italy and the end of communism in Albania, 1989– 1991. Cold War History, 12(4):615–635, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2011.650633.

vanDijk:2012:MPN

[vD12] Ruud van Dijk. ‘‘A mass psychosis”: The Netherlands and NATO’s dualtrack decision, 1978–1979. Cold War History, 12(3):381–405, 2012. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682745.2011.621750.

vanderHoog:2022:RNN

[vdH22] Tycho van der Hoog. Research note: the Nordic Africa Institute (NAI) Pamphlet Collection. Cold War History, 22(3):363–368, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2076345.

Velazquez:2019:MCF

[Vel19] Roberto Velázquez. Modernism came flying: a micro-history of artistic internationalism and cultural encounters in US–Chilean relations, 1968. Cold War History, 19(2):209–232, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1520212.

Vita:2017:ORT

[Vit17] Lorena De Vita. Overlapping rivalries: the two Germanys, Israel and the Cold War. Cold War History, 17(4):351–366, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2017.1322580.

Voss:2016:PDM

[Voß16] Klaas Voß. Plausibly deniable: mercenaries in US covert interventions during the Cold War, 1964–1987. Cold War History, 16(1): 37–60, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1078312.

vanWyk:2007:ACU

[vW07] Martha S. van Wyk. Ally or critic? The United States’ response to South African nuclear development, 1949–1980. Cold War History, 7 (2):195–225, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701284124.

vanWyk:2010:SAA

[vW10] Martha van Wyk. Sunset over atomic apartheid: United States– South African nuclear relations, 1981–93. Cold War History, 10(1): 51–79, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740902764569.

vanWyk:2011:NAQ

[vW11] Anna-Mart van Wyk. Nuclear apartheid: the quest for American supremacy from World War II to the present. Cold War History, 11 (2):273–274, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569149.

vanWyk:2015:AAA

[vW15] Jo-Ansie van Wyk. Atoms, apartheid, and the agency: South Africa’s relations with the IAEA, 1957–1995. Cold War History, 15 (3):395–416, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.897697.

vanWyk:2024:ECW

[vW24] Anna-Mart van Wyk. The end of the cold war in Southern Africa: the fall of the nuclear wall. Cold War History, 24(2):329–333, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2024.2315654.

Warlouzet:2013:EMM

[War13] Laurent Warlouzet. A Europe made of money: The emergence of the European monetary system. Cold War History, 13(1):142– 144, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.762245.

Waters:2010:SFI

[WD10] Robert Anthony Waters Jr. and Gordon Oliver Daniels. Striking for freedom? International intervention and the Guianese sugar workers’ strike of 1964. Cold War History, 10(4):537–569, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682741003603102.

Webb:2013:CWR

[Web13] Alban Webb. Cold War radio and the Hungarian uprising, 1956. Cold War History, 13(2):221–238, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2012.746667.

Webb:2021:HRC

[Web21] Brandon Webb. ‘How to raise a curtain’: security, surveillance, and mobility in Canada’s Cold War-era exchanges, 1955–65. Cold War History, 21(2):215–233, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2020.1724101.

Webb:2022:RNM

[Web22] Mattie C. Webb. Research note: Mayibuye Archives and the Cold War in Southern Africa. Cold War History, 22(3):369–373, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2022.2076347.

Weeks:2021:LAG

[Wee21] Gregory Weeks. Latin America and the Global Cold War. Cold War History, 21(2):240–242, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1886909.

Wenderski:2022:AVP

[Wen22] Michał Wenderski. Art versus politics: Polish–Dutch international cultural relationships at the outset of the Cold War (1947–50). Cold War History, 22(1):103–121, 2022. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2021.1898378.

Westad:2006:DDU

[Wes06] O. A. Westad. Devices and desires: On the uses of Cold War history. Cold War History, 6(3):373–376, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740600810805.

Westad:2010:BRA

[Wes10] Odd Arne Westad. Book review: 1989: L’année où le monde a basculé. [1989: The year the world turned upside down] by Pierre Grosser, Paris, Perrin, 2009, 605 pp. Cold War History, 10(4):574–575, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.513509.

Westad:2012:BRS

[Wes12] Odd Arne Westad. Book review: Storia della guerra fredda: l’ultimo conflitto per l’Europa. by Federico Romero, Turin, Einaudi, 2009, 356 pp. [History of the Cold War: the last conflict for Europe]. Cold War History, 12(4):705–706, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2012.737560.

Westad:2017:CEC

[Wes17] Odd Arne Westad. China and the end of the Cold War in Europe. Cold War History, 17(2):111–113, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2017.1315956.

Wettig:2009:LSO

[Wet09] Gerhard Wettig. The last Soviet offensive in the Cold War: emergence and development of the campaign against NATO euromissiles, 1979– 1983. Cold War History, 9(1):79–110, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682740802638640.

White:2009:COU

[Whi09] Mark White. The Castro obsession: U.S. covert operations against Cuba, 1959–1965. Cold War History, 9(3):451–452, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740903135496.

White:2010:USI

[Whi10] Stephen White. Uncivil society: 1989 and the implosion of the Communist establishment. Cold War History, 10(4):572–574, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2010.513506.

White:2011:JFK

[Whi11] Mark J. White. John F. Kennedy: world leader. Cold War History, 11(3):474–475, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.599977.

Widmer:2018:NCC

[Wid18] Sabina Widmer. Neutrality challenged in a cold war conflict: Switzerland, the International Committee of the Red Cross, and the Angolan War. Cold War History, 18(2):203–220, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/14682745.2017.1408072.

Williams:2011:CWM

[Wil11] Rhodri Williams. The Cold War in the Middle East: regional conflict and the superpowers 1967–73. Cold War History, 11(2): 289–290, 2011. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2011.569166.

Williams:2024:DCW

[Wil24] Brandon Kirk Williams. Diasporic Cold Warriors: Nationalist China, anticommunism, and the Philippine Chinese, 1930s–1970s. Cold War History, 24(1):155–157, 2024. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2022.2121360.

Winokur:2009:SWJ

[Win09] Talia Winokur. ‘The Soviets were just an excuse’: why Israel did not destroy the Egyptian Third Army. Cold War History, 9(1): 59–78, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802573458.

Wenger:2008:SCB

[WN08] Andreas Wenger and Christian Nuenlist. A ‘special case’ between independence and interdependence: Cold War studies and Cold War politics in post-Cold War Switzerland. Cold War History, 8(2): 213–240, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740802018777.

Wohlforth:2006:SIT

[Woh06] William Wohlforth. Superpowers, interventions and the Third World. Cold War History, 6(3):365–371, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740600795535.

Wolff:2010:SCW

[Wol10] David Wolff. Stalin’s Cold War: Soviet foreign policy, democracy and communism in Bulgaria, 1941–48. Cold War History, 10(1): 131–132, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740903398987.

Woodroofe:2010:CCH

[Woo10] Louise P. Woodroofe. Carter’s conversion: the hardening of American defense policy. Cold War History, 10(2):280–281, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682741003711913.

Woodward:2013:ECY

[Woo13] Susan L. Woodward. Europe and the collapse of Yugoslavia: The role of non-state actors and European diplomacy. Cold War History, 13 (3):437–438, 2013. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2013.819650.

Wenger:2001:CDS

[WS01] A. Wenger and J. Suri. At the crossroads of diplomatic and social history: The nuclear revolution, dissent and Détente. Cold War History, 1(3):1–42, 2001. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999931.

Wyss:2012:NEC

[Wys12] Marco Wyss. Neutrality in the early Cold War: Swiss arms imports and neutrality. Cold War History, 12(1):25–49, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2010.536534.

Xiao:2019:RBE

[XZ19] Shao Xiao and Xiaoming Zhang. Reassessment of Beijing’s economic and military aid to Hanoi’s War, 1964–75. Cold War History, 19(4): 549–567, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1581768.

Yamamoto:2007:DIE

[Yam07] Takeshi Yamamoto. Détente or integration? EC response to Soviet policy change towards the Common Market, 1970–75. Cold War History, 7(1):75–94, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740701197680.

Yegorova:2006:RAP

[Yeg06] Natalia I. Yegorova. Russian archives: Prospects for Cold War studies. Cold War History, 6(4):543–548, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740600979311.

Young:2002:BLW

[You02] J. W. Young. Britain and ’LBJ’s war’, 1964–68. Cold War History, 2(3):63–92, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
713999965.

Young:2006:R

[You06a] John W. Young. Reviews. Cold War History, 6(1):137–152, 2006. CO-DEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/14682740500395535.

Young:2006:TTM

[You06b] Marilyn B. Young. Two, three, many Vietnams. Cold War History, 6(4):413–424, 2006. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740600979238.

Zaccaria:2020:YIE

[Zac20] Benedetto Zaccaria. Yugoslavia, Italy, and European integration: was Osimo 1975 a Pyrrhic victory? Cold War History, 20(4): 503–520, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1657094.

Zaidi:2009:ESA

[Zai09] Waqar Zaidi. Eisenhower, science advice, and the nuclear test-ban Debate, 1945–1963. Cold War History, 9(4):536–537, 2009. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682740903268602.

Zanchetta:2012:LMS

[Zan12] Barbara Zanchetta. L’appuntamento mancato: La sinistra italiana e il Dissenso nei regimi comunisti (1968–1989). (Italian) [The missed appointment: The Italian Left and dissent in Communist regimes (1968–1989)]. Cold War History, 12(1):178–179, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2012.655450.

Zaremba:2005:KWP

[Zar05] Marcin Zaremba. Karol Wojtyła the Pope: Complications for comrades of the Polish United Workers’ Party. Cold War History, 5(3):

317–336, 2005. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740500222044.

Zaum:2010:UNV

[Zau10] Dominik Zaum. The United Nations: a very short introduction. Cold War History, 10(2):273–274, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682741003711939.

Zubok:2021:CWE

[ZCP^{+}21] Vladislav Zubok, Michael Cox, Vladimir O. Pechatnov, Rodric Braithwaite, Kristina Spohr, Sergey Radchenko, Sergey Zhuravlev, Isaac Scarborough, Svetlana Savranskaya, and M. E. Sarotte. A Cold War endgame or an opportunity missed? Analysing the Soviet collapse thirty years later. Cold War History, 21(4):541–599, 2021. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/14682745.2021.1987871.

Zeilig:2017:MR

[Zei17] Leo Zeilig. Mecca of revolution. Cold War History, 17(3):321– 323, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2017.1341371.

Zou:2019:PEW

[ZF19] Yizheng Zou and Hongwei Fan. The political economy of Welfare State and the China–Burma relationship. Cold War History, 19(4): 569–585, 2019. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2019.1581765.

Zhang:2007:CPC

[Zha07] Shu Guang Zhang. Constructing ‘Peaceful Coexistence’: China’s diplomacy toward the Geneva and Bandung Conferences, 1954–55. Cold War History, 7(4):509–528, 2007. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740701621846.

Zhu:2012:HRO

[Zhu12] Dandan Zhu. The Hungarian revolution and the origins of China’s Great Leap policies, 1956–57. Cold War History, 12(3):451–472, 2012. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/14682745.
2011.626771.

Zhuk:2014:HIC

[Zhu14] Sergei Zhuk. Hollywood’s insidious charms: the impact of American cinema and television on the Soviet Union during the Cold War. Cold War History, 14(4):593–617, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2014.950252.

Ziauddin:2020:LAB

[ZM20] Silvia Berger Ziauddin and Sibylle Marti. Life after the Bomb: Nuclear fear, science, and security politics in Switzerland in the 1980s. Cold War History, 20(1):95–113, 2020. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2018.1536121.

Zofka:2018:CRM

[Zof18] Jan Zofka. China as a role model? The ‘Economic Leap’ campaign in Bulgaria (1958–1960). Cold War History, 18(3):325–342, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2017.1336158.

Zubok:2002:GEC

[Zub02] V. Zubok. Gorbachev and the end of the Cold War: Perspectives on history and personality. Cold War History, 2(2):61–100, 2002. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/713999954.

Zubok:2008:SUD

[Zub08] Vladislav Zubok. The Soviet Union and détente of the 1970s. Cold War History, 8(4):427–447, 2008. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682740802373537.

Zubok:2010:RFC

[Zub10] Vladislav M. Zubok. The rise and fall of communism. Cold War History, 10(3):457–458, 2010. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2010.493718.

Zubok:2014:HBA

[Zub14] Vladislav Zubok. With his back against the Wall: Gorbachev, Soviet demise, and German reunification. Cold War History, 14(4):

619–645, 2014. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2014.950251.

Zubok:2015:GAS

[Zub15] Vladislav Zubok. Goodbye to all that? The story of Europe since 1945. Cold War History, 15(3):424–426, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2015.1051365.

Zubok:2016:BRS

[Zub16] Vladislav Zubok. Book review: Stephen Kotkin, Stalin, Vol. I: Paradoxes of Power, 1878–1928 (New York: Allen Lane, 2014), xiii + 949 pp. Cold War History, 16(2):231–233, 2016. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/14682745.2016.1153851.

Zubok:2017:SUC

[Zub17] Vladislav Zubok. The Soviet Union and China in the 1980s: reconciliation and divorce. Cold War History, 17(2):121–141, 2017. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/14682745.2017.1315923.

Zofka:2018:BKR

[ZVU18] Jan Zofka, Péter Vámos, and Sören Urbansky. Beyond the Kremlin’s reach? Eastern Europe and China in the Cold War era. Cold War History, 18(3):251–256, 2018. CODEN ???? ISSN 1468-2745 (print), 1743-

7962 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/14682745.2018.1453253.

Zhai:2015:SPC

[ZX15] Xiang Zhai and Ruping Xiao. Shifting political calculation: the secret Taiwan–Soviet talks, 1963–1971. Cold War History, 15(4): 533–556, 2015. CODEN ???? ISSN 1468-2745 (print), 1743-7962

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682745.2015.1031115.